### v3

# Semiparametric Double Reinforcement Learning with Applications to Long-Term Causal Inference

Abstract
Long-term causal effects often must be estimated from short-term data due to limited follow-up in healthcare, economics, and online platforms. Markov Decision Processes (MDPs) provide a natural framework for capturing such long-term dynamics through sequences of states, actions, and rewards. Double Reinforcement Learning (DRL) enables efficient inference on policy values in MDPs, but nonparametric implementations require strong intertemporal overlap assumptions and often exhibit high variance and instability. We propose a semiparametric extension of DRL for efficient inference on linear functionals of the $Q$-function—such as policy values—in infinite-horizon, time-homogeneous MDPs. By imposing structural restrictions on the $Q$-function, our approach relaxes the strong overlap conditions required by nonparametric methods and improves statistical efficiency. Under model misspecification, our estimators target the functional of the best-approximating $Q$-function, with only second-order bias. We provide conditions for valid inference using sieve methods and data-driven model selection. A central challenge in DRL is the estimation of nuisance functions, such as density ratios, which often entail difficult minimax optimization. To address this, we introduce a novel plug-in estimator based on isotonic Bellman calibration, which combines fitted $Q$-iteration with an isotonic regression adjustment. The estimator is debiased without requiring estimation of additional nuisance functions and reduces high-dimensional overlap assumptions to a one-dimensional condition. Bellman calibration extends isotonic calibration—widely used in prediction and classification—to the MDP setting and may be of independent interest.
Keywords— Automatic debiasing, infinite-horizon MDPs, time-homogeneous dynamics, long-term causal inference, semiparametric restrictions, policy evaluation, isotonic calibration
1 Introduction
1.1 Motivation: Long-term causal inference
Randomized experiments—such as A/B tests and controlled trials—are widely used in healthcare, technology, and other industries to assess the impact of interventions on outcomes like survival, customer retention, and revenue. In industry, such experiments are often short-term due to practical constraints and the need for rapid decision-making. Consequently, analysts typically evaluate interventions using proxy metrics such as user engagement or click-through rates. These short-term outcomes inform decisions intended to improve long-term objectives. However, short-term experiments only yield unbiased estimates of short-term effects. This limitation has spurred growing interest in methods for inferring the long-term causal effects of policies from short-term data.
A common strategy for inferring long-term effects from short-term data involves surrogate methods. These approaches aim to link short-term experimental results to long-term outcomes by identifying intermediate variables, or “surrogates,” that are measured during the experiment (Athey et al., 2019). For example, a streaming platform may use engagement metrics—such as viewing hours or click rate—as surrogates for outcomes like annual membership retention. Given observational data containing long-term outcomes, these methods estimate the long-term causal effect by leveraging the surrogate–outcome relationship. A key assumption of surrogate methods is that the surrogate fully mediates the treatment effect. However, this assumption is violated when treatments involve sustained exposure beyond the experiment’s duration. For instance, evaluating a personalized recommendation algorithm—deployed continuously and adapting to user interactions—cannot rely on a short-term surrogate to capture the cumulative long-term effect, limiting the applicability of surrogate-based approaches.
These limitations motivate the use of dynamic modeling frameworks, which account for how long-term outcomes evolve in response to sequences of states, actions, and rewards under sustained treatment. For example, on a streaming platform, annual user retention under a new recommendation algorithm depends on daily engagement, shifting content preferences, and monthly subscription renewals. Recognizing these dynamics, Tran et al. (2023) proposed a method for estimating long-term effects of sustained treatments from short-term experiments, assuming that short-term observations sufficiently capture the long-term trajectory, even if they do not fully mediate the effect (see also Example 2 of Bibaut et al. (2021)). The core idea is to model the experiment’s temporal dynamics as a time-homogeneous Markov Decision Process (MDP) (Puterman, 1990) (see Figure 1), imposing Markov independence and stationarity on the state-action-outcome process. By linking long-term causal inference with offline reinforcement learning (Kaelbling et al., 1996), the authors develop nonparametric, efficient estimators of long-term treatment effects using Double Reinforcement Learning (van Der Laan et al., 2018; Kallus and Uehara, 2020, 2022).
Figure 1: DAG for trajectory under Markov Decision Process. The outcome $Y_{1}$, state $S_{2}$, and the trajectory need not be observed in the experiment.
Double Reinforcement Learning (DRL) enables statistically efficient inference on the value of a policy in nonparametric MDPs from off-policy trajectories. However, DRL faces fundamental challenges. First, DRL require sufficient overlap between the initial and future state distributions to ensure that all states relevant to the long-term trajectory are visited with positive probability. In high-dimensional or unbounded state spaces, this assumption is often violated (Mehrabi and Wager, 2024). Limited intertemporal overlap increases estimator variance, degrades stability, and necessitates large sample sizes. This challenge is analogous to, but distinct from, the poor performance of inverse propensity weighted estimators in cross-sectional studies with limited treatment overlap (D’Amour et al., 2021), as intertemporal state overlap cannot be ensured by randomization, given that future states are causally determined by prior states and actions. Second, DRL requires the estimation of complex nuisance functions via minimax optimization, which is computationally intensive and can be unstable in finite samples.
1.2 Contributions of this work
We develop a semiparametric extension of Double Reinforcement Learning (DRL) for inference on linear functionals of the $Q$-function—such as policy values—in infinite-horizon, time-homogeneous Markov Decision Processes (MDPs). The $Q$-function, which generalizes the regression function from static settings, encodes the expected cumulative outcome given the current state and action. Our approach addresses two key challenges in DRL: sensitivity to limited intertemporal overlap and the difficulty of estimating nuisance functions, such as density ratios. By imposing semiparametric restrictions on the $Q$-function, we relax the overlap conditions required for identification and improve the efficiency of resulting estimators.
Our key contributions are as follows:
  1. 1.
    We propose automatic debiased estimators for linear functionals of the $Q$-function—such as policy values—in time-homogeneous Markov Decision Processes (MDPs), under a semiparametric model for the $Q$-function. These estimators are doubly robust with respect to the estimation error of the $Q$-function and a certain Riesz representer of the linear functional. We further show how these estimators can be adjusted to enable valid inference for the best-approximating $Q$-function under model misspecification.
  2. 2.
    We show that misspecification of the $Q$-function incurs only second-order bias for the target functional, enabling valid inference under mild misspecification. We then provide conditions for valid inference using sieve methods and data-driven model selection.
  3. 3.
    We propose a novel debiased plug-in estimator for nonparametric inference that addresses two key challenges in DRL: (i) the computational burden and instability of minimax nuisance estimation, and (ii) sensitivity to model misspecification. The core of our approach is isotonic Bellman calibration, a generalization of isotonic calibration to MDPs, which may be of independent interest. We show that calibrating the $Q$-function estimator alone suffices to debias the plug-in estimator and provide valid nonparametric inference, without requiring additional nuisance estimation, such as density ratios.
Semiparametric restrictions have long been used to address instability in causal inference, particularly in cross-sectional settings. For example, nonparametric estimators of the average treatment effect (ATE), such as the augmented inverse probability weighted (AIPW) estimator (Robins et al., 1994), often suffer from high variance under limited treatment overlap. To mitigate this, prior work has leveraged working model assumptions (e.g., treatment effect homogeneity or partially linear models (Crump et al., 2006; Li et al., 2019; Robinson, 1988)) and dimension reduction techniques (Benkeser et al., 2020; D’Amour and Franks, 2021). Related ideas have been applied in data fusion for hidden confounding (Kallus et al., 2018), improving efficiency in randomized trials (van der Laan et al., 2024d), and enabling data-driven model selection (van der Laan et al., 2023). We extend these ideas to dynamic settings, addressing both treatment and intertemporal overlap, and the more complex nuisance estimation challenges that arise in MDPs.
Our work contributes to the growing literature on inference for off-policy evaluation (Murphy, 2003; Liu et al., 2018; Tang et al., 2019; Shi et al., 2022; Wang et al., 2023a), particularly through debiased and doubly robust estimation techniques (Tang et al., 2019; Kallus and Uehara, 2020; Shi et al., 2021; Kallus and Uehara, 2022; Mehrabi and Wager, 2024), as well as to research on long-term causal inference using MDPs and reinforcement learning methods (Liao et al., 2021; Tran et al., 2023; Nam et al., 2024). Existing approaches primarily target nonparametric, parametric, or sieve-based models of the $Q$-function, and focus on a specific linear functional: the policy value. We extend this literature by developing debiased inference procedures for general linear functionals of the $Q$-function under semiparametric restrictions. The challenge of limited intertemporal overlap was recently addressed by Mehrabi and Wager (2024), who proposed modified DRL estimators for policy values that adaptively truncate density ratios to improve performance when the ratio is unbounded but has finite variance. In contrast, our approach both relaxes overlap requirements and improves efficiency by imposing semiparametric structure on the $Q$-function. As a result, our methods remain valid even in settings where the density ratio has infinite variance or does not exist, rendering nonparametric identification impossible.
This paper is organized as follows. Section 2 introduces Markov decision processes, $Q$-functions, and the target estimand. Section 3 presents our semiparametric DRL estimators, the corresponding asymptotic theory, and extensions. Section 4 discusses nuisance estimation strategies and associated challenges. In Section 5, we propose our Bellman-calibrated plug-in estimator. Finally, Section 6 presents numerical experiments.
2 Preliminaries
2.1 Data Structure and Markov Decision Model
We consider a randomized experiment or observational study where participants sequentially receive treatments (actions) based on a policy. At each time $t$, a participant occupies a state $S_{t}$, which informs the choice of action $A_{t}$. The state-action pair determines an intermediate outcome $Y_{t}$, interpreted as an immediate reward or cost, and influences the transition to the next state $S_{t+1}$. Although participants are typically observed over a short time horizon, our goal is to estimate the long-term causal effects of a target policy $\pi$, which may differ from the behavior policy generating the observed data.
Formally, we represent an individual’s trajectory by the sequence $(S_{0},A_{0},Y_{0},S_{1},A_{1},Y_{1},S_{2},A_{2},\dots)\sim\mathbb{P}_{0}$, where states $S_{t}\in\mathcal{S}\subseteq\mathbb{R}^{d}$, actions $A_{t}\in\mathcal{A}$, and outcomes $Y_{t}\in\mathcal{Y}\subseteq\mathbb{R}$. The observed data consist of $n$ i.i.d. samples of a single state transition $(S_{0},A_{0},Y_{0},S_{1})$ from a distribution $P_{0}$ in a nonparametric model $\mathcal{P}$, forming the dataset $\mathcal{D}_{n}:=\{(S_{0,i},A_{0,i},Y_{0,i},S_{1,i}):i\in[n]\}$. This simplification entails little loss of generality, as multiple transitions from a single observation can be decomposed into individual transitions. Although such transitions may be dependent, our theoretical results extend using central limit theorems for Markov chains (Bibaut et al., 2021). We let $\pi(a\mid s)$ denote the (stationary) policy of interest, where $\pi(a\mid S_{t})$ represents the conditional probability of selecting action $a\in\mathcal{A}$ at time $t$ given the current state $S_{t}\in\mathcal{S}$. We denote by $\mathbb{P}_{0}$ the distribution of complete long-term trajectories induced by $P_{0}$, and by $P_{0,A_{0},S_{0}}$ the marginal distribution of $(A_{0},S_{0})$. To simplify notation, we write $S_{0}$ for any summary $S_{P_{0}}$ of the true distribution $P_{0}$.
We assume the distribution $P_{0}$ of short-term observations $(S_{0},A_{0},Y_{0},S_{1})$ fully determines the distribution $\mathbb{P}_{0}$ of the long-term trajectory. Formally, we posit that the state-action-outcome process follows a time-homogeneous Markov decision process (Puterman, 1990), with data sequentially generated according to the nonparametric structural equation model (NPSEM) (Pearl, 2012):
$A_{t}:=f_{A}(S_{t},U_{A_{t}});\quad Y_{t}:=f_{Y}(A_{t},S_{t},U_{Y_{t}});\quad S_{t+1}$ $:=f_{S}(Y_{t},A_{t},S_{t},U_{S_{t+1}}),$
where $f_{A}$, $f_{Y}$, and $f_{S}$ are unknown deterministic functions, and the latent variables $\{U_{A_{t}},U_{Y_{t}},U_{S_{t+1}}\}$ are unobserved, mutually independent, and stationary random variables. This model imposes the Markovian assumptions that $A_{t}$ depends on its history only through $S_{t}$, $Y_{t}$ depends only on $A_{t}$ and $S_{t}$, and $S_{t+1}$ depends only on $Y_{t}$, $A_{t}$, and $S_{t}$. It also assumes stationarity over time, meaning that the conditional distributions $(S_{t+1}\mid Y_{t},A_{t},S_{t})$ and $(Y_{t}\mid A_{t},S_{t})$ are time-invariant. These assumptions become more plausible with a richer state space. For example, the state can be augmented to include multiple past time points, such as $\widetilde{S}_{t}:=(S_{t},S_{t-1},\dots,S_{t-k})$, or constructed using fixed, finite-dimensional summaries of historical information (van der Laan and Malenica, 2018).
The NPSEM allows us to define counterfactual MDPs, as the following example illustrates.
Example 1 (Policy Value in an MDP).
Given a policy $\pi$, let $\{S_{t}(\pi),A_{t}(\pi),Y_{t}(\pi):t\in\mathbb{T}\}\sim\mathbb{P}_{0}^{\pi}$ denote the counterfactual trajectory under a Markov decision process, where $A_{t}$ is drawn from $\pi(\cdot\mid S_{t})$. This trajectory is defined by intervening on the structural equation for $A_{t}$ in the NPSEM so that $A_{t}^{\prime}$ follows $\pi(\cdot\mid S_{t})$. The value of the policy at time $t$ is $\mathbb{E}_{0}^{\pi}[Y_{t}(\pi)]$, and the expected discounted cumulative outcome, for discount factor $\gamma\in[0,1]$, is $\mathbb{E}_{0}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}Y_{t}(\pi)\right]$, where $\mathbb{E}_{0}^{\pi}$ denotes the expectation under $\mathbb{P}_{0}^{\pi}$. ∎
2.2 Inferential objective and challenges in identification
In this work, we aim to estimate and perform inference on linear functionals of the $Q$-function (Kaelbling et al., 1996), a fundamental quantity connecting the causal effects of long-term policies to the observed short-term data. Formally, given a discount factor $\gamma\in[0,1]$, the $Q$-function associated with a policy $\pi$ is defined as the mapping:
$q_{0}^{\pi}(a,s):=\mathbb{E}_{0}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}Y_{t}(\pi)\mid A_{0}=a,S_{0}=s\right],$
where $\mathbb{E}_{0}^{\pi}$ denotes expectation under the counterfactual MDP induced by policy $\pi$. Throughout, we omit the superscript $\pi$ and write $q_{0}$ in place of $q_{0}^{\pi}$ when the policy under consideration is understood to be $\pi$. Intuitively, the $Q$-function represents the expected cumulative reward (or cost) for an individual who begins in state $s$, takes initial action $a$, and thereafter follows the policy $\pi$. The discount factor $\gamma$ determines the weight assigned to future rewards, controlling the time horizon for evaluation. By convention, we set $10^{0}:=$ 1; thus, when $\gamma=0$, the $Q$-function reduces to the short-term outcome regression $(a,s)\mapsto E_{0}[Y_{0}\mid A_{0}=a,S_{0}=s]$, and we recover the well-studied problem of inference on linear functionals of the outcome regression (Chernozhukov et al., 2018a, 2022; van der Laan et al., 2024b).
[…]
Statistically efficient estimation of the policy value, $E_{0}\left[V^{\pi}(q_{0})(S_{0})\right]$—a special case of a linear functional—via DRL has been studied under nonparametric models by van Der Laan et al. (2018), Kallus and Uehara (2020, 2022), and Tran et al. (2023). However, nonparametric inference on policy value is challenging, requiring strong conditions for identifiability and root-$n$ estimation. In particular, it depends on the existence and finite variance of the importance-weighted state occupancy ratio:
$d_{0}(a,s):=\frac{\pi(a\mid s)}{b_{0}(a\mid s)}\sum_{t=0}^{\infty}\gamma^{t}\frac{d\mathbb{P}_{0}^{\pi}(S_{t}=s)}{d\mathbb{P}_{0}^{b_{0}}(S_{0}=s)},$ (2)
where $\pi(a\mid s)$ is the target policy and $b_{0}(a\mid s)=P_{0}(A_{0}=a\mid S_{0}=s)$ is the behavior policy (Mehrabi and Wager, 2024). This condition requires overlap between the target and behavior policies, as well as between future and initial state distributions. Such overlap is often difficult to satisfy in high-dimensional settings, especially when interventions induce rare or novel states (D’Amour et al., 2021; Mehrabi and Wager, 2024), and even near-violations can cause high estimator variability. The asymptotic variance of existing DRL estimators and the Cramér–Rao bound depend critically on the variability of $d_{0}$.
A notable example of a linear functional of the $Q$-function is the long-term causal effect in a randomized experiment (Tran et al., 2023).
Example 2 (Long-term causal effect in A/B test).
Consider an A/B test where individuals are randomly assigned to one of two arms: either the intervention of interest or a control (such as no intervention or an alternative). In this case, the data-generating policy $b_{0}$ is static: $A_{t}:=Z$ for all $t\geq 0$, where $Z$ is the treatment arm indicator, with $Z=1$ representing the treatment arm and $Z=0$ the control. Following Tran et al. (2023), the long-term causal effect of treatment is defined as $\psi_{0}:=\mathbb{E}_{0}\left[\sum_{t=0}^{\infty}\gamma^{t}\{Y_{t}(1)-Y_{t}(0)\}\right],$ where $Y_{t}(1)$ and $Y_{t}(0)$ are the potential outcomes at time $t$ under treatment and control, respectively. Decompose the state as $S_{t}:=(Z,\widetilde{S}_{t})$, where $\widetilde{S}_{t}$ is a subvector of the state excluding the study assignment $Z$. Define the behavior policy as $b_{0}(a\mid(z,\widetilde{s})):=\mathbbm{1}\{a=z\}$. Then, $\psi_{0}$ can be expressed as $\psi_{0}=E_{0}\left[q_{0}^{b_{0}}(1,(1,\widetilde{S}_{0}))-q_{0}^{b_{0}}(0,(0,\widetilde{S}_{0}))\right],$ which is a linear functional of the $Q$-function $q_{0}^{b_{0}}$ corresponding to $m((z,\widetilde{s}),a,q)=q(1,(1,\widetilde{s}))-q(0,(0,\widetilde{s}))$. ∎
2.3 The Bellman equation as an integral equation
In this work, it is useful to consider an alternative identification of $q_{0}$ as the solution to a linear inverse problem. Let $\lambda$ be a measure on $\mathcal{S}\times\mathcal{A}$ that dominates the distributions of $(A_{0},S_{0})$ and $(A_{0},S_{1})$ almost surely for all $P\in\mathcal{P}$, and let $L^{\infty}(\lambda)$ denote the Banach space induced by the $\lambda$-essential supremum norm. By rearranging terms in (1), the $Q$-function $q_{P}$ for each $P\in\mathcal{P}$ is identified as the solution to the integral equation
[…]
where the outcome regression is defined by $\mu_{P}(a,s):=E_{P}[Y_{0}\mid A_{0}=a,S_{0}=s]$, and the Bellman integral operator $\mathcal{T}_{P}:L^{\infty}(\lambda)\rightarrow L^{\infty}(\lambda)$ is defined as as $\mathcal{T}_{P}(h)(a,s):=h(a,s)-\gamma E_{P}\left[V_{h}^{\pi}(S_{1})\mid A_{0}=a,S_{0}=s\right].$ Consequently, our estimand $\psi_{0}$ corresponds to the $P_{0}$-evaluation of the target parameter $\Psi:P\mapsto E_{P}\left[m(S_{0},A_{0},q_{P})\right]$, defined over the nonparametric model $\mathcal{P}$. Inference on $\psi_{0}$ thus reduces to inference on a linear functional of the solution to a linear inverse problem (Ai and Chen, 2003, 2012; Bennett et al., 2022, 2023a, 2023b)—a perspective we exploit in our theoretical analysis.
3 Semiparametric double reinforcement learning
3.1 Proposed estimator
In this section, we propose automatic DRL estimators for the estimand $\psi_{0}$ that impose semiparametric restrictions on the $Q$-function and outline their asymptotic properties. These estimators are automatic in the sense that they require only the specification of a linear functional, from which a debiasing procedure is derived. In later sections, we develop formal theory for semiparametric DRL and introduce model-robust and model-adaptive variants that remain valid under misspecification.
Suppose the $Q$-function $q_{0}$ lies in a (possibly infinite-dimensional) subspace $H\subset L^{\infty}(\lambda)$ of $\lambda$-essentially bounded functions, and let $\mathcal{P}_{H}:=\{P\in\mathcal{P}:q_{P}\in H\}$ denote the corresponding semiparametric model for $P_{0}$. For example, $q_{0}$ may be partially linear, additive, or depend only on a subset of the state variables. A possible choice for $H$ is the partially linear model (Robinson, 1988), which assumes $q_{0}(A_{0},S_{0})=q_{0}(0,S_{0})+A_{0}\beta_{0}^{T}S_{0}$. Here, the control function $q_{0}(0,S_{0})$ is estimated nonparametrically, while the treatment effect $q_{0}(1,S_{0})-q_{0}(0,S_{0})$ is modeled linearly in $S_{0}$. A natural estimator of $\psi_{0}=\Psi(P_{0})$ is the plug-in estimator $n^{-1}\sum_{i=1}^{n}m(S_{0,i},q_{n,H})$, where $q_{n,H}\in H$ is an estimator of $q_{0}$. One approach to estimating $q_{0}$ is through fitted Q-iteration (FQI) (Munos and Szepesvári, 2008), which we discuss in Section 4. However, when $q_{n,H}$ is obtained via flexible learning methods, the plug-in estimator typically lacks $\sqrt{n}$-consistency and asymptotic normality due to first-order bias from $q_{n,H}-q_{0}$ (Van der Laan et al., 2011; Chernozhukov et al., 2018a). To address this, debiasing techniques are used to eliminate first-order bias and restore asymptotic properties.
To construct a bias correction, we require the linear functional $\psi_{0}=E_{0}[m(S_{0},A_{0},q_{0})]$ to be continuous not only in the $Q$-function $q_{0}$, but also in the outcome regression $\mu_{0}$. These quantities are linked through the linear inverse problem $\mathcal{T}_{0}(q_{0})=\mu_{0}$ in (3). To formalize this, we introduce the following notation and condition. For $P\in\mathcal{P}$, let $\langle\cdot,\cdot\rangle_{P}$ denote the $L^{2}(P_{A_{0},S_{0}})$ inner product, where $P_{A_{0},S_{0}}$ is the marginal distribution of $(A_{0},S_{0})$ under $P$, and let $\|\cdot\|_{P}$ be the associated norm. The closure of $H$ with respect to $\|\cdot\|_{P}$ is denoted $\overline{H}_{P}$, and we define $\mathcal{T}_{P}(\overline{H}_{P}):=\{\mathcal{T}_{P}(h):h\in\overline{H}_{P}\}$.
  1. (C1)
    For all $P\in\mathcal{P}$ in a Hellinger neighborhood of $P_{0}$, the following hold:
    1. (i)
      (Functional continuity) There exists $C<\infty$ such that $|E_{P}[m(S_{0},A_{0},q)]|\leq C\|q\|_{P}$ for all $q\in H$.
    2. (ii)
      (Bounded inverse) The operator $\mathcal{T}_{P}$ is continuous as a map from $(H,\|\cdot\|_{P})$ to $L^{2}(P_{A_{0},S_{0}})$, and its unique extension to $\overline{H}_{P}$ has a continuous inverse on its range.
As a consequence of Condition C1, we have $q_{0}=\mathcal{T}_{0}^{-1}(\mu_{0})$ and $\psi_{0}=E_{0}[m(S_{0},A_{0},\mathcal{T}_{0}^{-1}(\mu_{0}))]$, where the map $\mu\mapsto E_{0}[m(S_{0},A_{0},\mathcal{T}_{0}^{-1}(\mu))]$ is continuous on $\mathcal{T}_{0}(\overline{H}_{P_{0}})$. This condition implies that the functional $q\mapsto E_{0}[m(S_{0},A_{0},q)]$ is continuous with respect to the weak norm $\|\mathcal{T}_{0}(\cdot)\|_{P_{0}}$ induced by the Bellman operator, in the sense that
$\sup_{q\in H}\frac{E_{0}[m(S_{0},A_{0},q)]}{\|\mathcal{T}_{0}(q)\|_{P_{0}}}<\infty.$ (4)
In ill-posed inverse problems, conditions akin to (4) are often assumed directly to enable debiased estimation without requiring invertibility of $\mathcal{T}_{0}$ (Bennett et al., 2022). In our setting, however, the inverse problem defining $q_{0}$ is well-posed, and C1 holds under mild regularity conditions (Chen and Qi, 2022). Specifically, we show in Appendix A that $\mathcal{T}_{0}$ is a Fredholm operator of index zero, and that the inverse problem defining $q_{0}$ is a Fredholm equation of the second kind (Conway, 1994).
The boundedness property in (4) yields a dual representation of the target parameter $\psi_{0}$ as a weighted expectation of the outcome $Y_{0}$ (Bennett et al., 2022, 2023b). The Riesz representation theorem ensures the existence of a representer $\alpha_{0,H}\in\overline{H}_{P_{0}}$ such that
$\psi_{0}=\langle\mathcal{T}_{0}(\alpha_{0,H}),\mathcal{T}_{0}(q_{0,H})\rangle_{P_{0}}=\langle\mathcal{T}_{0}(\alpha_{0,H}),\mu_{0}\rangle_{P_{0}},$
and hence $\psi_{0}=E_{0}[\mathcal{T}_{0}(\alpha_{0,H})(A_{0},S_{0})Y_{0}]$. The weighting function $\mathcal{T}_{0}(\alpha_{0,H})$ equals the inverse propensity weight function for the counterfactual mean $E_{0}[\mu_{0}(1,S_{0})]$ when $\gamma=0$, and equals the state occupancy ratio $d_{0}$ in (2) for the policy value estimand more generally. The Riesz representer is characterized as the minimizer of
$\alpha_{0,H}\in\arg\min_{\alpha\in\overline{H}_{P_{0}}}E_{0}\left[\left\{\mathcal{T}_{0}(\alpha)(A_{0},S_{0})\right\}^{2}-2m(S_{0},A_{0},\alpha)\right],$ (5)
a key fact underlying automatic Debiased Machine Learning (DML) (Chernozhukov et al., 2022; Bennett et al., 2023b; van der Laan et al., 2025).
Building on the weighted outcome representation, we propose an automatic DRL estimator that augments the plug-in estimator with a bias correction term. Specifically, given estimators $\widehat{\mathcal{T}}_{n}$, $\alpha_{n,H}$, and $q_{n,H}$ of $\mathcal{T}_{0}$, $\alpha_{0,H}$, and $q_{0}$, respectively, the autoDRL estimator $\psi_{n,H}$ takes the form of a one-step debiased estimator:
$\psi_{n,H}:=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},q_{n,H})+\frac{1}{n}\sum_{i=1}^{n}\widehat{\mathcal{T}}_{n}(\alpha_{n,H})(A_{0,i},S_{0,i})\left\{Y_{0,i}+\gamma V^{\pi}(q_{n,H})(S_{1,i})-q_{n,H}(A_{0,i},S_{0,i})\right\},$ (6)
where the second term is as an influence-function-based bias correction. This estimator is automatic in the sense that its form is agnostic to the specific linear functional, requiring only an estimate of the Riesz representer, which can be learned using (5). Note that we do not require an estimator $\widehat{\mathcal{T}}_{n}$ of the entire operator $\mathcal{T}_{0}$, but only an estimate of the evaluation $\mathcal{T}_{0}(\alpha_{n,H})$. Estimation of the nuisance functions is discussed in Section 4. This estimator generalizes the nonparametric DRL estimator of Kallus and Uehara (2020) from policy value estimation to generic continuous linear functionals under semiparametric model restrictions. To accommodate the use of flexible machine learning methods, we recommend cross-fitting the nuisance estimators $q_{n,H}$ and $\widehat{\mathcal{T}}_{n}(\alpha_{n,H})$, following standard practice in debiased machine learning (van der Laan et al., 2011; Chernozhukov et al., 2018a). For simplicity, we omit the cross-fitting notation and refer the reader to these references for implementation details.
3.2 Asymptotic properties and efficiency considerations
We now study the asymptotic behavior of the DRL estimator $\psi_{n,H}$. We first show that $\Psi:P\mapsto E_{P}[m(S_{0},A_{0},q_{P})]$ is pathwise differentiable with respect to the semiparametric model $\mathcal{P}_{H}$ (Bickel et al., 1993). We then establish the asymptotic linearity of $\psi_{n,H}$ and discuss how leveraging semiparametric structure can yield efficiency gains over fully nonparametric estimators.
The following theorem establishes the pathwise differentiability of the parameter $\Psi$. For each $P\in\mathcal{P}_{H}$, we define the influence function
$\varphi_{P,H}(s,a,y,s^{\prime}):=\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y+\gamma V^{\pi}(q_{P})(s^{\prime})-q_{P}(a,s)\}+m(s,a,q_{P})-\Psi(P),$ (7)
where $\alpha_{P,H}\in\overline{H}_{P}$ is the Riesz representer of $h\mapsto E_{P}[m(S_{0},A_{0},h)]$ with respect to $\|\mathcal{T}_{P}(\cdot)\|_{P}$.
Theorem 1 (Pathwise differentiability).
Suppose Condition C1 holds and that $P_{0}\in\mathcal{P}_{H}$. Then, $\Psi:\mathcal{P}_{H}\rightarrow\mathbb{R}$ is pathwise differentiable at $P_{0}$, with influence function $\varphi_{0,H}$. Moreover, for any $\overline{P}\in\mathcal{P}_{H}$ for which $\varphi_{\overline{P},H}$ exists, the following von Mises expansion holds:
$\Psi(\overline{P})-\Psi(P_{0})+P_{0}\varphi_{\overline{P},H}=\left\langle\mathcal{T}_{\overline{P}}(\alpha_{\overline{P},H})-\mathcal{T}_{0}(\alpha_{0,H}),\,\mathcal{T}_{0}(q_{0}-q_{\overline{P}})\right\rangle_{P_{0}}.$
Given an estimator $\widehat{P}_{n}\in\mathcal{P}_{H}$ of $P_{0}$, Theorem 1 shows that the influence function $\varphi_{\widehat{P}_{n},H}$ characterizes the bias of the plug-in estimator, $\Psi(\widehat{P}_{n})-\Psi(P_{0})$, up to second-order remainder terms. To correct for this bias, the proposed DRL estimator $\psi_{n,H}$ augments the plug-in estimator with the empirical mean of the influence function. In a more general setting, we will show in Section 3.3 that $\psi_{n,H}$ is asymptotically linear for $\psi_{0}$, with influence function $\varphi_{0,H}$, under conditions. Consequently, $\psi_{n,H}-\psi_{0}=P_{n}\varphi_{0,H}+o_{p}(n^{-1/2})$, and $\sqrt{n}\{\psi_{n,H}-\psi_{0}\}$ converges in distribution to a $\mathcal{N}(0,\sigma_{0}^{2})$ random variable with variance $\sigma_{0}^{2}:=E_{0}[\{\varphi_{0,H}(S_{0},A_{0},Y_{0},S_{1})\}^{2}]$. We further show that the DRL estimator is rate doubly robust, meaning that asymptotic linearity holds provided the nuisance estimators are consistent and satisfy the rate condition:
$\|\widehat{\mathcal{T}}_{n}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{0,H})\|_{P_{0}}\cdot\|\mathcal{T}_{0}(q_{n,H}-q_{0})\|_{P_{0}}=o_{p}(n^{-1/2}).$
The limiting variance of $\psi_{n,H}$ equals the variance of its influence function, $\operatorname{Var}(\varphi_{0,H}(S_{0},A_{0},Y_{0},S_{1}))$. In the nonparametric case where $H=L^{\infty}(\lambda)$, Theorem 1 recovers the nonparametric efficient influence function (EIF) for the policy value $E_{0}[V^{\pi}(q_{0})(S_{0})]$ derived in Kallus and Uehara (2020, 2022). In this setting, $\psi_{n,H}$ is asymptotically efficient. However, under semiparametric restrictions, the influence function $\varphi_{0,H}$ in Theorem 1 generally differs from the EIF of $\Psi:\mathcal{P}_{H}\rightarrow\mathbb{R}$. As a result, $\psi_{n,H}$ may be inefficient under the model $\mathcal{P}_{H}$, though necessarily more efficient than fully nonparametric estimators under semiparametric restrictions. We note that efficient estimation under $\mathcal{P}_{H}$ often requires inverse weighting by the conditional variance of the outcome (Laan and Robins, 2003). Thus, while potentially inefficient, $\psi_{n,H}$ is often more stable and simpler to implement—much like ordinary least squares, which remains widely used despite its inefficiency under heteroscedasticity.
The limiting variance of $\psi_{n,H}$, and its potential efficiency gains over nonparametric estimators, are primarily governed by the variability of $\mathcal{T}_{0}(\alpha_{0,H})$, which reflects the degree of policy overlap and intertemporal state overlap. In the nonparametric case and for the policy value estimand, $\mathcal{T}_{0}(\alpha_{0,H})$ corresponds to the state occupancy ratio $d_{0}$ defined in (2), which may exhibit high variance or fail to exist when overlap is limited (Mehrabi and Wager, 2024). Semiparametric restrictions on the $Q$-function generally reduce the variability of $\mathcal{T}_{0}(\alpha_{0,H})$, becoming less sensitive to limited overlap. Specifically, $\mathcal{T}_{0}(\alpha_{0,H})$ corresponds to the $L^{2}(P_{0,A_{0},S_{0}})$-projection of its nonparametric counterpart onto the range $\mathcal{T}_{0}(\overline{H}_{P_{0}})$, assuming the latter exists. Since projection contracts norms, the variance of $\varphi_{0,H}$ typically decreases as the function class $H$ becomes more restrictive. For example, if $H$ constrains $q_{0}$ to depend only on a subvector $\widetilde{S}_{t}\subset S_{t}$, then $\mathcal{T}_{0}(\alpha_{0,H})(A_{0},S_{0})$ reduces to the conditional expectation of $d_{0}(A_{0},S_{0})$ given $(A_{0},\widetilde{S}_{0})$, requiring overlap only within this lower-dimensional subspace. These efficiency gains can be substantial, as the following example illustrates in the context of data fusion.
Example 3 (Long-term causal effect in a data fusion setting).
Let $A_{t}:=Z$ be a time-homogeneous study assignment, and define the state as $S_{t}=(Z,\widetilde{S}_{t})$, as in Example 2, with $P_{0}(Z=1\mid S_{0})=p$. We consider a data fusion setting where experimental data ($Z=1$) is augmented with historical controls ($Z=0$), extending the approach of Kallus et al. (2018) and van der Laan et al. (2024d) to MDPs. The expected cumulative reward in the treatment arm, $\mathbb{E}_{0}[\sum_{t=0}^{\infty}\gamma^{t}Y_{t}\mid Z=1]$, is identified by $E_{0}[q_{0}^{\pi_{0}}(1,(1,\widetilde{S}_{0}))\mid Z=1]$, where $\pi_{0}(a\mid(z,\widetilde{s})):=\mathbbm{1}(a=z)$ is the behavior policy. Assume the historical control dataset is large, so the control $Q$-function $V^{\pi}_{\text{hist}}(\widetilde{s}):=q_{0}(0,(0,\widetilde{s}))$ is effectively known. Consider the offset model $H:=\{q:(a,(z,\widetilde{s}))\mapsto V^{\pi}_{\text{hist}}(\widetilde{s})+\kappa+\beta z;\kappa,\beta\in\mathbb{R}\}$, which assumes $V^{\pi}_{\text{hist}}$ is known and posits that the data-combination bias $q_{0}(1,(1,\widetilde{S}_{0}))-q_{0}(0,(0,\widetilde{S}_{0}))$ is constant in $\widetilde{S}_{0}$. Although $H$ is affine, Theorem 2 still applies, with $\alpha_{0,H}$ replaced by $\alpha_{0,T_{0}(H)}$, where $T_{0}(H):=H-\{V^{\pi}_{\text{hist}}\}$ is the tangent space at $P_{0}$. Under standard MDP conditions, the nuisance term $\mathcal{T}_{0}(\alpha_{0,T_{0}(H)})$ in the influence function is $(a,(z,\widetilde{s}))\mapsto\frac{z}{p}\cdot E_{0}[d_{0}(Z,\widetilde{S}_{0})\mid Z=1]$, where $d_{0}(z,s)=\frac{d\mathbb{P}_{0}(\widetilde{S}_{t}=s\mid Z=z)}{d\mathbb{P}_{0}(\widetilde{S}_{0}=s\mid Z=z)}$. Since $(1-\gamma)E_{0}[d_{0}(Z,\widetilde{S}_{0})\mid Z=z]=1$ almost surely, the influence function—and thus the asymptotic variance of $\psi_{n,H}$—is unaffected by the degree of intertemporal state overlap. ∎
3.3 Extension to model-robust inference and asymptotic theory
In practice, the $Q$-function $q_{0}$ may not lie exactly in the model $H\subset L^{\infty}(\lambda)$, but may be well-approximated by elements of $H$, with potential efficiency gains justifying misspecification bias (Crump et al., 2006; van der Laan et al., 2023). Alternatively, $H$ may serve as a working model, with interest focused on a linear functional of the best approximation to $q_{0}$, such as the best linear predictor when $H$ is linear (Whitney et al., 2020; Vansteelandt and Dukes, 2022; Chambaz et al., 2012; Chernozhukov et al., 2018b). To enable valid inference without assuming $P_{0}\in\mathcal{P}_{H}$, we define a projection parameter $\Psi_{H}:\mathcal{P}\rightarrow\mathbb{R}$ that extends $\Psi:\mathcal{P}_{H}\rightarrow\mathbb{R}$ to the full nonparametric model. That is, $\Psi_{H}$ agrees with $\Psi$ on the submodel $\mathcal{P}_{H}$, but remains well-defined and interpretable under misspecification (Buja et al., 2019). We show how the DRL estimator $\psi_{n,H}$ can be adjusted to target $\Psi_{H}$, enabling model-robust inference. We then establish the asymptotic linearity and efficiency of both $\psi_{n,H}$ and its bias-corrected version as estimators of $\Psi_{H}(P_{0})$ under correct and incorrect specification, respectively.
We consider inference on the projection estimand $\psi_{0,H}:=E_{0}[m(S_{0},A_{0},q_{0,H})]$, where $q_{0,H}$ is a projection of the $Q$-function onto the model $H$. For each $P\in\mathcal{P}$, we define the Bellman projection $q_{P,H}$ as a solution to
$q_{P,H}\in\arg\min_{q\in\overline{H}_{P}}E_{P}\left[\left\{Y_{0}-\mathcal{T}_{P}(q)(A_{0},S_{0})\right\}^{2}\right],$ (8)
which exists and is unique under the invertibility condition on $\mathcal{T}_{P}$ in Condition C1(ii). The resulting working parameter $\Psi_{H}:\mathcal{P}\to\mathbb{R}$, defined by $\Psi_{H}(P):=E_{P}[m(S_{0},A_{0},q_{P,H})]$, extends $\Psi$ to the nonparametric model $\mathcal{P}$ while preserving its value on $\mathcal{P}_{H}$. Moreover, $q_{P,H}$ satisfies the projected Bellman equation $\mathcal{T}_{P}(q_{P,H})=\mu_{P,H}$, where $\mu_{P,H}:=\arg\min_{\mu\in\mathcal{T}_{P}(\overline{H}_{P})}\|\mu_{P}-\mu\|_{P}$ is the $L^{2}(P)$-projection of the outcome regression $\mu_{P}$ onto the range $\mathcal{T}_{P}(\overline{H}_{P})$. Hence, $q_{P,H}$ is the best approximation to $q_{P}$ within $H$ under the norm $\|\mathcal{T}_{P}(\cdot)\|_{P}$.
The following theorem extends the pathwise differentiability result for $\Psi$ in Theorem 1 to the projection parameter $\Psi_{H}$. In the following theorem, for each distribution $P\in\mathcal{P}$, we define the influence function $\varphi^{*}_{P,H}$ as the map:
$(s,a,y,s^{\prime})$ $\mapsto\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y+\gamma V^{\pi}(q_{P,H})(s^{\prime})-q_{P,H}(a,s)\}$
+$\quad+\{\alpha_{P,H}(a,s)-\gamma V^{\pi}(\alpha_{P,H})(s^{\prime})-\mathcal{T}_{P}(\alpha_{P,H})(a,s)\}\{\mu_{0}(a,s)-\mathcal{T}_{P}(q_{P,H})(a,s)\}$
$\quad+m(s,a,q_{P,H})-\Psi_{H}(P).$
Theorem 2 (Pathwise differentiability and efficient influence function).
Suppose C1 holds. Then the parameter $\Psi_{H}:\mathcal{P}\rightarrow\mathbb{R}$ is pathwise differentiable at $P_{0}$, with efficient influence function $\varphi^{*}_{0,H}$. Moreover, for any $\overline{P}\in\mathcal{P}$ for which $\varphi^{*}_{\overline{P},H}$ exist, the parameter satisfies the expansion: $\Psi_{H}(\overline{P})-\Psi_{H}(P_{0})=-P_{0}\varphi^{*}_{\overline{P},H}+R_{H}^{*}(\overline{P},P_{0}),$ where
$R_{H}^{*}(\overline{P},P_{0})$ $:=\left\langle\mathcal{T}_{\overline{P}}(\alpha_{\overline{P},H})-\mathcal{T}_{0}(\alpha_{0,H}),\,\mathcal{T}_{0}(q_{0,H}-q_{\overline{P},H})\right\rangle_{P_{0}}$
$\quad+\left\langle(\mathcal{T}_{\overline{P}}-\mathcal{T}_{0})(\alpha_{\overline{P},H}),\,\mu_{0}-\mu_{\overline{P}}+\mathcal{T}_{\overline{P}}(q_{\overline{P},H})-\mathcal{T}_{0}(q_{0,H})\right\rangle_{P_{0}}.$
When $q_{0}\in H$, the EIF $\varphi^{*}_{0,H}$ in Theorem 2 reduces to the influence function $\varphi_{0,H}$ from Theorem 1 for the restricted parameter $\Psi\colon\mathcal{P}_{H}\rightarrow\mathbb{R}$. Under misspecification of $H$, $\varphi^{*}_{0,H}$ includes additional terms involving the approximation error $\mathcal{T}_{0}(q_{0}-q_{0,H})$ and the residual $(s,a,s^{\prime})\mapsto\alpha_{0,H}(a,s)-\gamma V^{\pi}(\alpha_{0,H})(s^{\prime})-\mathcal{T}_{0}(\alpha_{0,H})(a,s).$ The remainder in the von Mises expansion also includes additional terms involving the errors $\mathcal{T}_{P}-\mathcal{T}_{0}$ and $\mu_{P}-\mu_{0}$.
The model-robust EIF of Theorem 2 demonstrates that, to correct the bias of the DRL estimator $\psi_{n,H}$ for the projection estimand $\Psi_{H}(P_{0})$ under model misspecification, it suffices to add an additional bias correction term. In particular, a model-robust estimator of $\Psi_{H}(P_{0})$ is given by
$\psi_{n,H}^{*}$ $=\psi_{n,H}+\frac{1}{n}\sum_{i=1}^{n}\big{(}\mu_{n}-\widehat{\mathcal{T}}_{n}(q_{n,H})\big{)}(A_{0,i},S_{0,i})\left\{\alpha_{n,H}(A_{0,i},S_{0,i})-\gamma V^{\pi}(\alpha_{n,H})(S_{1,i})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H})(A_{0,i},S_{0,i})\right\}.$
where $q_{n,H}\in H$, $\alpha_{n,H}\in H$, $\widehat{\mathcal{T}}_{n}(q_{n,H})$, $\widehat{\mathcal{T}}_{n}(\alpha_{n,H})$, and $\mu_{n}$ are estimators of $q^{\pi}_{0,H}$, $\alpha_{0,H}$, $\mathcal{T}_{0}(q_{0,H})$, $\mathcal{T}_{0}(\alpha_{0,H})$, and $\mu_{0}$, respectively. The estimator $\psi_{n,H}^{*}$ generalizes the DRL estimator $\psi_{n,H}$ from the previous section, which is recovered under correct specification by setting $\mu_{n}:=\widehat{\mathcal{T}}_{n}(q_{n,H})$. Unlike $\psi_{n,H}$, however, the model-robust estimator $\psi_{n,H}^{*}$ enables efficient inference for $\Psi_{H}(P_{0})$ regardless of whether $H$ is correctly specified.
In the following conditions and theorem, let $\varphi_{n,H}^{*}$ denote the estimator of the EIF $\varphi_{0}^{*}$ from Theorem 2, obtained by plugging in our nuisance estimators.
  1. (C2)
    Consistency: $n^{-\frac{1}{2}}(P_{n}-P_{0})\{\varphi_{n,H}^{*}-\varphi_{0,H}^{*}\}=o_{p}(1)$.
  2. (C3)
    Nuisance estimation rate: Each of the following hold:
    1. (a)
      $\left\|\widehat{\mathcal{T}}_{n}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{0,H})\right\|_{P_{0}}\cdot\left\|\mathcal{T}_{0}(q_{n,H})-\mathcal{T}_{0}(q_{0,H})\right\|_{P_{0}}=o_{p}(n^{-\frac{1}{2}})$
    2. (b)
      $\left\|\mathcal{T}_{0}(\alpha_{n,H})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H})\right\|_{P_{0}}\cdot\left\|\widehat{\mathcal{T}}_{n}(q_{n,H})-\mathcal{T}_{0}(q_{0,H})+\mu_{0}-\mu_{n}\right\|_{P_{0}}=o_{p}(n^{-\frac{1}{2}})$
Theorem 3.
Assume C1, C2, and C3. Then, $\psi_{n,H}^{*}-\psi_{0,H}=(P_{n}-P_{0})\varphi_{0,H}^{*}+o_{p}(n^{-\frac{1}{2}})$. Moreover, $\psi_{n,H}^{*}$ is a $P_{0}$-regular and efficient estimator for the working parameter $\Psi_{H}$ under the nonparametric model .
Condition C2 is an empirical process condition that holds if $\|\varphi_{n,H}^{*}-\varphi_{0,H}^{*}\|_{P_{0}}=o_{p}(1)$ and the difference lies in a Donsker class, or if sample-splitting or cross-fitting is used to bypass Donsker conditions (van der Laan et al., 2011; Chernozhukov et al., 2018a). Condition C3 is a doubly robust rate condition requiring that the nuisance estimators converge at sufficiently fast rates. The first rate condition, $\|\widehat{\mathcal{T}}_{n}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{0,H})\|_{P_{0}}\cdot\|\mathcal{T}_{0}(q_{n,H})-\mathcal{T}_{0}(q_{0,H})\|_{P_{0}}=o_{p}(n^{-1/2})$, commonly appears in nonparametric inference for functionals of solutions to inverse problems (Kallus and Uehara, 2022; Bennett et al., 2023a; Li et al., 2024). Under C1, this condition holds if both $\widehat{\mathcal{T}}_{n}(\alpha_{n,H})$ and $q_{n,H}$ converge to their targets in $\|\cdot\|_{P_{0}}$-norm at rates faster than $n^{-1/4}$. Under correct specification ($q_{0}\in H$), the second rate condition is trivially satisfied when using the DRL estimator $\psi_{n,H}$, since $\mu_{n}=\widehat{\mathcal{T}}_{n}(q_{n,H})$ and $\mu_{0}=\mathcal{T}_{0}(q_{0,H})$. These conditions can be satisfied under appropriate smoothness assumptions using machine learning algorithms such as neural networks, random forests, and gradient-boosted trees.
3.4 Model misspecification bias and data-driven model selection
Selecting an appropriate working model $H$ for $q_{0}$ is challenging and may compromise inference due to misspecification bias. While choosing a rich model $H$ may mitigate such bias, it comes at the cost of increased variance and greater reliance on intertemporal overlap conditions. To navigate this bias-variance trade-off, it is natural to estimate the functional form of the $Q$-function $q_{0}$ from data. However, conventional wisdom holds that such data-driven choices can invalidate inference (Leeb and Pötscher, 2005). In this section, we propose an adaptive estimator (van der Laan et al., 2023) that learns a data-dependent model $H_{n}$ and establish conditions for valid inference using sieve methods and model selection. To do so, we show that misspecification of the $Q$-function induces only second-order bias in the target functional.
Let $H_{n}\subseteq H$ be a data-driven working model for the $Q$-function $q_{0}$. For example, $H_{n}$ may be selected via cross-validated FQI over a sieve of models $H_{1}\subset H_{2}\subset\cdots\subset H_{\infty}:=H$, where $H$ is correctly specified and contains $q_{0}$. Alternatively, $H_{n}$ could result from variable selection or a learned feature transformation (Pritz et al., 2021; Pavse and Hanna, 2024). Given estimators $q_{n,H_{n}}\in H_{n}$ and $\alpha_{n,H_{n}}\in H_{n}$ of $q_{0,H_{n}}$ and $\alpha_{0,H_{n}}$, our proposed adaptive DRL estimator of $\Psi(P_{0})$ takes the form of a one-step estimator that naively assumes $H_{n}$:
$\psi_{n,H_{n}}=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H_{n}})+\frac{1}{n}\sum_{i=1}^{n}\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})\left\{Y_{0,i}+\gamma V^{\pi}(q_{n,H_{n}})(S_{1,i})-q_{n,H_{n}}(A_{0,i},S_{0,i})\right\}.$
To analyze its asymptotic behavior, we assume $H_{n}$ approximates a fixed but unknown oracle submodel $H_{0}\subseteq H$ containing $q_{0}$. Provided the approximation error between $H_{n}$ and $H_{0}$ vanishes suitably, the Adaptive Debiased Machine Learning (ADML) framework of van der Laan et al. (2023) implies that $\psi_{n,H_{n}}$ is asymptotically equivalent to the oracle estimator $\psi_{n,H_{0}}$, which assumes knowledge of $H_{0}$. We formally prove this result in Appendix B.
Our novel contribution is the following theorem, which shows that the parameter approximation bias $\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ is second order in the model approximation error and thus asymptotically negligible under suitable conditions. This result extends related bounds for linear functionals of the outcome regression derived in van der Laan et al. (2023) to the MDP and inverse problem setting.
Theorem 4 (Second-order model approximation error).
Suppose that $q_{0}\in H_{0}$ for some oracle submodel $H_{0}\subseteq H$, depending on $P_{0}$. Assume C1 holds for both $H:=H_{n}$ and $H:=H_{0}$. Then, the oracle approximation error of the working model $H_{n}$ satisfies:
$\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ $=-\langle\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n,0}}),\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\rangle_{P_{0}}$
where $H_{n,0}:=H_{n}\oplus H_{0}$ is the direct sum linear model.
A consequence of Theorem 4 is that the prespecified DML estimators $\psi_{n,H}$ and $\psi_{n,H}^{*}$ from the previous sections are robust to mild misspecification of the model $H$. Hence, these estimators may still provide valid inference for $\psi_{0}$, provided that the bias due to misspecification is negligible relative to the standard error of the estimator. In Appendix B, we show that the adaptive DRL estimator satisfies $\psi_{n,H_{n}}=\psi_{n,H_{0}}+o_{p}(n^{-1/2})$, provided that $\|\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n,0}})\|\cdot\|\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\|_{P_{0}}=o_{p}(n^{-1/2})$, and that the data-dependent influence function $\varphi_{0,H_{n}}$ converges in probability to the oracle influence function $\varphi_{0,H_{0}}$. This latter condition imposes a mild requirement on the consistency and stability of the model selection procedure but can be avoided if the model is learned using data independent of that used to evaluate the estimator (van der Laan et al., 2023).
For the approximation error $\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ to vanish, both $\mathcal{T}_{0}(\alpha_{0,H_{n}})$ and $\mathcal{T}_{0}(\alpha_{0,H_{n,0}})$ must converge to one another in $L^{2}(P_{0})$, and $\mathcal{T}_{0}(q_{0,H_{n}})$ must converge to $\mathcal{T}_{0}(q_{0})$. This convergence requires that the learned model $H_{n}$ approximates both the true $Q$-function $q_{0}$ and the union model representer $\alpha_{0,H_{n,0}}$ with vanishing error in the $\|\mathcal{T}_{0}(\cdot)\|$ norm. In sieve-based model selection, the event $H_{n}\subseteq H_{0}$ typically holds with high probability, in which case $\mathcal{T}_{0}(\alpha_{0,H_{n,0}})=\mathcal{T}_{0}(\alpha_{0,H_{0}})$, and it suffices for $H_{n}$ to grow sufficiently fast. For general model selection procedures, convergence of $\mathcal{T}_{0}(\alpha_{0,H_{n}})$ to $\mathcal{T}_{0}(\alpha_{0,H_{n,0}})$ further requires that any directions in $H_{n,0}\cap H_{n}^{\perp}$ contribute negligibly to the union model representer $\alpha_{0,H_{n,0}}$.
To further clarify these conditions, suppose the working model $H_{n}:=H_{\phi_{n}}$ and the oracle model $H_{0}:=H_{\phi_{0}}$ are induced by feature transformations. For a transformation $\phi:\mathcal{A}\times\mathcal{Z}\times\mathcal{S}\rightarrow\mathbb{R}^{m}$, define $H_{\phi}:=\{f\circ\phi:f:\mathbb{R}^{m}\rightarrow\mathbb{R}\}$. The combined model $H_{n,0}$ is given by $H_{(\phi_{n},\phi_{0})}$, where $(\phi_{n},\phi_{0})$ denotes the feature map formed by stacking $\phi_{n}$ and $\phi_{0}$. Theorem 4 implies that the approximation bias vanishes if the nuisance functions derived from $\phi_{n}$ and $(\phi_{n},\phi_{0})$ converge to those derived from $\phi_{0}$. In Lemma 14 of Appendix H.1, we show that with features $X$ and outcome $Y$, the mean squared error between $E_{0}[Y\mid\phi_{0}(X)]$ and $E_{0}[Y\mid\phi_{n}(X),\mathcal{D}_{n}]$ is bounded by the feature approximation error $\int\|\phi_{n}(x)-\phi_{0}(x)\|_{\mathbb{R}^{m}}^{2}\,P_{0,X}(dx)$. A sufficient condition for this bound is that the map $(t_{1},t_{2})\mapsto E_{0}[Y\mid\phi_{n}(X)=t_{1},\phi_{0}(X)=t_{2},\mathcal{D}_{n}]$ is almost surely Lipschitz continuous.
We revisit this setup in Section 5, where we introduce an adaptive DRL estimator that leverages $Q$-function estimates as a data-adaptive form of dimension reduction.
4 Methods and challenges in nuisance estimation
In this section, we describe how the $Q$-function $q_{0,H}$ and the Riesz representer $\alpha_{0,H}$ can be estimated using flexible machine learning methods. The $Q$-function can be readily estimated using standard reinforcement learning methods. In contrast, estimating the Riesz representer is more challenging, as it involves solving a convex–concave min–max problem. Recognizing this challenge, the next section proposes a novel DRL estimator that achieves debiasing without requiring estimation of the Riesz representer.
A popular algorithm for estimation of $q_{0,H}$ in reinforcement learning is fitted Q-iteration (FQI) (Munos and Szepesvári, 2008), an iterative method for solving the Bellman integral equation (1). FQI is based on the observation that if $q_{0,H}$ were known, it could be estimated by regressing $Y_{0}+\gamma V^{\pi}(q_{0,H})(S_{1})$ on $(A_{0},S_{0})$, since rearranging (1) yields $E_{0}[Y_{0}+\gamma V^{\pi}(q_{0,H})(S_{1})\mid A_{0},S_{0}]=q_{0,H}(A_{0},S_{0})$ almost surely. Since $q_{0,H}$ is unknown, FQI initializes $q_{n,H}^{\pi,(0)}:=0$ and iteratively updates it by regressing the Bellman outcome $Y_{0}+\gamma V^{\pi}(q_{n,H}^{\pi,(k)})(S_{1})$ on $(A_{0},S_{0})$ over the model class $H$ at each iteration $k+1\in\mathbb{N}$. Iteration stops when the $\ell^{2}$ norm between consecutive updates is sufficiently small or when out-of-sample or cross-validated risk ceases to improve. Algorithm 1 details the procedure. Theoretical guarantees on validity and convergence rates of FQI using generic function approximation methods, such as neural networks (Bishop, 1994), random forests (Breiman, 2001), and gradient boosted trees (Friedman, 2001), are provided in Munos and Szepesvári (2008) and Agarwal et al. (2019).
Algorithm 1 Fitted Q-Iteration
0:  Data, Function class $H$, number of iterations $K$;
1:  Initialize $q_{n,H}^{\pi,(0)}:=0$;
2:  for $k=0,1,\ldots,K-1$ do
[…]
4:     Update $q_{n,H}^{\pi,(k+1)}\in H$ by estimating:
$\arg\min_{q\in H}E_{0}\left[\{Y_{0}+\gamma V_{n,H}^{\pi,(k)}(S_{1})-q(A_{0},S_{0})\}^{2}\right];$
5:  end for
6:  Set $q_{n,H}:=q_{n,H}^{\pi,(K)}$;
7:  return  $q_{n,H}$;
One commonly used approach to estimate the Riesz representer $\alpha_{0,H}$ is to recast the minimization objective in (5) as the following convex-concave min-max optimization problem (Liu et al., 2018; Uehara et al., 2020; Dikkala et al., 2020; Kallus and Uehara, 2020):
$\alpha_{0,H}=\operatorname*{argmin}_{\alpha\in H}\max_{f\in L^{2}(P_{S_{0},A})}L_{0}(\alpha,f),$ (9)
where the objective function is given by:
$L_{0}(\alpha,f)=E_{0}\Big{[}\{\alpha(A_{0},S_{0})\}^{2}-$ $2\gamma\alpha(A_{0},S_{0})V^{\pi}(\alpha)(S_{1})-2m(S_{0},A_{0},\alpha)$
$-\frac{\gamma^{2}}{2}\left[\{f(A_{0},S_{0})\}^{2}-2V^{\pi}(\alpha)(S_{1})f(A_{0},S_{0})\right]\Big{]}.$
Given an estimator $\alpha_{n,H}$ of $\alpha_{0,H}$, the remaining nuisance component in $\mathcal{T}_{0}(\alpha_{0,H})$ can be estimated by regressing $V^{\pi}(\alpha_{n,H})(S_{1})$ on $(A_{0},S_{0})$, or equivalently, by computing $\arg\max_{f}L_{0}(\alpha_{n,H},f)$. Minimax estimation of solutions to conditional moment restrictions using empirical risk minimization techniques has been studied in Dikkala et al. (2020) and Bennett et al. (2023a).
To mitigate the computational challenges of min–max optimization, a common approach is to replace the inner unconstrained maximization with a constrained maximization over a linear model or a reproducing kernel Hilbert space (RKHS). In this case, the inner maximization can be efficiently computed in closed form using methods such as kernel ridge regression (Kallus and Uehara, 2020). The min–max optimization then reduces to a standard minimization problem based on the profiled loss (Murphy and Van der Vaart, 2000). A limitation of this approach is that constraining the inner maximization to a linear model or RKHS, while computationally attractive, can introduce substantial bias if the model is misspecified.
5 Automatic debiasing via Bellman Calibration
5.1 Leveraging the Q-function for dimension reduction
A central challenge in DRL is estimating the Riesz representer $\mathcal{T}_{0}(\alpha_{0,H})$ of the linear functional, such as the density ratio $d_{0}$ in (2). This quantity must typically be estimated by solving the min–max problem in (9), which is both computationally intensive and potentially unstable. To address this issue, we propose a specific instance of the adaptive DRL estimator developed in Section 3.4, which leverages a data-adaptive dimension reduction based on the $Q$-function. This approach only requires estimation of the $Q$-function, which can be readily obtained via fitted $Q$-iteration, and avoids estimating the Riesz representer altogether. In addition to simplifying estimation, it retains the efficiency benefits of semiparametric DRL while achieving the robustness guarantees of nonparametric DRL.
Our key insight is that the $Q$-function $q_{0}$ provides a sufficient one-dimensional summary for the Bellman equation, which enables us to reduce the conditional moment condition to
$q_{0}(A_{0},S_{0})-\gamma\,E_{0}[V^{\pi}(q_{0})(S_{1})\mid q_{0}(A_{0},S_{0})]=E_{0}[Y_{0}\mid q_{0}(A_{0},S_{0})],$
so that the equation conditions on the scalar $q_{0}(A_{0},S_{0})$ rather than the full tuple $(A_{0},S_{0})$. This identity follows from the law of total expectation, since
$q_{0}(A_{0},S_{0})=E_{0}\left[E_{0}\left[Y_{0}+\gamma V^{\pi}(q_{0})(S_{1})\mid A_{0},S_{0}\right]\mid q_{0}(A_{0},S_{0})\right]=E_{0}\left[Y_{0}+\gamma V^{\pi}(q_{0})(S_{1})\mid q_{0}(A_{0},S_{0})\right].$
This dimension reduction naturally motivates inference for an oracle projection parameter defined in terms of $q_{0}$. Specifically, if $q_{0}$ were known a priori, we could target the parameter $\Psi_{q_{0}}\colon P\mapsto E_{P}[m(S_{0},A_{0},q_{P,q_{0}})]$, where, for each $q\in L^{\infty}(\lambda)$, we define the Bellman projection
$q_{P,q}:=\operatorname*{argmin}_{f\circ q;\ f\colon\mathbb{R}\rightarrow\mathbb{R}}E_{P}\left[\left\{Y_{0}-\mathcal{T}_{P,q}(f\circ q)(A_{0},S_{0})\right\}^{2}\right],$
[…]
We now introduce notation used in the theorem. For a feature map $\phi:\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}^{m}$, define the Riesz representer $\alpha_{0,\phi}:=\operatorname*{argmin}_{f\in\overline{H}_{P_{0},\phi}}E_{0}\left[\left\{\mathcal{T}_{0,\phi}(f\circ\phi)(A_{0},S_{0})\right\}^{2}-2m(S_{0},A_{0},f\circ\phi)\right]$, where $\overline{H}_{P_{0},\phi}$ is the closure of the model class $H_{\phi}:=\{f\circ\phi\text{ with }f:\mathbb{R}^{m}\rightarrow\mathbb{R}\}\cap L^{\infty}(\lambda)$ with respect to $\|\cdot\|_{P_{0}}$. Let $\widetilde{\alpha}_{0,q_{n}}:=\operatorname*{argmin}_{f\in\overline{H}_{P_{0},q_{n}}}\left\|\mathcal{T}_{0,(q_{n},q_{0})}(\alpha_{0,(q_{n},q_{0})})-\mathcal{T}_{0,q_{n}}(f\circ q_{n})\right\|_{P_{0}},$ where $\mathcal{T}_{0,q_{n}}(\widetilde{\alpha}_{0,q_{n}})$ is the projection of $\mathcal{T}_{0,(q_{n},q_{0})}(\alpha_{0,(q_{n},q_{0})})$ onto the range $\mathcal{T}_{0,q_{n}}(\overline{H}_{q_{n}})$.
  1. (D1)
    (Continuity and invertibility) For $\phi=(q_{0}),$ $(q_{n})$, and $(q_{0},q_{n})$, the following hold:
    1. (a)
      There exists $C<\infty$ such that $|E_{0}[m(S_{0},A_{0},q)]|\leq C\|q\|_{P_{0}}$ for all $q\in H_{\phi}$.
    2. (b)
      $\mathcal{T}_{0,\phi}:\overline{H}_{P_{0},\phi}\rightarrow\overline{H}_{P_{0},\phi}$ is continuous and invertible.
In the following condition, we define the random variable $D_{n,0}:=\mathcal{T}_{0,(q_{n},q_{0})}\left(\alpha_{0,(q_{n},q_{0})}\right)(A_{0},S_{0})$.
  1. (D2)
    (Lipschitz continuity) There exists a constant $L\in(0,\infty)$ such that, for all sufficiently large $n$, the bivariate function $(t_{1},t_{2})\mapsto E_{0}\left[D_{n,0}\,\middle|\,q_{n}(A_{0},S_{0})=t_{1},\ q_{0}(A_{0},S_{0})=t_{2},\ \mathcal{D}_{n}\right]$ is almost surely Lipschitz continuous with Lipschitz constant $L$.
Theorem 5 (Parameter approximation error is second-order).
Suppose that D1 holds. Then,
$\Psi_{q_{n}}(P_{0})-\Psi(P_{0})$ $=\langle\mathcal{T}_{0,(q_{n},q_{0})}(\alpha_{0,(q_{n},q_{0})})-\mathcal{T}_{0,q_{n}}(\widetilde{\alpha}_{0,q_{n}}),\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}}-q_{0})\rangle_{P_{0}}.$
If D2 also holds, then $\Psi_{q_{n}}(P_{0})-\Psi(P_{0})=O_{p}\left(\|q_{n}-q_{0}\|_{P_{0}}\|\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\|_{P_{0}}\right).$
Condition D1 is a variant of C1 for the dimension-reduced Bellman operator. Theorem 5 states that the parameter approximation bias vanishes if the optimal transformation of the estimated $Q$-function $q_{n}$ converges to $q_{0}$, and the difference between $\mathcal{T}_{0,(q_{n},q_{0})}(\alpha_{0,(q_{n},q_{0})})$—a function of $(q_{n}(A_{0},S_{0}),q_{0}(A_{0},S_{0}))$—and its projection onto functions of $q_{n}(A_{0},S_{0})$ also vanishes in $\|\cdot\|_{P_{0}}$. Heuristically, this requires that conditioning on $(q_{n}(A_{0},S_{0}),q_{0}(A_{0},S_{0}))$ asymptotically provides the same information as conditioning on either component alone. We formalize this by imposing D2, a mild regularity requirement that a certain bivariate function is Lipschitz continuous. Related smoothness conditions for debiased machine learning with estimated features in regression have been studied in Benkeser et al. (2017, 2020); Wang et al. (2023b); Bonvini et al. (2024); van der Laan et al. (2024b).
5.2 Proposed calibrated DRL estimator
In this section, we propose a DRL estimator of the oracle parameter $\Psi_{q_{0}}(P_{0})$, which can be interpreted as an adaptive estimator of $\Psi(P_{0})$ in the sense of Section 3.4. Instead of using an influence-function-based bias correction, we introduce a novel debiased plug-in estimator inspired by calibrated DML (van der Laan et al., 2024b). Our approach leverages a new form of calibration—Bellman calibration—which automatically corrects bias for any continuous linear functional, without requiring explicit estimation of the Riesz representer.
Our proposed Bellman-calibrated estimator of $\Psi(P_{0})$ is given by the plug-in estimator
$\psi_{n}^{*}:=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n}^{*}),$
where $q_{n}^{*}$ is a Bellman-calibrated estimator of the $Q$-function $q_{0}$ constructed to satisfy the empirical Bellman equation:
$q_{n}^{*}(a,s)=\frac{\sum_{i=1}^{n}1(q_{n}^{*}(A_{0,i},S_{0,i})=q_{n}^{*}(a,s))\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})\}}{\sum_{i=1}^{n}1(q_{n}^{*}(A_{0,i},S_{0,i})=q_{n}^{*}(a,s))}.$ (10)
As one approach to constructing a Bellman-calibrated estimator, we propose isotonic Bellman calibration, which is outlined in Algorithm 2 and discussed in more detail below. Bellman calibration implies that the Bellman residuals $\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}^{*}(A_{0,i},S_{0,i})\}_{i=1}^{n}$ are empirically orthogonal to the transformed $Q$-function estimates $\{f(q_{n}^{*}(A_{0,i},S_{0,i}))\}_{i=1}^{n}$ for each $f:\mathbb{R}\rightarrow\mathbb{R}$. A consequence of this property is the following lemma.
Lemma 6 (Calibration corrects plug-in bias).
Suppose $q_{n}^{*}$ is Bellman calibrated (10). Then,
$\psi_{n}^{*}=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n}^{*})+\frac{1}{n}\sum_{i=1}^{n}\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})(A_{0,i},S_{0,i})\left\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}^{*}(A_{0,i},S_{0,i})\right\}.$
A consequence of Lemma 6 is that the Bellman-calibrated estimator $\psi_{n}^{*}$ is implicitly debiased for the data-dependent parameter $\Psi_{q_{n}^{*}}$. In particular, it equals a DRL estimator of $\Psi_{q_{n}^{*}}$ that uses $q_{n}^{*}$ as an estimate of $q_{0}$ and the true Riesz representer nuisance $\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})$. By Theorem 5, it is also debiased for the oracle parameter $\Psi_{q_{0}}$, up to second-order terms. Consequently, $\psi_{n}^{*}$ is an example of an adaptive DRL estimator, as described in Section 3.4, corresponding to the working model $H_{n}=\{f\circ q_{n}^{*}:f\}$ and the oracle model $H_{0}=\{f\circ q_{0}:f\}$. The debiasing achieved via calibration is agnostic to the specific functional and does not require knowledge of $\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})$.
Algorithm 2 Isotonic Bellman Calibration
0:  Initial estimator $q_{n}$ of $q_{0}$, stopping threshold $\varepsilon\approx 0$;
1:  initialize $q_{n}^{*(0)}:=q_{n}$;
2:  for $k=0,1,2,\ldots$ do
3:     compute $f_{n}^{(k+1)}$ by solving:
$\operatorname*{argmin}_{f\in\mathcal{F}_{iso}}\sum_{i=1}^{n}\{Y_{0,i}+\gamma V_{q_{n}^{*(k)}}^{\pi}(S_{1,i})-f(q_{n}(A_{0,i},S_{0,i}))\}^{2};$
4:     update $q_{n}^{*(k+1)}:=f_{n}^{(k+1)}\circ q_{n}$;
5:     if $\|q_{n}^{*(k+1)}-q_{n}^{*(k)}\|_{P_{n}}<\varepsilon$ then
6:        set $q_{n}^{*}:=q_{n}^{*(k+1)}$;
7:        break;
8:     end if
9:  end for
10:  return  $q_{n}^{*}$;
Our novel calibration algorithm, Isotonic Bellman Calibration, is outlined in Algorithm 2, where $\mathcal{F}_{\text{iso}}$ denotes the space of all monotone non-decreasing (isotonic) functions. Isotonic Bellman calibration combines isotonic regression—a distribution-free calibration method widely used in prediction (Niculescu-Mizil and Caruana, 2005; Van Der Laan et al., 2023; van der Laan and Alaa, 2025)—with fitted $Q$-iteration for solving the Bellman equation (Munos and Szepesvári, 2008). Since isotonic regression solutions are generally non-unique, we follow Groeneboom and Lopuhaa (1993) and select the unique càdlàg, piecewise constant solution with jumps only at observed values of $q_{n}$. When paired with fitted $Q$-iteration, isotonic regression acts as a data-driven histogram estimator that bins the one-dimensional space $\{q_{n}(a,s):a\in\mathcal{A},\,s\in\mathcal{S}\}$. Upon convergence, the calibrated estimator $q_{n}^{*}(a,s)$ is evaluated as the empirical mean of the Bellman outcome $Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})$ over observations where $q_{n}(A_{0,i},S_{0,i})$ falls into the same bin as $q_{n}(a,s)$, thereby ensuring empirical calibration as defined in (10).
Since calibration involves additional fitting, it is important that the initial Q-function estimator $q_{n}$ in Algorithm 2 is obtained from a dataset independent of the data used for calibration. In practice, this can be achieved via sample splitting, wherein one half of the data is used to estimate $q_{n}$, and the other half is used to calibrate $q_{n}$ and compute the plug-in estimator of the linear functional. To improve data efficiency, cross-fitting techniques for nuisance estimation can be employed (van der Laan et al., 2011; Chernozhukov et al., 2018a). We provide a cross-fitted variant of Algorithm 2 in Appendix C.
5.3 Asymptotic theory
In this section, we show that the estimator $\psi_{n}^{*}$ is asymptotically linear and superefficient for $\Psi$, while achieving nonparametric efficiency for the oracle parameter $\Psi_{q_{0}}$.
Our main result of this section is the following theorem. We discuss its conditions in Appendix C.4. We introduce the following notation. Let $\varphi_{n,q_{n}^{*}}$ denote the map $(s,a,y,s^{\prime})\mapsto m(s,a,q_{n}^{*})+\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})(a,s)\{y+\gamma q_{n}^{*}(a,s^{\prime})-q_{n}^{*}(a,s)\}-\psi_{n}^{*}$, and let $\varphi_{0,q_{0}}$ denote $(s,a,y,s^{\prime})\mapsto m(s,a,q_{0})+\mathcal{T}_{0,q_{0}}(\alpha_{0,q_{0}})(a,s)\{y+\gamma V^{\pi}(q_{0})(s^{\prime})-q_{0}(a,s)\}-\Psi_{q_{0}}(P_{0})$.
  1. (D3)
    Nuisance estimation rate: $\|q_{n}^{*}-q_{0}\|_{P_{0}}\|\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0})\|_{P_{0}}=o_{p}(n^{-\frac{1}{2}})$.
  2. (D4)
    Empirical process condition: $n^{-\frac{1}{2}}(P_{n}-P_{0})\{\varphi_{n,q_{n}^{*}}-\varphi_{0,q_{0}}\}=o_{p}(1)$.
Theorem 7 (Asymptotic linearity and superefficiency).
Suppose $q_{n}^{*}$ satisfies the empirical Bellman calibration condition (10), and D1D4 hold. Then:
  1. (i)
    $\psi_{n}^{*}$ is asymptotically linear for $\psi_{0}$ with influence function $\varphi_{0,q_{0}}$.
If, in addition, D1 holds in a Hellinger neighborhood of $P_{0}$, then $\varphi_{0,q_{0}}$ is the EIF for $\Psi_{q_{0}}$, and:
  1. (ii)
    $\psi_{n}^{*}$ is a $P_{0}$-regular and efficient estimator for $\Psi_{q_{0}}$ under the nonparametric model.
As a consequence of Theorem 7, the calibrated estimator $\psi_{n}^{*}$ satisfies the asymptotic expansion $\psi_{n}^{*}-\psi_{0}=(P_{n}-P_{0})\varphi_{0,q_{0}}+o_{p}(n^{-1/2})$, where $\varphi_{0,q_{0}}$ is the EIF for the oracle parameter $\Psi_{q_{0}}$. Therefore, $\sqrt{n}(\psi_{n}^{*}-\Psi(P_{0}))$ is asymptotically normal with limiting variance $\sigma_{0}^{2}:=\mathrm{Var}_{0}(\varphi_{0,q_{0}})$. Given a consistent estimator of $\sigma_{0}^{2}$, inference can be conducted using Wald-type confidence intervals and tests. Under regularity conditions, this variance can be consistently estimated by computing the empirical variance of an estimate of the influence function $\varphi_{0,q_{n}^{*}}$, which requires estimating the unknown quantity $\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})$. The function $\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})$ is known up to a one-dimensional transformation of $q_{n}^{*}$, and, by properties of isotonic regression, the calibrated estimator $q_{n}^{*}$ takes on finitely many values. Consequently, $\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})$ can be computed efficiently using matrix formulas for discrete Markov chains (see Appendix C.2). Alternatively, $\sigma_{0}^{2}$ can be estimated via a bootstrap procedure that resamples the calibration step while holding the initial $Q$-function estimator fixed, following van der Laan et al. (2024b), thereby avoiding additional nuisance estimation altogether.
The limiting variance of $\psi_{n}^{*}$ is equal to the efficiency bound for the oracle parameter $\Psi_{q_{0}}$ under $P_{0}$. Consequently, $\psi_{n}^{*}$ is generally superefficient for the original parameter $\Psi$, as its limiting variance may be strictly below the generalized Cramér–Rao efficiency bound for $\Psi$ under the nonparametric model. The efficiency gains of $\psi_{n}^{*}$ relative to nonparametric estimators can be substantial.In particular, the key quantity $\mathcal{T}_{0,q_{0}}(\alpha_{0,q_{0}})$ appearing in the limiting variance depends on the degree of intertemporal state overlap within the level sets of the $Q$-function, rather than across the entire state space. Specifically for the policy value estimand, the relevant term $\mathcal{T}_{0}(\alpha_{0,q_{0}})$ in the influence function $\varphi_{0,q_{0}}$ reduces to the aggregated density ratio $E_{0}[d_{0}(A_{0},S_{0})\mid q_{0}(A_{0},S_{0})]$, where $d_{0}$ denotes the state occupancy ratio defined in (2).
The superefficiency of $\psi_{n}^{*}$ comes at the cost of irregularity: it is not a regular estimator for the target parameter $\Psi$ and may exhibit non-vanishing asymptotic bias under sampling from local alternatives to $P_{0}$ in the nonparametric model (van der Laan et al., 2023, Theorem 7). Nevertheless, Theorem 7 guarantees that $\psi_{n}^{*}$ is regular and efficient for the oracle parameter $\Psi_{q_{0}}$ at $P_{0}$, and locally asymptotically equivalent to the semiparametric DRL estimator that assumes the oracle submodel $H_{q_{0}}$ a priori. Thus, even under a local alternative $P_{0,n^{-1/2}}$, the estimator $\psi_{n}^{*}$ enables valid inference for the projection $\Psi_{q_{0}}(P_{0,n^{-1/2}})$. In contrast, semiparametric DRL estimators are also irregular for $\Psi$ under the nonparametric model but lack adaptivity and may fail to achieve $\sqrt{n}$-consistency when the assumed model is misspecified. For additional discussion of this trade-off between superefficiency and irregularity, see van der Laan et al. (2023).
5.4 Related work
In the special case of a static setting ($\gamma=0$) and the average treatment effect (ATE) functional, our estimator is asymptotically equivalent to the targeted minimum loss estimator (TMLE) proposed by Benkeser et al. (2020). That approach uses an estimate of the outcome regression as a data-adaptive dimension reduction to construct a superefficient estimator. We extend this idea to the MDP setting, replacing the outcome regression with the $Q$-function. Unlike their explicit TMLE update, we achieve debiasing via isotonic calibration, avoiding additional nuisance estimation. Our method is also inspired by the calibrated DML framework of van der Laan et al. (2024c), which shows how nuisance calibration yields higher-order debiased estimators for linear functionals of the outcome regression. The objective of that paper was to construct estimators that remain asymptotically normal even when one of the nuisance functions is inconsistently estimated. While we generalize this idea to the MDP setting, our goal is distinct: we use calibration to construct a superefficient plug-in estimator without estimating the Riesz representer. In particular, whereas van der Laan et al. (2024c) calibrate both nuisances for nonparametric inference, we calibrate a single nuisance to avoid the complexity of high-dimensional min-max estimation.
6 Numerical experiments
We consider Example 2 described in Section 2.2, where the parameter $\Psi$ represents the long-term causal effect of an A/B test. Participants are randomly assigned to either a treatment group, receiving a specific intervention, or a control group, receiving an alternative intervention or no intervention. We let $A_{t}$ denote the study assignment at time $t$, where $A_{t}=Z$ almost surely, and consider the behavior policy $\pi$ that sets the treatment $A_{t}$ equal to $Z$. We write the state at time $t$ as $S_{t}=(Z,\widetilde{S}_{t})$, and write a generic realization of $S_{t}$ as $s=(z,\widetilde{s})$. In this case, the $Q$-function $q_{0}^{\pi}$ equals the V-function $(\widetilde{s},z)\mapsto\mathbb{E}_{0}\left[\sum_{t=0}^{\infty}\gamma^{t}Y_{t}\mid\widetilde{S}_{0}=\widetilde{s},Z=z\right]$. Following Tran et al. (2023), our parameter of interest is the long-term ATE $E_{0}[V^{\pi}(q_{0})((1,\widetilde{S}_{0}))-V^{\pi}(q_{0})((0,\widetilde{S}_{0}))]$.
To mimic an A/B test on online platforms, we simulate a discrete-state Markov process with four state variables: engagement, churn risk, tenure, and an overlap variable parameterized by $\beta\in\mathbb{R}$. Each variable takes values in $\{0,1,2\}$. The degree of intertemporal state overlap is governed by the parameter $\beta$, with higher values of $\beta$ inducing less overlap between treatment arms and states over time. This setup allows us to assess estimator performance under varying degrees of intertemporal overlap—a key challenge in long-term causal inference. The simulation details are provided in Appendix D.2.
We estimate the $V$-function using fitted $Q$-iteration with gradient-boosted regression trees implemented in lightgbm. We compare three estimators: the Bellman-calibrated plug-in estimator from Section 5, an adaptive DRL estimator that learns a data-driven model for the $V$-function, and the nonparametric DRL estimator of Kallus and Uehara (2020) and Tran et al. (2023). The adaptive DRL approach, described in Section 3.4, constructs the model class $\mathcal{H}_{n}$ by one-hot encoding the leaf nodes of a gradient-boosted sum-of-trees model for the $V$-function (see, e.g, Section 3.1 of He et al. (2014)). The Riesz representer is estimated via min-max optimization using gradient-boosted trees; the inner maximization is approximated in closed form via ridge regression on the induced tree features. Confidence intervals for the Bellman-calibrated estimator are obtained by bootstrap, following van der Laan et al. (2024c).
(a) Limited overlap ($\beta=0.6$).
(b) Varying overlap ($\gamma=0.8$).
Figure 2: Bias, standard error (SE), and coverage across discount factors $\gamma$ for setting with limited intertemporal overlap. Subfigure (d) compares Bellman calibration and nonparametric methods in low-overlap settings ($\beta=0.7,0.8,0.9$); adaptive DRL (tree) results closely resemble the nonparametric method and are omitted for clarity.
Figure 2 summarizes the performance of the three estimators across settings with varying intertemporal overlap ($\beta$) and discount factors ($\gamma$). Appendix D.2 provides results for the good ($\beta=0$) and moderate ($\beta=0.3$) overlap settings. Estimator performance depends critically on the degree of overlap: with good overlap, all methods exhibit low bias, small standard errors, and near-nominal coverage. As overlap deteriorates ($\beta=0.3$, $0.6$), the nonparametric estimator performs poorly—showing high variance and undercoverage—reflecting its sensitivity to inverse weighting. Subfigure (d) compares the nonparametric and Bellman-calibrated estimators under low overlap at $\gamma=0.8$. The nonparametric method exhibits greater bias, variance, and coverage error, while Bellman calibration yields substantially better coverage that approaches 95% with increasing sample size. Across all settings, Bellman calibration is the most stable, consistently achieving low bias and variance. While the tree-based adaptive DRL estimator improves over the nonparametric baseline, it remains more variable than Bellman calibration under poor overlap, likely due to the higher complexity of $\mathcal{H}_{n}$.
The discount factor $\gamma$ determines how far into the future the estimator must extrapolate to evaluate long-term causal effects. As $\gamma$ increases, estimation becomes more difficult, with greater bias and standard error observed near $\gamma=0.75$. The nonparametric estimator becomes increasingly unstable at higher $\gamma$, reflecting its sensitivity to limited intertemporal overlap. In contrast, Bellman calibration maintains low bias and variance across all values of $\gamma$, underscoring the benefits of calibration and dimension reduction in overcoming overlap limitations.
7 Conclusion
In this work, we leveraged $Q$-function calibration to develop calibrated plug-in estimators. A promising direction for future research is to extend the calibrated debiased machine learning framework of van der Laan et al. (2024c) to linear functionals of solutions to integral equations. This extension could demonstrate that calibrating both the $Q$-function and the Riesz representer yields doubly robust asymptotically linear estimators (Benkeser et al., 2017), enabling valid inference, including confidence intervals and hypothesis tests, even if either component is estimated inconsistently or at a slow rate. We leave the development of such a doubly robust inference procedure for future work.
Appendix A Sufficient conditions for C1
In this section, we provide sufficient conditions for the invertibility of $\mathcal{T}_{P}$ in Condition C1. The key idea is that $\mathcal{T}_{P}$ differs from the identity operator by a compact operator under mild conditions. As a result, $\mathcal{T}_{P}$ is a Fredholm operator of index zero on $\overline{H}_{P}$, and the inverse problem defining $q_{P,H}$ is a Fredholm equation of the second kind (Conway, 1994).
In what follows, we define the operator $\mathcal{K}_{P}:L^{\infty}(\lambda)\rightarrow L^{\infty}(\lambda)$ pointwise as $\mathcal{K}_{P}(h)(a,s):=E_{P}\left[V^{\pi}(h)(S_{1})\mid A_{0}=a,S_{0}=s\right]$. A key property of the Bellman operator $\mathcal{T}_{P}:=I-\gamma\mathcal{K}_{P}$ is that it is a $\gamma$-perturbation of the identity operator $I:L^{\infty}(\lambda)\rightarrow L^{\infty}(\lambda)$, which maps each function to itself, by the conditional expectation operator $\mathcal{K}_{P}$. The following condition ensures existence of $q_{P,H}$.
  1. (A1)
    Compactness and Fredholm property: The operator $\mathcal{K}_{P}$ is continuous on $H$ with respect to $\|\cdot\|_{P}$, and its unique continuous extension is compact on the closure $(\overline{H}_{P},\|\cdot\|_{P})$.
Condition A1 is a mild requirement that holds under appropriate assumptions on the state transition probabilities. Suppose that the operator $\mathcal{K}_{P}$ admits the following integral representation for $h\in L^{\infty}(\lambda)$:
[…]
where the kernel is given by $K_{P}(a^{\prime},s^{\prime}\mid a,s):=\frac{\pi(a^{\prime}\mid s^{\prime})}{b_{P}(a^{\prime}\mid s^{\prime})}\,\frac{dP(S_{1}=s^{\prime}\mid A_{0}=a,S_{0}=s)}{dP(S_{0}=s^{\prime})}.$ Condition A1 holds with $H=L^{\infty}(\lambda)$ when the state and action spaces are compact subsets of $\mathbb{R}^{d}$, and the kernel function $K_{P}$ is continuous and bounded. In this case, the kernel is square-integrable with respect to the product measure $P_{A_{0},S_{0}}\otimes P_{A_{0},S_{0}}$, so $\mathcal{K}_{P}$ defines a Hilbert–Schmidt operator on $L^{2}(P_{A_{0},S_{0}})$, and is therefore compact (Conway, 1994). Related conditions for the nonparametric well-posedness of the integral equation for the $Q$-function were proposed in Section 3.2 of Chen and Qi (2022). Notably, Condition A1 becomes even less stringent under stronger semiparametric restrictions imposed through $H$, and holds trivially when $H$ is finite-dimensional.
The following theorem shows that Condition A1 ensures not only the existence but also the uniqueness of the Bellman projection for almost all discount factors $\gamma$. We begin by introducing the following condition.
  1. (A2)
    Invertibility: $\gamma^{-1}$ is not an eigenvalue of $\mathcal{K}_{P}$ when restricted to $(\overline{H}_{P},\|\cdot\|_{P})$.
Theorem 8 (Existence and uniqueness of solution).
Assume A1 holds at $P\in\mathcal{P}$. Then, the range $\mathcal{T}_{P}(\overline{H}_{P})$ is a closed subspace of $L^{2}(P_{A_{0},S_{0}})$ and there exists an element $q_{P,H}\in H$ satisfying (8). Moreover, if A2 also holds, then $\mathcal{T}_{P}\colon(\overline{H}_{P},\|\cdot\|_{P})\rightarrow L^{2}(P_{A_{0},S_{0}})$ has a bounded inverse on its range, and $q_{P,H}=\mathcal{T}_{P}^{-1}(\mu_{P,H})$ is the unique solution.
Condition A1 ensures that the Bellman integral operator $\mathcal{T}_{P}$ can be continuously extended to a map from $\overline{H}_{P}$ to $\overline{H}_{P}$, and this map differs from the identity by a compact operator. Consequently, $\mathcal{T}_{P}=I-\gamma\mathcal{K}_{P}$ is a Fredholm operator of index zero on $\overline{H}_{P}$, and the inverse problem defining $q_{P,H}$ is a Fredholm equation of the second kind (Conway, 1994). The closedness of the range $\mathcal{T}_{0}(\overline{H}_{0})$, guaranteed by Theorem 8, ensures the existence of the Bellman projection $q_{P,H}$ in (8). Condition A2 further guarantees uniqueness, and holds whenever $\gamma<\|\mathcal{K}_{P}\|_{P}^{-1}$. By the compactness of $\mathcal{K}_{P}$, the spectrum of $\mathcal{T}_{P}$ is countable with $\gamma^{-1}$ as the only possible accumulation point. Hence, the Bellman projection $q_{P,H}$ is unique for almost all discount factors $\gamma$.
Historically, the existence and uniqueness of the $Q$-function $q_{0}$ for all $\gamma\in(0,1)$ in $L^{\infty}(\lambda)$ is established using Banach’s fixed point theorem, leveraging the fact that $\mathcal{K}_{P}$ is a contraction on $L^{\infty}(\lambda)$. However, this argument does not apply to the Bellman projection $q_{0,H}$, since the $L^{2}$-projection operator is not a contraction in the supremum norm. Moreover, even for $q_{0}$, this approach does not guarantee uniqueness of the solution in $L^{2}(P_{0,A_{0},S_{0}})$, as it only ensures injectivity of $\mathcal{T}_{P}$ on the dense subspace $L^{\infty}(\lambda)$.
Appendix B Model selection with adaptive debiased machine learning
B.1 General Approach
Selecting an appropriate working model is challenging and can compromise inference due to model misspecification bias. Adaptive debiased machine learning (ADML) (van der Laan et al., 2023) provides a unified framework that combines debiased estimation with data-driven model selection to construct superefficient estimators of smooth functionals, adapting to the structure of the nuisance components. By learning model assumptions or feature representations directly from data, ADML facilitates valid inference while mitigating misspecification bias. The calibrated plug-in estimator introduced in the previous section is a special case of ADML. In this section, we extend the ADML framework to the MDP setting, showing how semiparametric DRL can be combined with model selection to construct estimators that adapt to the functional form of the $Q$-function $q_{0}$, going beyond calibration alone. We refer to this extension as Adaptive DRL (ADRL).
Let $H_{n}\subseteq H$ be a data-dependent working model for the $Q$-function $q_{P}$, selected via model selection. ADRL posits the existence of a fixed but unknown oracle submodel $H_{0}\subseteq H$, determined by the true $Q$-function $q_{0}$, such that the approximation error between $H_{n}$ and $H_{0}$ vanishes asymptotically. Suppose we have estimators $q_{n,H_{n}}\in H_{n}$ and $\alpha_{n,H_{n}}\in H_{n}$ for $q_{0,H_{n}}$ and $\alpha_{0,H_{n}}$, respectively. The ADRL estimator of $\Psi(P_{0})$ is the DRL estimator:
$\psi_{n,H_{n}}=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H_{n}})+\frac{1}{n}\sum_{i=1}^{n}\widehat{\mathcal{T}}_{n}(\alpha_{n,H})\left\{Y_{0,i}+\gamma V^{\pi}(q_{n,H_{n}})(S_{1,i})-q_{n,H_{n}}(A_{0,i},S_{0,i})\right\},$
which targets the data-adaptive parameter $\Psi_{H_{n}}(P_{0})$. Unlike the model-robust estimator $\psi_{n,H}^{*}$ in Section 3.3, we omit the bias correction term, as it is unnecessary when the model approximation error vanishes asymptotically. Using a novel expansion of the approximation error $\Psi_{H_{n}}(P_{0})-\Psi_{H_{0}}(P_{0})$, we show that $\psi_{n,H_{n}}$ remains $\sqrt{n}$-consistent, asymptotically normal, and efficient for the oracle parameter $\Psi_{H_{0}}$. This oracle parameter coincides with the target $\Psi(P_{0})$ when $q_{0}\in H_{0}$, and often has a smaller efficiency bound, yielding less variable estimates and narrower confidence intervals while preserving unbiasedness. The adaptive plug-in estimator $\psi_{n}^{*}$ based on $Q$-function calibration is a special case corresponding to $H_{n}=\{f\circ q_{n}:f\}$ and $H_{0}=\{f\circ q_{0}:f\}$.
For example, the working model $H_{n}$ could be selected via cross-validated FQI over a sieve of models, i.e., a sequence of increasingly complex classes $H_{1}\subset H_{2}\subset H_{3}\subset\dots\subset H_{\infty}:=H$, where $H$ is a correctly specified model containing $q_{0}$. A plausible oracle submodel $H_{0}$ is the smallest correctly specified class in the sieve that contains $q_{0}$, which can feasibly be approximated via cross-validation. Alternatively, $H_{n}$ could result from a variable selection procedure or a learned feature transformation, with $H_{0}$ corresponding to the asymptotically selected variables or limiting transformation. Data-adaptive methods for learning state-action feature representations have been proposed in Pritz et al. (2021) and Pavse and Hanna (2024). Such transformations can also be derived directly from the fitted FQI model $q_{n,H_{n}}^{\pi}$, for example, by one-hot encoding the leaf nodes of trees in a random forest or gradient-boosted tree model, as in Section 3.1 of He et al. (2014).
Example 4 (Domain adaptation and confounding-robust data fusion).
Continuing from Example 3, suppose $S_{t}=(\widetilde{S}_{t},Z)$, where $Z=1$ denotes a randomized experiment with limited data and $Z=0$ denotes abundant but confounded observational data. To improve efficiency, we augment the unbiased experimental data with biased observational data. In Appendix E, we extend the ADML framework of van der Laan et al. (2024d), originally developed for cross-sectional data fusion, to the MDP setting. Following van der Laan et al. (2024d), we may use the Highly Adaptive Lasso (Benkeser and Van Der Laan, 2016) to learn a model class $\mathcal{H}_{n}$ that constrains the difference between the experimental and observational $Q$-functions, i.e., $q_{0}(a,s,1)-q_{0}(a,s,0)$. As shown in Example 3, these constraints enable confounding-robust information sharing across domains while avoiding misspecification by not imposing fixed structural assumptions. Crucially, without such constraints, no efficiency gain is possible from observational data. Learning them from data is therefore essential for improving efficiency while preserving nonparametric validity. See Appendix E for details.
B.2 Asymptotic theory
The following theorem is key to establishing the validity of our ADRL estimator, showing that the parameter approximation bias $\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ is second-order in the model approximation error and thus asymptotically negligible under certain conditions.
Theorem 9 (Second-order model approximation bias).
Suppose that $q_{0}\in H_{0}$ for some oracle submodel $H_{0}\subseteq H$, depending on $P_{0}$. Assume C1 holds for both $H:=H_{n}$ and $H:=H_{0}$. Then, the oracle approximation error of the working model $H_{n}$ satisfies:
$\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ $=-\langle\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n,0}}),\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\rangle_{P_{0}}$
where $H_{n,0}:=H_{n}\oplus H_{0}$ is the direct sum linear model.
For the approximation error $\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ to vanish, both $\mathcal{T}_{0}(\alpha_{0,H_{n}})$ and $\mathcal{T}_{0}(\alpha_{0,H_{n,0}})$ must converge in $L^{2}(P_{0})$, and $\mathcal{T}_{0}(q_{0,H_{n}})$ must converge to $\mathcal{T}_{0}(q_{0})$. This requires that the learned model $H_{n}$ approximates both the true $Q$-function $q_{0}$ and the union model representer $\alpha_{0,H_{n,0}}$ with vanishing error in the norm $\|\mathcal{T}_{0}(\cdot)\|$. In sieve-based model selection, the event $H_{n}\subseteq H_{0}$ typically holds with high probability, in which case $\mathcal{T}_{0}(\alpha_{0,H_{n,0}})=\mathcal{T}_{0}(\alpha_{0,H_{0}})$, and the condition reduces to requiring that $H_{n}$ grows sufficiently fast. For general model selection procedures, convergence of $\mathcal{T}_{0}(\alpha_{0,H_{n}})$ to $\mathcal{T}_{0}(\alpha_{0,H_{n,0}})$ further requires that any directions (e.g., variables or basis functions) in $H_{n,0}\cap H_{n}^{\perp}$ contribute negligibly to the union model representer $\alpha_{0,H_{n,0}}$.
To further clarify these conditions, suppose the working model $H_{n}:=H_{\phi_{n}}$ and the oracle model $H_{0}:=H_{\phi_{0}}$ are induced by feature transformations. For a transformation $\phi:\mathcal{A}\times\mathcal{Z}\times\mathcal{S}\rightarrow\mathbb{R}^{m}$, define $H_{\phi}:=\{f\circ\phi:f:\mathbb{R}^{m}\rightarrow\mathbb{R}\}$. The combined model $H_{n,0}$ is given by $H_{(\phi_{n},\phi_{0})}$, where $(\phi_{n},\phi_{0})$ denotes the feature map formed by stacking $\phi_{n}$ and $\phi_{0}$. Theorem 4 implies that the approximation bias vanishes if the nuisance functions derived from $\phi_{n}$ and $(\phi_{n},\phi_{0})$ converge to those derived from $\phi_{0}$. The special case $\phi_{n}=q_{n}$ and $\phi_{0}=q_{0}$ recovers Theorem 5 as a corollary. In Lemma 14 (Appendix H.1), we show that with features $X$ and outcome $Y$, the $L^{2}(P_{0})$ error of estimating $E_{0}[Y\mid\phi_{0}(X)]$ using either $E_{0}[Y\mid\phi_{n}(X),\mathcal{D}_{n}]$ or $E_{0}[Y\mid\phi_{n}(X),\phi_{0}(X),\mathcal{D}_{n}]$ is bounded by the feature approximation error $\sqrt{\int\|\phi_{n}(x)-\phi_{0}(x)\|_{\mathbb{R}^{m}}^{2}P_{0,X}(dx)}$. A sufficient condition for this bound is that the map $(t_{1},t_{2})\mapsto E_{0}[Y\mid\phi_{n}(X)=t_{1},\phi_{0}(X)=t_{2},\mathcal{D}_{n}]$ is almost surely Lipschitz continuous, generalizing Condition D2.
We now present our main result on the asymptotic linearity and superefficiency of the ADRL estimator $\psi_{n,H_{n}}$ for $\Psi(P_{0})$. In the following conditions, we define for each $P\in\mathcal{P}$ and model $H$:
$\varphi_{P,H}(s,a,y,s^{\prime}):=\mathcal{T}_{P}(\alpha_{P,H})(a,s)\left\{y+\gamma V^{\pi}(q_{P,H}^{\pi})(s^{\prime})-q_{P,H}^{\pi}(a,s)\right\}+m(s,a,q_{P,H})-\Psi_{H}(P),$
which aligns with the influence function in Theorem 1 for $P\in\mathcal{P}_{H}$. Let $\varphi_{n,H_{n}}$ denote the estimator of the influence function $\varphi_{0,H_{n}}$, obtained by plugging in our nuisance estimators.
  1. (C4)
    Consistency: $n^{-\frac{1}{2}}(P_{n}-P_{0})\{\varphi_{n,H_{n}}-\varphi_{0,H_{n}}\}=o_{p}(1)$.
  2. (C5)
    Nuisance estimation rate: $\|\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\|_{P_{0}}\|\mathcal{T}_{0}(\lambda_{n,H_{n}})-\mathcal{T}_{0}(\lambda_{0,H_{n}})\|_{P_{0}}=o_{p}(n^{-1/2})$.
  3. (C6)
    Stabilization of selected model: $n^{-\frac{1}{2}}(P_{n}-P_{0})\{\varphi_{0,H_{n}}-\varphi_{0,H_{0}}\}=o_{p}(1)$.
  4. (C7)
    Model approximation error: $\|\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n,0}})\|_{P_{0}}\|\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\|_{P_{0}}=o_{p}(n^{-\frac{1}{2}}).$
Theorem 10.
Assume C1 holds with $H=H_{n}$ and $H=H_{0}$. Suppose that $H_{n}$ converges to an oracle submodel $H_{0}$ with $q_{0}\in H_{0}$ in the sense that conditions C4-C7 hold. Then, $\psi_{n,H_{n}}-\Psi(P_{0})=(P_{n}-P_{0})\varphi_{0,H_{0}}+o_{p}(n^{-\frac{1}{2}})$. If, in addition, the conditions of Theorem 2 hold with $H:=H_{0}$, then $\psi_{n,H_{n}}$ is a $P_{0}$-regular and efficient estimator for the oracle parameter $\Psi_{H_{0}}$ under the nonparametric statistical model.
Together, C4 and C5 imply that $\psi_{n,H_{n}}-\Psi_{H_{n}}(P_{0})=(P_{n}-P_{0})\varphi_{0,H_{n}}+o_{p}(n^{-\frac{1}{2}})$, so that $\psi_{n,H_{n}}$ is debiased for the working parameter $\Psi_{H_{n}}(P_{0})$. Conditions C6 and C7, which ensures that data-driven model selection preserves the validity of the debiased machine learning estimator, appear in prior works on ADML (van der Laan et al., 2023, 2024d). Condition C6 is an asymptotic stability condition requiring the EIF for the learned model $H_{n}$ to converge to the EIF for the oracle submodel $H_{0}$, which necessitates that $\mathcal{T}_{0}(\alpha_{0,H_{n}})$ and $q_{0,H_{n}}$ are asymptotically consistent with their oracle counterparts. Condition C7 ensures the parameter approximation bias satisfies $\Psi_{n}(P_{0})-\Psi(P_{0})=o_{p}(n^{-\frac{1}{2}})$ in view of Theorem 4.
Appendix C Additional details for Section 5
C.1 Empirical Bellman calibration of Alg. 2
Lemma 11.
Suppose that $f_{n}^{*}$ is the fixed point isotonic regression solution to the calibrated fitted Q-iteration algorithm in Alg. 2, such that:
$f_{n}^{*}=\operatorname*{argmin}_{f\in\mathcal{F}_{iso}}\sum_{i=1}^{n}\{Y_{0,i}+\gamma V_{f_{n}^{*}\circ q_{n}}^{\pi}(S_{1,i}))-f(q_{n}(A_{0,i},S_{0,i}))\}^{2}.$
Then, $q_{n}^{*}:=f_{n}^{*}\circ q_{n}$ satisfies, for each transformation $f:\mathbb{R}\rightarrow\mathbb{R}$, the empirical orthogonality condition:
$\sum_{i=1}^{n}f(q_{n}^{*}(A_{0,i},S_{0,i}))\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}^{*}(A_{0,i},S_{0,i})\}=0.$
Proof.
The proof follows from Lemma 4 in Van Der Laan et al. (2023) with minor notational changes. Recall that $f_{n}^{*}$ is the unique càdlàg piecewise constant solution of the isotonic regression problem with jumps occurring only at observed values of $q_{n}$. For any transformation $f:\mathbb{R}\rightarrow\mathbb{R}$, we claim that $f_{n}^{*}+\varepsilon(h\circ f_{n}^{*})$ is monotone nondecreasing for $\varepsilon$ sufficiently close to zero. To see this, note that $f_{n}^{*}$ is a step function with only finitely many jumps. As a consequence, $h\circ f_{n}^{*}$ is also a step function with the same jump points as $f_{n}^{*}$. By taking $\varepsilon$ close enough to zero, we can guarantee that the maximum jump size of $\varepsilon(h\circ f_{n}^{*})$ is smaller than the minimum jump size of $f_{n}^{*}$. For all $\varepsilon$ sufficiently close to zero, it must then be the case that $f_{n}^{*}+\varepsilon(h\circ f_{n}^{*})$ is also monotone nondecreasing and, thus, an element of $\mathcal{F}_{\text{iso}}$. Since $f_{n}^{*}$ is the empirical risk minimizer over $\mathcal{F}_{\text{iso}}$, we must have that
$\frac{d}{d\varepsilon}\sum_{i=1}^{n}\left\{Y_{0,i}+\gamma V_{f_{n}^{*}\circ q_{n}}^{\pi}(S_{1,i}))-(f_{n}^{*}+\varepsilon(h\circ f_{n}^{*}))(q_{n}(A_{0,i},S_{0,i}))\right\}^{2}\bigg{|}_{\varepsilon=0}=0,$
which implies that
$\sum_{i=1}^{n}f(q_{n}^{*}(A_{0,i},S_{0,i}))\left\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}^{*}(A_{0,i},S_{0,i})\right\}=0.$
Since the transformation $f$ was arbitrary, the result then follows. ∎
C.2 Estimation of Riesz representer for calibrated fitted Q-iteration
An empirical plug-in estimator $d_{n}^{*}$ of $\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})$ is given by $T_{n,q_{n}^{*}}(\alpha_{n,q_{n}^{*}})$, where
$T_{n,q_{n}^{*}}(\alpha)=(a,s)\mapsto\alpha(a,s)-\gamma E_{P_{n}}[\alpha(A,S_{1})\mid q_{n}^{*}(A_{0},S_{0})=q_{n}^{*}(a,s)]$
is the empirical Bellman operator induced by the empirical distribution $P_{n}$ of $\{(S_{0,i},A_{0,i},S_{1,i})\}_{i=1}^{n}$, and $\alpha_{n,q_{n}^{*}}$ is obtained by solving
$\operatorname*{argmin}_{f\circ q_{n}^{*};f:\mathbb{R}\rightarrow\mathbb{R}}\frac{1}{n}\sum_{i=1}^{n}\left[\{T_{n,q_{n}^{*}}(f\circ q_{n}^{*})(A_{0,i},S_{0,i})\}^{2}-2m(S_{0,i},A_{0,i},f\circ q_{n}^{*})\right],$
which is a parametric M-estimation problem that can be computed using numerical solvers.
C.3 Bellman Calibration with Cross-Fitting
Algorithm 3 presents a cross-fitted variant of isotonic Bellman calibration (Algorithm 2). We note that the isotonic calibration step in Algorithm 2 should not itself be cross-fitted. Instead, isotonic regression should be applied using the cross-fitted estimates obtained by pooling the out-of-fold predictions. Importantly, this additional fitting step does not compromise the theoretical guarantees of DML, as $\mathcal{F}_{\text{iso}}$ has controlled complexity, being a Donsker class (van der Laan et al., 2024c).
Algorithm 3 Isotonic Bellman Calibration with Cross-Fitting
1:  Input: Data $\{(S_{0,i},A_{0,i},Y_{0,i},S_{1,i})\}_{i=1}^{n}$
2:        Cross-fitted estimators $\{q_{n}^{(-i)}\}_{i=1}^{n}$, with each $q_{n}^{(-i)}$ independent of $(S_{0,i},A_{0,i},Y_{0,i},S_{1,i})$
3:        Stopping threshold $\varepsilon\approx 0$
4:  Initialize $q_{n}^{*(-i,0)}(a,s):=q_{n}^{(-i)}(a,s)$ for each $i$;
5:  for $k=0,1,2,\ldots$ do
6:     Compute $f_{n}^{(k+1)}$ by solving:
$\operatorname*{argmin}_{f\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\left\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*(-i,k)})(S_{1,i})-f(q_{n}^{(-i)}(A_{0,i},S_{0,i}))\right\}^{2};$
7:     Update $q_{n}^{*(-i,k+1)}(a,s):=f_{n}^{(k+1)}(q_{n}^{(-i)}(a,s))$ for each $i$;
8:     if $\|q_{n}^{*(-i,k+1)}-q_{n}^{*(-i,k)}\|_{P_{n}}<\varepsilon$ then
9:        Set $q_{n}^{*(-i)}:=q_{n}^{*(-i,k+1)}$ for each $i$;
10:        break;
11:     end if
12:  end for
13:  return  $\{q_{n}^{*(-i)}\}_{i=1}^{n}$;
The procedure begins with an initial collection of Q-function estimators $\{q_{n}^{(-i)}\}_{i=1}^{n}$, each trained on a subsample that excludes the $i$th observation to preserve independence. These out-of-fold predictions serve as the input to an iterative calibration procedure that updates each fold-specific Q-function through composition with an isotonic regression fit. At each iteration, a global isotonic calibrator $f_{n}^{(k+1)}\in\mathcal{F}_{\text{iso}}$ is trained to regress the pseudo-outcomes $Y_{0,i}+\gamma V^{\pi}(q_{n}^{*(-i,k)})(S_{1,i})$ on the out-of-fold predictions $q_{n}^{(-i)}(A_{0,i},S_{0,i})$. The updated estimate $q_{n}^{*(-i,k+1)}$ is then defined as the composition $f_{n}^{(k+1)}\circ q_{n}^{(-i)}$. The process continues until convergence in $L_{2}(P_{n})$, at which point the final cross-fitted calibrated estimators $\{q_{n}^{*(-i)}\}_{i=1}^{n}$ are returned. This calibration procedure leverages the entire dataset while preserving fold-level independence in each update, thereby recovering full-sample efficiency without violating the theoretical guarantees of cross-fitted debiased machine learning. We refer the reader to van der Laan et al. (2024a) and van der Laan et al. (2024c) for additional examples of calibration on cross-fitted estimates. Importantly, this additional fitting from calibration does not compromise the theoretical guarantees of DML, as $\mathcal{F}_{\text{iso}}$ has controlled complexity, being a Donsker class (van der Laan et al., 2024c; Rabenseifner et al., 2025).
C.4 Discussions of conditions of Theorem 7
Condition D1 ensures pathwise differentiability of $\Psi_{q_{0}}$, $\Psi_{q_{n}^{*}}$, and $\Psi_{(q_{n}^{*},q_{0})}$, and requires overlap only in the lower-dimensional feature space induced by the $Q$-function. It is therefore significantly weaker than the condition required for pathwise differentiability of $\Psi$ in the full nonparametric model. Condition D3 holds if the calibrated estimator $q_{n}^{*}$ is $o_{p}(n^{-1/4})$-consistent for $q_{0}$ in $L^{2}(P_{0})$, and if the best approximation $q_{0,q_{n}^{*}}^{\pi}$ to $q_{0}$ given $q_{n}^{*}$ satisfies $\|\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}}^{\pi})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0})\|_{P_{0}}=o_{p}(n^{-1/4})$. Together, these imply both $\|q_{n}^{*}-q_{0}\|_{P_{0}}=o_{p}(n^{-1/4})$ and the required approximation rate in the transformed space. General results on isotonic calibration suggest that $q_{n}^{*}$ converges at least as fast—and possibly faster—than the initial estimator $q_{n}$, up to an asymptotically negligible error of order $n^{-1/3}$ (van der Laan and Alaa, 2025). Condition D4 is satisfied if (i) $\|\varphi_{n,q_{n}^{*}}-\varphi_{0,q_{0}}\|_{P_{0}}=o_{p}(1)$, and (ii) the difference lies in a Donsker class or if the initial estimator $q_{n}$ is estimated via sample splitting or cross-fitting (van der Laan et al., 2024a; Rabenseifner et al., 2025). Under boundedness and D2, the first condition holds whenever $\|q_{n}^{*}-q_{0}\|_{P_{0}}=o_{p}(1)$. This empirical process condition holds under mild conditions when Alg. 3 is used.
Appendix D Additional details on experiments
D.1 Simulation design
We generate data from a discrete-state Markov process, where each individual is characterized by a state $S=(\text{engagement},\text{churn risk},\text{tenure},\text{overlap})$, with each variable taking values in $\{0,1,2\}$. The initial state follows $\text{engagement}\sim\text{Multinom}(0.5,0.3,0.2)$, $\text{churn risk}\sim\text{Multinom}(0.25,0.25,0.25)$, $\text{tenure}\sim\text{Multinom}(0.25,0.25,0.25)$, and $\text{overlap}\sim\text{Multinom}(0.7,0.3,0.2)$. Treatment is assigned as $Z\sim\text{Bernoulli}(\pi)$ with $\pi=0.25$. State transitions evolve as follows: tenure increments deterministically as $T_{t+1}=\min(T_{t}+1,2)$. Engagement follows a random walk, where the probability of decrementing is $p_{0}(s_{t})=0.8-C_{t}/5$ for $Z=0$ and $\min(0.1+(0.8-C_{t}/5),1)$ for $Z=1$, with $E_{t+1}=\min(\max(E_{t}+2B_{0}-1,0),2)$, where $B_{0}\sim\text{Bernoulli}(p_{0}(Z))$. Churn risk evolves similarly, with $p_{1}(Z)=0.6$ for $Z=0$ and $0.4$ for $Z=1$, and updates as $C_{t+1}=\min(\max(C_{t}+2B_{1}-1,0),2)$, where $B_{1}\sim\text{Bernoulli}(p_{1}(Z))$. Overlap updates as $O_{t+1}=\min(O_{t}+1,2)$ if $Z=1$ and $B_{2}=1$, otherwise $O_{t+1}=0$, where $B_{2}\sim\text{Bernoulli}(\beta)$. The overlap parameter $\beta$ controls the degree of overlap between states over time, with larger values indicating less overlap. The reward is generated as $Y_{t}\mid(S_{t},Z)\sim\text{Bernoulli}(\sigma(-0.5+1\{O_{t}>0\}+T_{t}/2+0.3Z+1\{E_{t}>0\}/2-C_{t}/2))$, where $\sigma(x)=1/(1+e^{-x})$.
D.2 Additional experimental results
(a) Good overlap ($\beta=0$).
(b) Moderate overlap ($\beta=0.3$).
Figure 3: Bias, standard error (SE), and coverage across discount factors $\gamma$ for various values of $\beta$.
Appendix E Details on data-fusion application
E.1 Background on confounding-robust data-fusion
We consider a data fusion setting in which experimental data ($Z=1$) is augmented with historical control data ($Z=0$), adapting the frameworks of Kallus et al. (2018) and van der Laan et al. (2024d) to Markov decision processes. Define the state as $S_{t}=(Z,\widetilde{S}_{t})$, as in Example 2. Suppose the study indicator $Z$ denotes whether a study unit belongs to an observational study $(Z=0)$, such as historical data, or a randomized experiment $(Z=1)$, such as a randomized control trial. We address the data-fusion problem of augmenting experimental data with potentially biased observational data to increase power by effectively enlarging the sample size. While randomization ensures unbiasedness in the experimental study, incorporating observational data can introduce bias from unmeasured confounding unless strong, untestable assumptions are made. Our goal is to combine these data sources in a confounding-robust manner that retains the unbiasedness of the “gold-standard” experiment while enhancing statistical efficiency.
A flexible approach for integrating randomized and observational data involves generating a biased estimate from the pooled data, learning a bias function between the data sources, and adjusting the biased estimate to obtain an unbiased causal effect estimate (Kallus et al., 2018; Wu and Yang, 2022; Shyr et al., 2023; van der Laan et al., 2024d). However, without model assumptions on the data-generating distribution, nonparametric efficient estimators like those constructed via debiased machine learning asymptotically gain no efficiency from including biased observational data (Dang et al., 2022). The efficiency gain from a larger sample size is completely offset by increased variance from learning the confounding bias function. When restricted to regular estimators, efficiency gains require imposing parametric or semiparametric restrictions on the bias function, which may induce estimation and confounding bias if these assumptions are violated. To address this limitation in cross-sectional studies, van der Laan et al. (2024d) proposed the ADML framework, which uses the highly adaptive lasso to learn model assumptions data-adaptively, yielding nonparametric superefficient estimators.
E.2 ADML Methodology
[…]
As outlined in Section B, an ADML estimator can be constructed by learning a model $H_{n}$ and performing inference for the data-adaptive parameter $\Psi_{H_{n}}$. According to the theory in Section B, the ADML estimator is, under certain conditions, asymptotically linear and efficient for the oracle parameter $\Psi_{H_{0}}$ corresponding to a limiting oracle model $H_{0}$ to which $H_{n}$ converges. For example, $H_{n}$ can be learned using fitted Q-iteration for $q_{0}(a,1,s)$ with the Highly Adaptive Lasso (Benkeser and Van Der Laan, 2016), as in van der Laan et al. (2024d), where a preliminary estimate of the observational $Q$-function $q_{0}(a,0,s)$ is used as an offset, allowing the bias function to be directly modeled. Consequently, the ADML estimator is asymptotically linear and superefficient for $\Psi(P_{0})$ under the nonparametric model, maintaining robustness to model misspecification while gaining efficiency by pooling the two studies when there is learnable structure in the bias function. This approach directly generalizes the approach of van der Laan et al. (2024d) for short-term causal effects in cross-sectional studies, which corresponds to the case where $\gamma=0$.
We now propose a specific ADML estimator that leverages isotonic Bellman calibration, as described in Section 5, to construct superefficient estimators of $\Psi(P_{0})$ while avoiding the computational challenges and instability of estimating the Riesz representer via min-max optimization. Suppose the observational data is far larger than the experimental study, as often occurs in industrial applications where historical data vastly outnumbers randomized data from A/B tests. In this setting, the observational $Q$-function $q_{0}(a,0,s)$ can be estimated very accurately and is effectively known. Assuming $q_{0}(a,0,s)$ is known, we define the oracle model:
[…]
To construct an ADML estimator, we use a modified version of calibrated FQI that incorporates $q_{0}$ as an offset. Define the modified outcome $\widetilde{Y}_{0}=Y_{0}+\gamma V_{q_{0}}^{\pi}(A_{1},0,S_{1})-q_{0}(A_{0},0,S_{0})$. By the Bellman equation for $q_{0}$, we have $E_{0}[\widetilde{Y}_{0}\mid A_{0},Z,\widetilde{S}_{0},Z=0]=0$ and $E_{0}[\widetilde{Y}_{0}\mid A_{0},S_{0},Z=1]=E_{0}[b^{\pi}_{0}(A_{0},Z,\widetilde{S}_{0})-\gamma V_{b^{\pi}_{0}}^{\pi}(1,S_{1})\mid A_{0},S_{0},Z=1]$, so $b^{\pi}_{0}$ satisfies the Bellman equation for $\widetilde{Y}_{0}$ given $Z=1$. In our modification of Algorithm 2, $q_{n}$ is replaced by $b_{n}^{\pi}$, each $Y_{0,i}$ is replaced by $\widetilde{Y}_{0,i}:=Y_{0,i}+\gamma V_{q_{0}}^{\pi}(0,S_{1,i})-q_{0}(A_{0,i},0,S_{0,i})$, and calibration is applied only using observations with $Z_{i}=1$. This corresponds to calibrated fitted Q-iteration where calibration uses the class $\{(a,s)\mapsto q_{0}(a,0,s)+zf(b_{n}(a,s)):f\text{ is an isotonic function}\}$. Applying this procedure, we obtain an isotonic-calibrated bias function $b_{n}^{*}$ that satisfies the empirical orthogonality condition for each transformation $f:\mathbb{R}\rightarrow\mathbb{R}$:
$\sum_{i=1}^{n}Z_{i}f(b_{n}^{*}(A_{0,i},S_{0,i}))\{\widetilde{Y}_{0,i}+\gamma V_{b^{*}_{n}}^{\pi}(1,S_{1,i})-b_{n}^{*}(A_{0,i},S_{0,i})\}=0.$
A debiased plug-in estimator is then given by $\frac{1}{\sum_{i=1}^{n}Z_{0,i}}\sum_{i=1}^{n}Z_{0,i}m(A_{0,i},S_{0,i},q^{*}_{n,0})$, where $q^{*}_{n,0}(a,s)=q_{0}(a,0,s)+zb_{n}^{*}(a,s)$ is the calibrated estimator of $q_{0}^{(\pi)}$. An application of the results from Section 5 with the offset outcome $\widetilde{Y}_{0}:=Y_{0}+\gamma V_{q_{0}}^{\pi}(0,S_{1})-q_{0}(A_{0},0,S_{1,i})$ conditional on $Z_{0}=1$ establishes the asymptotic linearity and superefficiency of this estimator under the stated conditions.
Appendix F Proofs for Section 3
F.1 Proof of Theorem 8 on uniqueness of Bellman projection
Proof of Theorem 8.
(Existence) By Condition A1, the operator $\mathcal{K}_{P}$ is compact on $\overline{H}_{P}$, so $\mathcal{T}_{P}=I_{P}-\gamma\mathcal{K}_{P}$ is a Fredholm operator of index zero on the Banach space $(\overline{H}_{P},\|\cdot\|_{P})$ (Riesz–Schauder theory, Conway (1994)). In particular, its range $\mathcal{R}:=\mathcal{T}_{P}(\overline{H}_{P})$ is a closed subspace of $L^{2}(P_{A_{0},S_{0}})$.
Now write $Y_{0}=\mu_{P}(S_{0},A_{0})+\epsilon$, where $E_{P}[\epsilon\mid A_{0},S_{0}]=0$. Then for any $q\in\overline{H}_{P}$,
$E_{P}\left[(Y_{0}-\mathcal{T}_{P}(q))^{2}\right]=E_{P}[\epsilon^{2}]+\|\mu_{P}-\mathcal{T}_{P}(q)\|_{P}^{2}.$
Hence the minimizer in (8) corresponds to the projection $\mu_{P,H}:=\operatorname*{argmin}_{\mu\in\mathcal{R}}\|\mu_{P}-\mu\|_{P}^{2}$, which satisfies $\mu_{P,H}\in\mathcal{R}$. Therefore, there exists $q_{P,H}\in\overline{H}_{P}$ such that $\mathcal{T}_{P}(q_{P,H})=\mu_{P,H}$, i.e., a solution to (8) exists. Here, we used the standard fact that the projection onto a closed subspace of a Hilbert space always exists and is unique.
(Uniqueness) Under Condition A2, $\gamma^{-1}$ is not an eigenvalue of $\mathcal{K}_{P}$, so $\mathcal{T}_{P}=I_{P}-\gamma\mathcal{K}_{P}$ has a trivial null space on $\overline{H}_{P}$. Since $\mathcal{T}_{P}$ is Fredholm of index zero (Conway, 1994), injectivity implies surjectivity. Thus, $\mathcal{T}_{P}$ is bijective with a bounded inverse $\mathcal{T}_{P}^{-1}:\mathcal{R}\to\overline{H}_{P}$, and $q_{P,H}=\mathcal{T}_{P}^{-1}(\mu_{P,H})$ is the unique solution to (8). ∎
F.2 Derivation of EIF in Theorem 1 and Theorem 2
Proof of EIF in Theorem 2.
By Theorem 8, we have $q_{P,H}=\mathcal{T}_{P}^{-1}(\mu_{P,H})$, where $\mathcal{T}_{P}^{-1}$ is a bounded linear operator and $\mu_{P,H}:=\operatorname*{argmin}_{\mu\in\mathcal{T}_{P}(\overline{H}_{P})}\|\mu_{P}-\mu\|_{P}$ is the projection of $\mu_{P}$ onto the range of $\mathcal{T}_{P}$. Since $\mathcal{T}_{P}$ is invertible on $L^{2}(P)$, its range is closed, and $\overline{H}_{P}$ denotes the closure of $H$ in $L^{2}(P)$. Consequently, $q_{P,H}$ is uniquely identified as an element of the $L^{2}(P)$ closure of $H$.
Let $P\in\mathcal{P}$ be arbitrary, and let $(P_{\varepsilon,\phi}:\varepsilon\in\mathbb{R})$ denote a regular submodel satisfying: (i) $\frac{dP_{\varepsilon,\phi}}{dP}$ exists; (ii) $P_{\varepsilon,\phi}=P$ at $\varepsilon=0$; and (iii) the score at $\varepsilon=0$ is $\phi\in T_{\mathcal{P}}(P)$. We now show that the parameter $\Psi_{H}$ is pathwise differentiable along any such path and satisfies the inner product representation:
$\frac{d}{d\varepsilon}\Psi_{H}(P_{\varepsilon,\phi})\big{|}_{\varepsilon=0}=\langle\varphi_{P},\phi\rangle_{P},$
where $\varphi_{P}$ denotes the efficient influence function (EIF) of $\Psi_{H}$. To compute the pathwise derivative of $\Psi_{H}$, we will use the representation $\Psi_{H}(P)=\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P}(q_{P,H})\rangle_{P}=\langle\mathcal{T}_{P}(\alpha_{P,H}),\Pi_{P}(\mu_{P})\rangle_{P}$, which is guaranteed by C1.
We adopt the following notation. Let $W=(S_{0},A_{0},Y_{0},S_{1})$, and let $w=(s,a,y,s^{\prime})$ denote a generic realization of $W$. For each function $h$, define the next-state value function $\overline{V}^{\pi}(h)$ as $w\mapsto V^{\pi}(h)(s^{\prime})$. Throughout, we will view $\overline{V}^{\pi}(h)$ as an element of $L^{2}(P_{S_{1}})$ and $V^{\pi}(h)$ as an element of $L^{2}(P_{S_{0}})$. Define $\Pi_{P}:L^{2}(P)\rightarrow H$ as the $L^{2}(P_{A_{0},S_{0}})$ projection operator onto $H$, given pointwise by $\Pi_{P}f:=\operatorname*{argmin}_{h\in H}\|f-h\|_{P}$.
For each $h\in\mathcal{H}$, we denote the pathwise derivative $d\mathcal{T}_{P}(h):T_{\mathcal{P}}(P)\rightarrow L^{2}(P)$ of $\mathcal{T}_{P}(h)$ by the map $\phi\mapsto\frac{d}{d\varepsilon}\mathcal{T}_{P_{\varepsilon,\phi}}(h)\big{|}_{\varepsilon=0}$. We can compute this pathwise derivative as follows:
$\frac{d}{d\varepsilon}\mathcal{T}_{P_{\varepsilon,\phi}}(h)\big{|}_{\varepsilon=0}$ $=-\gamma\frac{d}{d\varepsilon}\int V^{\pi}(h)(s^{\prime},Z)P_{\varepsilon,\phi}(S_{1}=ds^{\prime}\mid A_{0},S_{0})\big{|}_{\varepsilon=0}$
$=-\gamma E_{P}[V^{\pi}(h)(S_{1})\left\{\phi(W)-E_{P}[\phi(W)\mid A_{0},S_{0}]\right\}\mid A_{0},S_{0}]$
$=-\gamma E_{P}[\left\{V^{\pi}(h)(S_{1})-E_{P}[V^{\pi}(h)(S_{1})\mid A_{0},S_{0}]\right\}\phi(W)\mid A_{0},S_{0}]$
$=-E_{P}[\left\{\gamma V^{\pi}(h)(S_{1})+\mathcal{T}_{P}(h)(A_{0},S_{0})-h(A_{0},S_{0})\right\}\phi(W)\mid A_{0},S_{0}]$
$=E_{P}[\left\{h(A_{0},S_{0})-\gamma\overline{V}^{\pi}(h)(S_{1})-\mathcal{T}_{P}(h)(A_{0},S_{0})\right\}\phi(W)\mid A_{0},S_{0}].$
In the final equality, we used the fact that $V^{\pi}(h)(S_{1})=\overline{V}^{\pi}(h)(S_{1})$ by definition. We will make use of the following expression:
$\langle f,d\mathcal{T}_{P}(h)(\phi)\rangle_{P}=\langle f,\phi\left\{h-\gamma\overline{V}^{\pi}(h)-\mathcal{T}_{P}(h)\right\}\rangle_{P}\quad\text{for all }f\in L^{2}(P),$
where $\overline{V}^{\pi}(h)$ is viewed as a function of $S_{1}$.
The first-order conditions of the optimization problem defining $q_{P,H}$ imply that $q_{P,H}$ satisfies the restricted moment equation:
$\langle\mathcal{T}_{P}(h),\mu_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}=0\text{ for all }h\in\mathcal{H}.$
By the product rule of differentiation, we have
$\frac{d}{d\varepsilon}\Psi_{H}(P_{\varepsilon,\phi})\Big{|}_{\varepsilon=0}$ $=\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P_{\varepsilon,\phi}}(q_{P_{\varepsilon,\phi}})\rangle_{P}\Big{|}_{\varepsilon=0}+\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P}(q_{P,H})\rangle_{P_{\varepsilon,\phi}}\Big{|}_{\varepsilon=0}.$
First Term.
We know $\mathcal{T}_{P}(q_{P,H})$ is determined by:
$\langle\mathcal{T}_{P}(h),\mathcal{T}_{P}(q_{P,H})\rangle_{P}=\langle\mathcal{T}_{P}(h),Y_{0}\rangle_{P}\text{ for all }h\in\mathcal{H}.$
Hence, taking the pathwise derivative of both sides, we find, for all $h\in\mathcal{H}$, that
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(h),\mathcal{T}_{P_{\varepsilon,\phi}}(q_{P_{\varepsilon,\phi}})\rangle_{P_{\varepsilon,\phi}}\Big{|}_{\varepsilon=0}=\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(h),Y_{0}\rangle_{P_{\varepsilon,\phi}}\Big{|}_{\varepsilon=0}.$
Thus, by the chain rule, we have
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(h),\mathcal{T}_{P_{\varepsilon,\phi}}(q_{P_{\varepsilon,\phi}})\rangle_{P}\Big{|}_{\varepsilon=0}+\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(h),\mathcal{T}_{P}(q_{P,H})\rangle_{P_{\varepsilon,\phi}}\Big{|}_{\varepsilon=0}=\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(h),Y_{0}\rangle_{P_{\varepsilon,\phi}}\Big{|}_{\varepsilon=0}.$
Therefore,
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(h),\mathcal{T}_{P_{\varepsilon,\phi}}(q_{P_{\varepsilon,\phi}})\rangle_{P}\Big{|}_{\varepsilon=0}$ $=\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(h),Y_{0}\rangle_{P}\Big{|}_{\varepsilon=0}+\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(h),Y_{0}\rangle_{P}\Big{|}_{\varepsilon=0}$
$\quad-\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(h),\mathcal{T}_{P}(q_{P,H})\rangle_{P}\Big{|}_{\varepsilon=0}-\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(h),\mathcal{T}_{P}(q_{P,H})\rangle_{P}\Big{|}_{\varepsilon=0}$
$=\langle d\mathcal{T}_{P}(h)(\phi),Y_{0}\rangle_{P}+\langle\phi,\mathcal{T}_{P}(h)Y_{0}-E_{P}[\mathcal{T}_{P}(h)Y_{0}]\rangle_{P}$
$\quad-\langle\phi,\mathcal{T}_{P}(h)\mathcal{T}_{P}(q_{P,H})-E_{P}[\mathcal{T}_{P}(h)\mathcal{T}_{P}(q_{P,H})]\rangle_{P}-\langle d\mathcal{T}_{P}(h)(\phi),\mathcal{T}_{P}(q_{P,H})\rangle_{P}$
$=\langle d\mathcal{T}_{P}(h)(\phi),\mu_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}+\langle\phi,\mathcal{T}_{P}(h)\{Y_{0}-\mathcal{T}_{P}(q_{P,H})\}\rangle_{P}$
$\quad+\langle\phi,E_{P}[\mathcal{T}_{P}(h)\mathcal{T}_{P}(q_{P,H})]-E_{P}[\mathcal{T}_{P}(h)Y_{0}]\rangle_{P}.$
The above holds for all $h\in H$, and therefore also for $h=\alpha_{P,H}$ by continuity of the inner product and of $\mathcal{T}_{P}$, since $\alpha_{P,H}$ lies in the $L^{2}(P)$-closure of $H$. Hence, taking $h=\alpha_{P,H}$ and using that $E_{P}[\mathcal{T}_{P}(\alpha_{P,H})\mathcal{T}_{P}(q_{P,H})]=E_{P}[\mathcal{T}_{P}(\alpha_{P,H})Y_{0}]$, it follows that
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P_{\varepsilon,\phi}}(q_{P_{\varepsilon,\phi}})\rangle_{P}\Big{|}_{\varepsilon=0}$ $=\langle d\mathcal{T}_{P}(\alpha_{P,H})(\phi),\mu_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}+\langle\phi,\mathcal{T}_{P}(\alpha_{P,H})\{Y_{0}-\mathcal{T}_{P}(q_{P,H})\}\rangle_{P}$
$\quad+\langle\phi,E_{P}[\mathcal{T}_{P}(\alpha_{P,H})\mathcal{T}_{P}(q_{P,H})]-E_{P}[\mathcal{T}_{P}(\alpha_{P,H})Y_{0}]\rangle_{P}$
$=\langle d\mathcal{T}_{P}(\alpha_{P,H})(\phi),\mu_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}+\langle\phi,\mathcal{T}_{P}(\alpha_{P,H})\{Y_{0}-\mathcal{T}_{P}(q_{P,H})\}\rangle_{P}$
$=\langle\phi\{\alpha_{P,H}-\gamma\overline{V}_{\alpha_{P,H}}^{\pi}-\mathcal{T}_{P}(\alpha_{P,H})\},\mu_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}$
$\quad+\langle\phi,\mathcal{T}_{P}(\alpha_{P,H})\{Y_{0}-\mu_{P}\}\rangle_{P}.$
Consequently, this derivative component can be expressed as the inner product $\langle\varphi_{P,1},\phi\rangle$ for the gradient component:
$\varphi_{1,P}:w\mapsto\{\alpha_{P,H}(a,s)-\gamma V_{\alpha_{P,H}}^{\pi}(s^{\prime})\}\{\mu_{P}(a,s)-\mathcal{T}_{P}(q_{P,H})(a,s)\}+\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y-\mu_{P}(a,s)\}.$
Second Term.
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P}(q_{P,H})\rangle_{P_{\varepsilon,\phi}}\Big{|}_{\varepsilon=0}$ $=\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P}(q_{P,H})\rangle_{P}\Big{|}_{\varepsilon=0}$
$\quad+\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P}(q_{P,H})\rangle_{P_{\varepsilon,\phi}}\Big{|}_{\varepsilon=0}.$
To compute this term, we use the Riesz representation property of $\alpha_{P,H}$, which implies:
$\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P}(h)\rangle_{P}=E_{P}[m(S_{0},A_{0},h)]\text{ for all }h\in\mathcal{H}.$
Taking the pathwise derivative of both sides and applying the chain rule, we find:
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P_{\varepsilon,\phi}}(h)\rangle_{P_{\varepsilon,\phi}}\big{|}_{\varepsilon=0}$ $=\frac{d}{d\varepsilon}E_{P_{\varepsilon,\phi}}[m(S_{0},A_{0},h)]\big{|}_{\varepsilon=0};$
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P}(h)\rangle_{P_{\varepsilon,\phi}}\big{|}_{\varepsilon=0}+\langle\mathcal{T}_{P}(\alpha_{P,H}),d\mathcal{T}_{P}(h)(\phi)\rangle_{P}\big{|}_{\varepsilon=0}$ $=\frac{d}{d\varepsilon}E_{P_{\varepsilon,\phi}}[m(S_{0},A_{0},h)]\big{|}_{\varepsilon=0},$
and, hence,
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P}(h)\rangle_{P_{\varepsilon,\phi}}\big{|}_{\varepsilon=0}=\frac{d}{d\varepsilon}E_{P_{\varepsilon,\phi}}[m(S_{0},A_{0},h)]\big{|}_{\varepsilon=0}-\langle\mathcal{T}_{P}(\alpha_{P,H}),d\mathcal{T}_{P}(h)(\phi)\rangle_{P}\big{|}_{\varepsilon=0},$
where we compute
$\frac{d}{d\varepsilon}E_{P_{\varepsilon,\phi}}[m(S_{0},A_{0},h)]\big{|}_{\varepsilon=0}=E_{P}[\phi_{S_{0}}(S_{0})m(S_{0},A_{0},h)]=\langle\phi,m(S_{0},A_{0},h)-E_{P}[m(S_{0},A_{0},h)]\rangle_{P}.$
The above holds for all $h\in H$ and, therefore, also for $q_{P,H}$ by continuity of the inner product and of $\mathcal{T}_{P}$, since $q_{P,H}$ lies in the $L^{2}(P)$-closure of $H$ by Theorem 8. Thus, setting $h=q_{P,H}$, we find that:
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P}(q_{P,H})\rangle_{P_{\varepsilon,\phi}}\big{|}_{\varepsilon=0}=\langle\phi,m(S_{0},A_{0},q_{P,H})-E_{P}[m(S_{0},A_{0},q_{P,H})]\rangle_{P}-\langle\mathcal{T}_{P}(\alpha_{P,H}),d\mathcal{T}_{P}(q_{P,H})(\phi)\rangle_{P}.$
By the definition of $d\mathcal{T}_{P}(q_{P,H})(\phi)$, we have:
$\langle\mathcal{T}_{P}(\alpha_{P,H}),d\mathcal{T}_{P}(q_{P,H})(\phi)\rangle_{P}$ $=-E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\gamma V_{q_{P,H}}^{\pi}(S_{1})+\mathcal{T}_{P}(q_{P,H})-q_{P,H}(A_{0},S_{0})\right\}\phi(Z)\right].$
Thus, $\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P}(q_{P,H})\rangle_{P_{\varepsilon,\phi}}\big{|}_{\varepsilon=0}:=\langle\varphi_{P,2},\phi\rangle_{P}$ for the gradient component:
$\varphi_{P,2}:w\mapsto\mathcal{T}_{P}(\alpha_{P,H})(a,s)\left\{\gamma V_{q_{P,H}}(a,s^{\prime})+\mathcal{T}_{P}(q_{P,H})(a,s)-q_{P,H}(a,s)\right\}\,+\,m(s,a,q_{P,H})-\Psi_{H}(P).$
EIF.
Putting it all together, the EIF $\varphi_{P}:=\varphi_{P,1}+\varphi_{P,2}$ is:
$w\mapsto$ $\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y-\mu_{P}(a,s)\}+\{\alpha_{P,H}(a,s)-\gamma V_{\alpha_{P,H}}^{\pi}(a,s^{\prime})\}\{\mu_{P}(a,s)-\mathcal{T}_{P}(q_{P,H})(a,s)\}$
$\quad+\mathcal{T}_{P}(\alpha_{P,H})(a,s)\left\{\gamma V_{q_{P,H}}(a,s^{\prime})+\mathcal{T}_{P}(q_{P,H})(a,s)-q_{P,H}(a,s)\right\}$
$\quad+m(s,a,q_{P,H})-\Psi_{H}(P).$
Assuming a correct model ($\mathcal{T}_{P}(q_{P,H})=\mu_{P}$), it simplifies to:
$\varphi_{P}:w\mapsto$ $\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y+\gamma V_{q_{P,H}}^{\pi}(a,s^{\prime})-q_{P,H}(a,s)\}$
$\quad+m(s,a,q_{P,H})-\Psi_{H}(P).$
F.3 Derivation of von Mises expansion in Theorem 1 and Theorem 2
We establish the following generalization of the von Mises expansion in Theorem 2, which does not require the nuisance components in the influence function to be compatible with any single distribution $P\in\mathcal{P}$. Let $\widehat{q}_{H},\widehat{\alpha}_{H}\in H$, $\widehat{\mu}_{P}\in L^{\infty}(\lambda)$, and let $\widehat{\mathcal{T}}:L^{\infty}(\lambda)\rightarrow L^{\infty}(\lambda)$ be an arbitrary map. Let $\widehat{P}\in\mathcal{P}$ be a distribution such that $q_{\widehat{P}}^{\pi}=\widehat{q}_{H}$, so that $\Psi(\widehat{P}):=E_{\widehat{P}}[m(S_{0},A_{0},\widehat{q}_{H})]$. Define $\widehat{\varphi}^{*}_{H}$ as the function
$(s,a,y,s^{\prime})\mapsto\;$ $\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(a,s)\left\{y+\gamma V^{\pi}(\widehat{q}_{H})(s^{\prime})-\widehat{q}_{H}(a,s)\right\}$
+$+\left\{\widehat{\alpha}_{H}(a,s)-\gamma V^{\pi}(\widehat{\alpha}_{H})(s^{\prime})-\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(a,s)\right\}\left\{\mu_{0}(a,s)-\widehat{\mathcal{T}}(\widehat{q}_{H})(a,s)\right\}$
$+m(s,a,\widehat{q}_{H})-\Psi(\widehat{P}).$
Theorem 12 (Functional von Mises expansion).
Assume that C1 holds at $P_{0}$. Then, the parameter expansion satisfies: $\Psi_{H}(\widehat{P})-\Psi_{H}(P_{0})=-P_{0}\widehat{\varphi}^{*}_{H}+\widehat{R}_{H}^{*}(P_{0})$, where:
$\widehat{R}_{H}^{*}(P_{0})$ $:=P_{0}\left[\{\mathcal{T}_{0}(\widehat{\alpha}_{H})-\mathcal{T}_{0}(\alpha_{0,H})\}(\mathcal{T}_{0}(q_{0,H})-\mathcal{T}_{0}(\widehat{q}_{H}))\right]$
$\quad+P_{0}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})-\mathcal{T}_{0}(\widehat{\alpha}_{H})\right\}(\mu_{0}-\widehat{\mu})\right]$
$\quad+P_{0}\left[\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})-\mathcal{T}_{0}(\widehat{\alpha}_{H})\}(\widehat{\mathcal{T}}(\widehat{q}_{H})-\mathcal{T}_{0}(\widehat{q}_{H}))\right].$
Proof of Theorem 12.
Let $\widehat{P}$ be a distribution compatible with $\widehat{q}_{H}$ and the marginal distribution of $(S_{0},A_{0})$ used to compute the term $\Psi_{H}(\widehat{P})$ in $\widehat{\varphi}^{*}$. By the law of iterated expectations, it holds that
$\widehat{R}_{H}^{*}(P)$ $:=\Psi_{H}(\widehat{P})-\Psi_{H}(P)+P\widehat{\phi}^{*}_{H}$
$=E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\{\mu_{P}(A_{0},S_{0})-\widehat{\mu}(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\{\widehat{\alpha}_{H}(A_{0},S_{0})-\gamma V^{\pi}(\widehat{\alpha}_{H})(S_{1})\}\{\widehat{\mu}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\left\{\gamma V^{\pi}(\widehat{q}_{H})(S_{1})+\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\widehat{q}_{H}(A_{0},S_{0})\right\}\right]$
$\quad+E_{P}\left[m(S_{0},A_{0},\widehat{q}_{H})-m(S_{0},A_{0},q_{P,H})\right].$
By Riesz representation theorem, it holds that
$E_{P}\left[m(S_{0},A_{0},\widehat{q}_{H})-m(S_{0},A_{0},q_{P,H})\right]$ $=E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(q_{P,H})(A_{0},S_{0})\right\}\right]$
$=E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})-\mu_{P}(A_{0},S_{0})\right\}\right],$
where we used that $\mathcal{T}_{P}(q_{P,H})$ is the $L^{2}(P)$ projection of $\mu_{P}$ onto $\mathcal{T}_{P}(\overline{H}_{P})$. In addition, applying the law of iterated expectations applied to the second and third terms, we find:
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\{\mu_{P}(A_{0},S_{0})-\widehat{\mu}(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\{\widehat{\mu}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right]$
$\quad+E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})-\mu_{P}(A_{0},S_{0})\right\}\right].$
Next, adding and subtracting, the first and third term can be rewritten as
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{\mu_{P}(A_{0},S_{0})-\widehat{\mu}(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\{\mu_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right]$
$\quad+E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})-\mu_{P}(A_{0},S_{0})\right\}\right].$
Adding and subtracting again, the third and fourth terms can be rewritten as
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{\mu_{P}(A_{0},S_{0})-\widehat{\mu}(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\{\mu_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right]$
$\quad+E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mu_{P}(A_{0},S_{0})\right\}\right].$
Combining the second and fourth term, we find
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{\mu_{P}(A_{0},S_{0})-\widehat{\mu}(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\left\{\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\{\mu_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right].$
Using that $\mathcal{T}_{P}(q_{P,H})$ is the $L^{2}(P)$ projection of $\mu_{P}$ onto $\mathcal{T}_{P}(\overline{H}_{P})$, we can show that
$E_{P}\left[\left\{\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\{\mu_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\hskip 28.45274pt=E_{P}\left[\left\{\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\{\mathcal{T}_{P}(q_{P})(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right].$
Substituting this expression, we conclude that
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{\mu_{P}(A_{0},S_{0})-\widehat{\mu}(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\left\{\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\{\mathcal{T}_{P}(q_{P})(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\quad+E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right].$
The first result then follows. In the case where $P,\widehat{P}\in\mathcal{P}_{H}$, we have that $\mu_{P}=\mathcal{T}_{P}(q_{P,H})$ and $\widehat{\mu}=\widehat{\mathcal{T}}(\widehat{q}_{H})$. In this case, the expression simplifies to:
$\widehat{R}_{H}^{*}(P)=E_{P}\left[\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})-\mathcal{T}_{P}(\alpha_{P,H})\}(\mathcal{T}_{P}(q_{P,H})-\mathcal{T}_{P}(\widehat{q}_{H}))]\right].$
Appendix G Proofs for Section 3 on semiparametric DRL
G.1 Asymptotic linearity of DRL estimator under correct specification
In the following conditions and theorem, let $\varphi_{n,H}$ denote the estimator of the IF $\varphi_{0}$ from Theorem 1, obtained by plugging in our nuisance estimators.
  1. (C8)
    Consistency: $n^{-\frac{1}{2}}(P_{n}-P_{0})\{\varphi_{n,H}-\varphi_{0,H}\}=o_{p}(1)$.
  2. (C9)
    Nuisance estimation rate: $\left\|\mathcal{T}_{0}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{0,H})\right\|_{P_{0}}\cdot\|\mathcal{T}_{0}(q_{n,H})-\mathcal{T}_{0}(q_{0,H})\|_{P_{0}}=o_{p}(n^{-\frac{1}{2}})$
Theorem 13 (Asymptotic linearity under correct specification).
Suppose that $P_{0}\in\mathcal{P}_{H}$, meaning $q_{0,H}\in\overline{H}_{P}$. Assume C1 holds, as well as C8 and C9. Then, $\psi_{n,H}-\Psi_{H}(P_{0})=(P_{n}-P_{0})\varphi_{0,H}+o_{p}(n^{-\frac{1}{2}})$. Moreover, $\psi_{n,H}$ is locally robust to misspecification as it is a $P_{0}$-regular and efficient estimator for the working parameter $\Psi_{H}$ under the nonparametric model .
Proof of Theorem 13 .
By C1, we can apply the von Mises expansion in Theorem 1 to conclude that
$\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H}^{\pi})-\Psi(P_{0})=-P_{0}\varphi_{n,H}+R_{n,H}(P_{0}),$
where:
$R_{n,H}(P_{0})$ $:=P_{0}\left[\left\{\widehat{\mathcal{T}}_{n}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{n,H})\right\}\mathcal{T}_{0}(q_{n,H}^{\pi}-q_{0})\right]$
By C9 and the Cauchy-Schwarz inequality, we have that $R_{n,H}(P_{0})=o_{p}(n^{-1/2})$. Thus,
$\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H}^{\pi})-\Psi(P_{0})=-P_{0}\varphi_{n,H}+o_{p}(n^{-1/2}).$
Using the definition of the one-step estimator $\psi_{n,H}=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H}^{\pi})+P_{n}\varphi_{n,H}$, we have
$\psi_{n,H}-\Psi(P_{0})=(P_{n}-P_{0})\varphi_{n,H}+o_{p}(n^{-1/2}).$
By C8, it follows that
$\psi_{n,H}-\Psi(P_{0})$ $=(P_{n}-P_{0})\varphi_{0,H}+(P_{n}-P_{0})\{\varphi_{n,H}-\varphi_{0,H}\}+o_{p}(n^{-1/2})$
$=(P_{n}-P_{0})\varphi_{0,H}+o_{p}(n^{-1/2}),$
as desired. Under A2 and correct specification of $H$, we have that the influence function $\varphi_{0,H}=\varphi_{0,H}^{*}$ is the $P_{0}$-EIF of $\Psi_{H}$ by Theorem 2. Thus, $\psi_{n,H_{n}}$ is a $P_{0}$-regular and efficient estimator for the working parameter $\Psi_{H}$ under the nonparametric model. ∎
G.2 Asymptotic linearity and efficiency of model-robust DRL estimator
Proof of Theorem 3.
By C1, we can apply the von Mises expansion in Theorem 2 to conclude that
$\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H}^{\pi})-\Psi_{H}(P_{0})=-P_{0}\varphi_{n,H}^{*}+R_{n,H}^{*}(P_{0}),$
where:
$R_{n,H}^{*}(P_{0})$ $:=P_{0}\left[\left\{\widehat{\mathcal{T}}_{n}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{n,H})\right\}(\mu_{0}-\mu_{n})\right]$
$\quad+P_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{0,H})\}(\mathcal{T}_{0}(q_{0,H})-\mathcal{T}_{0}(q_{n,H}^{\pi}))\right]$
$\quad+P_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H})\}(\mathcal{T}_{0}(q_{n,H}^{\pi})-\widehat{\mathcal{T}}_{n}(q_{n,H}^{\pi}))\right].$
By C3 and the Cauchy-Schwarz inequality, we have that $R_{n,H}^{*}(P_{0})=o_{p}(n^{-1/2})$. Thus,
$\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H}^{\pi})-\Psi_{H}(P_{0})=-P_{0}\varphi_{n,H}^{*}+o_{p}(n^{-1/2}).$
Using the definition of the one-step estimator $\psi_{n,H}^{*}=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H}^{\pi})+P_{n}\varphi_{n,H}^{*}$, we have
$\psi_{n,H}^{*}-\Psi_{H}(P_{0})=(P_{n}-P_{0})\varphi_{n,H}^{*}+o_{p}(n^{-1/2}).$
By C2, it follows that
$\psi_{n,H}^{*}-\Psi_{H}(P_{0})$ $=(P_{n}-P_{0})\varphi_{0,H}^{*}+(P_{n}-P_{0})\{\varphi_{n,H}^{*}-\varphi_{0,H}^{*}\}+o_{p}(n^{-1/2})$
$=(P_{n}-P_{0})\varphi_{0,H}^{*}+o_{p}(n^{-1/2}),$
as desired. We have that the influence function $\varphi_{0,H}^{*}$ is the EIF of $\Psi_{H}$ by Theorem 2. Thus, $\psi_{n,H_{n}}^{*}$ is a $P_{0}$-regular and efficient estimator for the working parameter $\Psi_{H}$ under the nonparametric model.
Appendix H Proofs for Section 5 on calibrated FQI
H.1 Lemma bounding approximation error of estimated features
In the following lemma, let $X$ be a covariate and $Y\in\mathbb{R}$ be an outcome. For feature transformations $\varphi_{n},\varphi_{0}$, denote $f_{(\varphi_{n},\varphi_{0})}:x\mapsto E_{0}[Y_{0}\mid\varphi_{n}(X)=\varphi_{n}(x),\varphi_{0}(X)=\varphi_{0}(x)]$, $f_{\varphi_{n}}:x\mapsto E_{0}[Y_{0}\mid\varphi_{n}(X)=\varphi_{n}(x)]$, and $f_{\varphi_{0}}:x\mapsto E_{0}[Y_{0}\mid\varphi_{0}(X)=\varphi_{0}(x)]$.
Lemma 14.
Suppose that $(t_{1},t_{2})\mapsto E_{0}[f_{(\varphi_{n},\varphi_{0})}(X)\mid\varphi_{n}(X)=t_{1},\varphi_{0}(X)=t_{2},\mathcal{D}_{n}]$ is almost surely $L$-Lipschitz continuous. Then,
$\|f_{\varphi_{n}}-f_{(\varphi_{n},\varphi_{0})}\|_{P_{0}}\lesssim\|\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}\|_{P_{0}}\text{ and }\|f_{\varphi_{n}}-f_{\varphi_{0}}\|_{P_{0}}\lesssim\|\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}\|_{P_{0}}.$
Proof.
For any real-valued function $f:\mathcal{X}\rightarrow\mathbb{R}$ and a vector-valued function $v:\mathcal{X}\rightarrow\mathbb{R}^{k}$ with $k\in\mathbb{N}$, we define the conditional expectation projection operator $\Pi_{v}:\mathcal{H}\rightarrow\mathcal{H}$ pointwise as $\Pi_{v}f:=\operatorname*{argmin}_{\theta\in\Theta}\|f-\theta\circ v\|$, where $\Theta$ consists of all functions from $\mathbb{R}^{k}\rightarrow\mathbb{R}$. Whenever $v$ and $f$ are nonrandom functions, we have that $\Pi_{v}f:(a,w)\mapsto E_{0}[f(A,W)\,|\,v(A,W)=v(a,w)]$.
Let $g:\mathbb{R}^{k}\times\mathbb{R}^{k}\rightarrow\mathbb{R}$ be a Lipschitz continuous function with constant $L>0$. By Lipschitz continuity, we have that
$\left|g(\varphi_{n}(x),\varphi_{0}(x))-E[g(\varphi_{n}(X),\varphi_{0}(X))|\varphi_{n}(X)=\varphi_{n}(x)]\right|$ $=\left|E[g(\varphi_{n}(x),\varphi_{0}(x))-g(\varphi_{n}(x),\varphi_{0}(X))|\varphi_{n}(X)=\varphi_{n}(x)]\right|$
$\leq E[\left|g(\varphi_{n}(x),\varphi_{0}(x))-g(\varphi_{n}(x),\varphi_{0}(X))\right||\varphi_{n}(X)=\varphi_{n}(x)]$
$\leq LE[\left\|\varphi_{0}(x)-\varphi_{0}(X)\right\|_{\mathbb{R}^{d}}|\varphi_{n}(X)=\varphi_{n}(x)].$
On the event $\{\varphi_{n}(X)=\varphi_{n}(x)\}$, we know
$\|\varphi_{0}(x)-\varphi_{0}(X)\|_{\mathbb{R}^{d}}$ $\leq\|\varphi_{0}(x)-\varphi_{n}(x)\|_{\mathbb{R}^{d}}+\|\varphi_{n}(x)-\varphi_{n}(X)\|_{\mathbb{R}^{d}}+\|\varphi_{0}(X)-\varphi_{n}(X)\|_{\mathbb{R}^{d}}$
$\leq\|\varphi_{0}(x)-\varphi_{n}(x)\|_{\mathbb{R}^{d}}+\|\varphi_{0}(X)-\varphi_{n}(X)\|_{\mathbb{R}^{d}}.$
Therefore,
$\big{|}g(\varphi_{n}(x),\varphi_{0}(x))-$ $E[g(\varphi_{n}(X),\varphi_{0}(X))|\varphi_{n}(X)=\varphi_{n}(x)]\big{|}$
$\lesssim E[\left\|\varphi_{0}(x)-\varphi_{0}(X)\right\|_{\mathbb{R}^{d}}\mid\varphi_{n}(X)=\varphi_{n}(x)]$
$\lesssim E[\|\varphi_{0}(x)-\varphi_{n}(x)\|_{\mathbb{R}^{d}}]+E[\|\varphi_{0}(X)-\varphi_{n}(X)\|_{\mathbb{R}^{d}}\mid\varphi_{n}(X)=\varphi_{n}(x)].$
Now, for some function $f$, suppose that $(\varphi_{n}(x),\varphi_{0}(x))\mapsto(\Pi_{\varphi_{n},\varphi_{0}}f)(x)$ is Lipschitz continuous. Then, defining $g:(\widehat{m},m)\mapsto E_{0}[f(X)|\varphi_{n}(X)=\widehat{m},\varphi_{0}(X)=m,\mathcal{D}_{n}]$ and noting by the law of iterated expectation that $\Pi_{\varphi_{n}}\Pi_{\varphi_{n},\varphi_{0}}f=\Pi_{\varphi_{n}}f$, we obtain the following pointwise error bound:
$\left|\Pi_{\varphi_{n},\varphi_{0}}f-\Pi_{\varphi_{n}}f\right|\lesssim\left\|\varphi_{n}-\varphi_{0}\right\|_{\mathbb{R}^{d}}+\Pi_{\varphi_{n}}(\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}).$
Since $\|\Pi_{\varphi_{n}}(\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}})\|_{L^{2}(P_{0})}\leq\|\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}\|_{L^{2}(P_{0})}$ by the properties of projections, it follows that
$\|\Pi_{\varphi_{n},\varphi_{0}}f-\Pi_{\varphi_{n}}f\|_{L^{2}(P)}\lesssim\|\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}\|_{L^{2}(P)}.$
Taking $f:=f_{(\varphi_{n},\varphi_{0})}$ and noting that $\Pi_{\varphi_{n},\varphi_{0}}f_{(\varphi_{n},\varphi_{0})}:=f_{(\varphi_{n},\varphi_{0})}$ and that $\Pi_{\varphi_{n}}f_{(\varphi_{n},\varphi_{0})}:=f_{(\varphi_{n})}$, we conclude that
$\|f_{(\varphi_{n},\varphi_{0})}-f_{\varphi_{n}}\|_{L^{2}(P)}\lesssim\|\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}\|_{L^{2}(P)}.$
By an symmetric argument, swapping $\varphi_{n}$ with $\varphi_{0}$, we conclude that
$\|f_{(\varphi_{n},\varphi_{0})}-f_{\varphi_{0}}\|_{L^{2}(P)}\lesssim\|\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}\|_{L^{2}(P)}.$
Hence, by the triangle inequality, we have that
$\|f_{\varphi_{n}}-f_{\varphi_{0}}\|_{L^{2}(P)}\lesssim\|\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}\|_{L^{2}(P)}.$
H.2 Proofs of main results
Proof of Theorem 5.
We introduce the following notation. For any feature transformation $\phi:\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}^{m}$, let $d_{0,\phi}=T_{\phi}(\alpha_{0,\phi})$, where $\alpha_{0,\phi}:=\operatorname*{argmin}_{\alpha\in\overline{H}_{\phi}}E_{0}\left[\{\mathcal{T}_{0,\phi}(f\circ\phi)(A_{0},S_{0})\}^{2}-2m(S_{0},A_{0},\alpha)\right]$ is the Riesz representer for the function class $H_{\phi}:=\{f\circ\phi;f:\mathbb{R}^{m}\rightarrow\mathbb{R}\}\cap L^{\infty}(\lambda)$ induced by $\phi$. Denote $\mu_{0,\phi}$ by the map $(a,s)\mapsto E_{0}[Y_{0}\mid\phi(A_{0},S_{0})=\phi(a,s)]$. Let $\widetilde{\alpha}_{0,q_{n}}:=\operatorname*{argmin}_{f\in\overline{H}_{q_{n}}}\left\|\mathcal{T}_{0,(q_{n},q_{0})}(\alpha_{0,(q_{n},q_{0})})-\mathcal{T}_{0,q_{n}}(f\circ q_{n})\right\|_{P_{0}}$. Let $\widetilde{d}_{0,q_{n}}$ be either $\mathcal{T}_{0,q_{n}}(\alpha_{0,q_{n}})$ or $\mathcal{T}_{0,q_{n}}(\widetilde{\alpha}_{0,q_{n}})$.
By Riesz representation theorem, we have that:
$\Psi_{n}(P_{0})-\Psi(P_{0})$ $=\langle d_{0,(q_{n},q_{0})},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}$
$=\langle d_{0,(q_{n},q_{0})}-\widetilde{d}_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}$
$\quad+\langle\widetilde{d}_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}.$
Note $\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})=\mu_{0}$ by correct specification of $H_{(q_{n},q_{0})}$ for $q_{0}$. Thus, by the law of total expectation, we have
$\langle\widetilde{d}_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}$ $=\langle\widetilde{d}_{0,q_{n}},\mathcal{T}_{0,q_{n}}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}$
$=\langle\widetilde{d}_{0,q_{n}},\mathcal{T}_{0,q_{n}}(q_{0,q_{n}})-\mu_{0}\rangle_{P_{0}}$
$=0,$
where the final equality follows from the orthogonality conditions of the Bellman projection $q_{0,q_{n}}$ of $q_{0}$ with respect to the norm $\|\mathcal{T}_{0,q_{n}}(\cdot)\|_{P_{0}}$ and the fact that $\widetilde{d}_{0,q_{n}}$ is in the closed range of $\mathcal{T}_{0,q_{n}}$. Hence,
$\Psi_{n}(P_{0})-\Psi(P_{0})$ $=\langle d_{0,(q_{n},q_{0})}-\widetilde{d}_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}.$
Moreover, by the Cauchy–Schwarz inequality, it holds that
$|\Psi_{n}(P_{0})-\Psi(P_{0})|$ $\leq\|d_{0,(q_{n},q_{0})}-\widetilde{d}_{0,q_{n}}\|_{P_{0}}\cdot\|\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\|_{P_{0}}.$
For the second part of the theorem, take $\widetilde{d}_{0,q_{n}}$ equal to $\mathcal{T}_{0,q_{n}}(\widetilde{\alpha}_{0,q_{n}})$. By the invertibility condition in D1, $\mathcal{T}_{0,q_{n}}(\overline{H}_{q_{n}})$ equals $\overline{H}_{q_{n}}$ from the first part of (D2). Hence, we have that $\widetilde{d}_{0,q_{n}}$ equals the projection $\operatorname*{argmin}_{d\in\overline{H}_{q_{n}}}\|d_{0,(q_{n},q_{0})}-d\|_{P_{0}}.$ Next, using the second part of (D2), we apply Lemma 14 with $Y:=d_{0,(q_{n},q_{0})}$, $\varphi_{n}:=q_{n}$, and $\varphi_{0}:=q_{0}$ to conclude that
$\|d_{0,(q_{n},q_{0})}-\widetilde{d}_{0,q_{n}}\|_{P_{0}}\lesssim\|q_{n}-q_{0}\|_{P_{0}}.$
Thus,
$|\Psi_{n}(P_{0})-\Psi(P_{0})|$ $\leq\|q_{n}-q_{0}\|_{P_{0}}\|\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\|_{P_{0}},$
as desired. ∎
Proof of Lemma 6.
By empirical calibration, for any transformation $f:\mathbb{R}\rightarrow\mathbb{R}$, we have that
$\sum_{i=1}^{n}f(q_{n}(A_{0,i},S_{0,i}))\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}(A_{0,i},S_{0,i})\}=0.$
Taking $f$ such that $f\circ q_{n}=\widetilde{d}_{0,q_{n}}$, we find that
$\sum_{i=1}^{n}\widetilde{d}_{0,q_{n}}(A_{0,i},S_{0,i})\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}(A_{0,i},S_{0,i})\}=0.$
Therefore, the plug-in estimator $\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n})$ is equal to the DRL estimator:
$\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n})+\sum_{i=1}^{n}\widetilde{d}_{0,q_{n}}(A_{0,i},S_{0,i})\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}(A_{0,i},S_{0,i})\}.$
Proof of Theorem 7.
By Lemma 11, $q_{n}^{*}$ is empirically calibrated for $q_{0}$. Thus, by Lemma 6, it holds that $P_{n}\varphi_{n,q_{n}^{*}}=0$ and, therefore,
$\psi_{n}^{*}-\Psi_{q_{n}^{*}}(P_{0})$ $=\psi_{n}^{*}+P_{n}\varphi_{n,q_{n}^{*}}-\Psi_{q_{n}^{*}}(P_{0})$
$=P_{n}\varphi_{0,q_{0}}+(P_{n}-P_{0})\{\varphi_{n,q_{n}^{*}}-\varphi_{0,q_{0}}\}$
$\quad+\psi_{n}^{*}-\Psi_{q_{n}^{*}}(P_{0})+P_{0}\varphi_{n,q_{n}^{*}}.$
We first inspect the term $\psi_{n}^{*}-\Psi_{q_{n}^{*}}(P_{0})+P_{0}\varphi_{n,q_{n}^{*}}$. Note,
$\psi_{n}^{*}-\Psi_{q_{n}^{*}}(P_{0})+P_{0}\varphi_{n,q_{n}^{*}}$ $=P_{0}m(\cdot,q_{n}^{*})-P_{0}m(\cdot,q_{0,q_{n}^{*}})$
[…]
$=\langle\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}}),\mathcal{T}_{0,q_{n}^{*}}(q_{n}^{*})-\mathcal{T}_{0,q_{n}^{*}}(q_{0,q_{n}^{*}})\rangle_{P_{0}}$
[…]
where the final equality uses the Riesz representation property of ${d}_{0,q_{n}^{*}}$ and that $q_{n}^{*}\in H_{q_{n}^{*}}$ and $q_{0,q_{n}^{*}}\in H_{q_{n}^{*}}$. Next, note, by the law of iterated expectation, that
[…]
Putting it all together, we find that
$\psi_{n}^{*}-\Psi_{q_{n}^{*}}(P_{0})+P_{0}\varphi_{n,q_{n}^{*}}$ $=\langle{d}_{0,q_{n}^{*}},\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{n}^{*})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}})\rangle_{P_{0}}$
$\quad+\langle{d}_{0,q_{n}^{*}},\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{n}^{*})\rangle_{P_{0}}$
$=0.$
Using that $\psi_{n}^{*}-\Psi_{q_{n}^{*}}(P_{0})+P_{0}\varphi_{n,q_{n}^{*}}=0$, we find that
$\psi_{n}^{*}-\Psi_{q_{n}^{*}}(P_{0})$ $=P_{n}\varphi_{0,q_{0}}+(P_{n}-P_{0})\{\varphi_{n,q_{n}^{*}}-\varphi_{0,q_{0}}\}$
$=P_{n}\varphi_{0,q_{0}}+o_{p}(n^{-\frac{1}{2}}),$
where we used that $(P_{n}-P_{0})\{\varphi_{n,q_{n}^{*}}-\varphi_{0,q_{0}}\}=o_{p}(n^{-\frac{1}{2}})$ by D4.
Finally, applying Theorem 5 and D3, we find that
$\Psi_{q_{n}}(P_{0})-\Psi(P_{0})=O_{p}\left(\|q_{n}^{*}-q_{0}\|_{P_{0}}\|\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0})\|_{P_{0}}\right)=o_{p}(n^{-\frac{1}{2}}).$
Consequently,
$\psi_{n}^{*}-\Psi(P_{0})$ $=P_{n}\varphi_{0,q_{0}}+o_{p}(n^{-\frac{1}{2}}).$
Thus, $\psi_{n}^{*}$ is an asymptotically linear estimator of $\Psi(P_{0})=\Psi_{q_{0}}(P_{0})$ with influence function given by the $P_{0}$-efficient influence function of $\Psi_{q_{0}}$. It follows that $\psi_{n}^{*}$ is a regular and efficient estimator for $\Psi_{q_{0}}$ at $P_{0}$. The result then follows.
Appendix I Proofs for ADML in Section B
Proof of Theorem 4.
By C1 applied with $H:=H_{n,0}$ and Riesz representation theorem, we have that
$\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ $=\langle\mathcal{T}_{0}(\alpha_{0,H_{n,0}}),\mathcal{T}_{0}(q_{0,H_{n}})\rangle_{P_{0}}-\langle\mathcal{T}_{0}(\alpha_{0,H_{n,0}}),\mathcal{T}_{0}(q_{0})\rangle_{P_{0}}$
$=\langle\mathcal{T}_{0}(\alpha_{0,H_{n,0}}),\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\rangle_{P_{0}}.$
Note that $\mathcal{T}_{0}(q_{0,H_{n}})$ is the orthogonal projection in $L^{2}(P)$ of $\mathcal{T}_{0}(q_{0})$ onto $\mathcal{T}_{P}(H_{n})$. The orthogonality conditions of the projection imply that
$\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ $=\langle\mathcal{T}_{0}(\alpha_{0,H_{n,0}})-\mathcal{T}_{0}(\alpha_{0,H_{n}}),\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\rangle_{P_{0}}$
$=-\langle\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n,0}}),\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\rangle_{P_{0}}.$
In the event $H_{n}\subseteq H_{0}$, we have that $H_{n,0}=H_{0}$ and, hence,
$\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ $=-\langle\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{0}}),\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\rangle_{P_{0}},$
as desired. ∎
I.1 Proof of Theorem 10
Proof of Theorem 10.
Note that
$\psi_{n,H_{n}}-\Psi_{H_{0}}(P_{0})=\psi_{n,H_{n}}-\Psi_{H_{n}}(P_{0})+\Psi_{H_{n}}(P_{0})-\Psi_{H_{0}}(P_{0}).$
Observe that $\psi_{n,H_{n}}=\Psi_{H_{n}}(\widehat{P_{n}})+P_{n}\varphi_{n,H_{n}}$, where $\widehat{P}_{n}\in\mathcal{P}$ is any distribution such that $q_{\widehat{P}_{n},H_{n}}=q_{n,H_{n}}$, $\mu_{\widehat{P}_{n}}=\widehat{\mathcal{T}}_{n}(q_{n,H_{n}})$, and $T_{\widehat{P}_{n}}(\alpha_{\widehat{P}_{n},H_{n}})=T_{\widehat{P}_{n}}(\alpha_{n,H_{n}})$. Thus, it holds that:
$\psi_{n,H_{n}}-\Psi_{H_{n}}(P_{0})$ $=\Psi_{H_{n}}(\widehat{P_{n}})+P_{n}\varphi_{n,H_{n}}-\Psi_{H_{n}}(P_{0})$
$=P_{n}\varphi_{0,H_{n}}+(P_{n}-P)\{\varphi_{n,H_{n}}-\varphi_{0,H_{n}}\}+R_{n,H_{n}}(P_{0}),$
where $R_{n,H_{n}}(P_{0})=\Psi_{H_{n}}(\widehat{P}_{n})-\Psi_{H_{n}}(P_{0})+P_{0}\varphi_{n,H_{n}}$. By a direct application of C4, we have that $(P_{n}-P)\{\varphi_{n,H_{n}}-\varphi_{0,H_{n}}\}=o_{p}(n^{-\frac{1}{2}})$. Moreover, by application of C6,
$P_{n}\varphi_{0,H_{n}}$ $=(P_{n}-P_{0})\varphi_{0,H_{n}}$
$=(P_{n}-P_{0})\varphi_{0,H_{0}}+(P_{n}-P_{0})\{\varphi_{n,H_{n}}-\varphi_{0,H_{n}}\}$
$=P_{n}\varphi_{0,H_{0}}+o_{p}(n^{-\frac{1}{2}}),$
where we used that $P_{0}\varphi_{0,H_{n}}=0$ and $P_{0}\varphi_{0,H_{0}}=0$. Thus,
$\psi_{n,H_{n}}-\Psi_{H_{n}}(P_{0})$ $=P_{n}\varphi_{0,H_{0}}+R_{n,H_{n}}(P_{0})+o_{p}(n^{-\frac{1}{2}}).$
Next, applying Theorem 2, we find that
$R_{n,H_{n}}(P_{0})$ $=E_{0}\left[\left\{\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{n,H_{n}})\right\}(\mathcal{T}_{0}(q_{0})-\widehat{\mathcal{T}}_{n}(q_{n,H_{n}})\right]$
$\quad+E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\}(\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{n,H_{n}}))\right]$
$\quad+E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})\}(\mathcal{T}_{0}(q_{n,H_{n}})-\widehat{\mathcal{T}}_{n}(q_{n,H_{n}}))\right]$
$=E_{0}\left[\left\{\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{n,H_{n}})\right\}(\mathcal{T}_{0}(q_{0})-\mathcal{T}_{0}(q_{n,H_{n}})\right]$
$\quad+E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\}(\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{n,H_{n}}))\right].$
Since $\mathcal{T}_{0}(q_{0,H_{n}})$ is the $L^{2}(P_{0})$ projection of $q_{0}$ onto $\mathcal{T}_{0}(H_{n})$, we have the orthogonality conditions:
$E_{0}\left[\left\{\mathcal{T}_{0}(\alpha)\right\}(\mathcal{T}_{0}(q_{0})-\mathcal{T}_{0}(q_{0,H_{n}})\right]=0\text{ for all }\alpha\in H_{n}.$
Hence, since $\alpha_{n,H_{n}},\alpha_{0,H_{n}}\in H_{n}$, we have that
$E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\}(\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{n,H_{n}}))\right]=E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\}(\mathcal{T}_{0}(q_{0})-\mathcal{T}_{0}(q_{n,H_{n}}))\right].$
Substituting the above expression, we find that
$R_{n,H_{n}}(P_{0})$ $=E_{0}\left[\left\{\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{n,H_{n}})\right\}(\mathcal{T}_{0}(q_{0})-\mathcal{T}_{0}(q_{n,H_{n}})\right]$
$\quad+E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\}(\mathcal{T}_{0}(q_{0})-\mathcal{T}_{0}(q_{n,H_{n}}))\right]$
$=E_{0}\left[\left\{\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\right\}(\mathcal{T}_{0}(q_{0})-\mathcal{T}_{0}(q_{n,H_{n}})\right]$
$=O_{p}\left(\|\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\|_{P_{0}}\|\mathcal{T}_{0}(q_{n,H_{n}})-\mathcal{T}_{0}(q_{0,H_{n}})\|_{P_{0}}\right)$
$=o_{p}(n^{-\frac{1}{2}}),$
where the final two equalities follow from the Cauchy-Schwarz inequality and C5. Thus,
$\psi_{n,H_{n}}-\Psi_{H_{n}}(P_{0})$ $=P_{n}\varphi_{0,H_{0}}+o_{p}(n^{-\frac{1}{2}}).$
Next we turn to the term $\Psi_{H_{n}}(P_{0})-\Psi_{H_{0}}(P_{0})$. Note, by Theorem 4 and C7, it holds that:
$\Psi_{H_{n}}(P_{0})-\Psi_{H_{0}}(P_{0})$ $=-\langle\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n,0}}),\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\rangle_{P_{0}}$
$=O_{p}(\|\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n,0}})\|_{P_{0}}\|\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\|_{P_{0}})$
$=o_{p}(n^{-\frac{1}{2}}),$
where the final two equalities follow from the Cauchy-Schwarz inequality and C7.
Putting it all together, we conclude that
$\psi_{n,H_{n}}-\Psi_{H_{0}}(P_{0})$ $=P_{n}\varphi_{0,H_{0}}+o_{p}(n^{-\frac{1}{2}}).$
Since $\varphi_{0,H_{0}}$ is the $P_{0}$-efficient influence function of $\Psi_{0}$, it follows that $\psi_{n,H_{n}}$ is an asymptotically linear, regular, and efficient estimator for $\Psi_{0}$ at $P_{0}$.