### v4

# Semiparametric Double Reinforcement Learning with Applications to Long-Term Causal Inference

Abstract
Double Reinforcement Learning (DRL) enables efficient inference for policy values in nonparametric Markov decision processes (MDPs), but existing methods face two major obstacles: (1) they require stringent intertemporal overlap conditions on state trajectories, and (2) they rely on estimating high-dimensional occupancy density ratios. Motivated by problems in long-term causal inference, we extend DRL to a semiparametric setting and develop doubly robust, automatic estimators for general linear functionals of the $Q$-function in infinite-horizon, time-homogeneous MDPs. By imposing structure on the $Q$-function, we relax the overlap conditions required by nonparametric methods and obtain efficiency gains. The second obstacle—density-ratio estimation—typically requires computationally expensive and unstable min–max optimization. To address both challenges, we introduce superefficient nonparametric estimators whose limiting variance falls below the generalized Cramér–Rao bound. These estimators treat the $Q$-function as a one-dimensional summary of the state–action process, reducing high-dimensional overlap requirements to a single-dimensional condition. The procedure is simple to implement: estimate and calibrate the $Q$-function using fitted $Q$-iteration, then plug the result into the target functional, thereby avoiding density-ratio estimation altogether.
Keywords— Double reinforcement learning, semiparametric inference, policy evaluation, long-term causal inference, automatic debiasing, superefficiency, calibration
1 Introduction and motivation
Randomized experiments such as A/B tests and clinical trials are widely used to assess the effects of interventions. In industry, econometrics, and clinical applications, these experiments are often short-term, even though the decisions they inform involve long-term outcomes such as survival, customer retention, or revenue. This gap has motivated a central question in recent research: How can we estimate the long-term causal effect of an intervention when only short-term experiments are available? (athey2019surrogate; van2023estimating; tran2023inferring; bibaut2023long; nam2024predicting; imbens2025long)
A common strategy is to use surrogate outcomes—intermediate variables observed during the experiment that are assumed to mediate the long-term effect (prentice1989surrogate; frangakis2002principal; athey2019surrogate). These methods combine short-term experimental data with long-term observational data by treating the surrogate as a proxy for the eventual outcome. For example, a streaming platform might use one-month watch time in an A/B test as a surrogate for annual subscription retention recorded in historical data. A key limitation of surrogate-based approaches is that the short-term surrogate must fully mediate the long-term treatment effect (vanderweele2013surrogate; weintraub2015perils; stijven2025evaluation). This assumption fails in settings where treatment effects accumulate over time (tran2023inferring). For instance, annual retention under a new recommendation algorithm depends on evolving user behavior that cannot be captured by a single short-term surrogate.
Figure 1: DAG for Markov Decision Process. $Y_{1}$ and $S_{2}$ need not be observed in the experiment.
These limitations motivate dynamic models that capture how long-term outcomes evolve in response to sequences of states, actions, and rewards under sustained treatment. To this end, tran2023inferring proposed estimating long-term effects from short-term experiments using a time-homogeneous Markov Decision Process (MDP) (Figure 1). Unlike surrogate-based methods, the MDP approach does not require the long-term effect to be mediated by a single short-term variable; instead, it models the full state–action trajectory. Earlier work in this direction includes battocchi2021estimating, liao2021off, and Example 2 of bibaut2021sequential. These approaches link long-term causal inference to offline reinforcement learning (kaelbling1996reinforcement) and enable efficient nonparametric estimation via Double Reinforcement Learning (DRL) (van2018online; kallus2020double; kallus2022efficiently).
DRL enables statistically efficient inference on policy values in nonparametric MDPs using off-policy data, but it faces two major practical challenges. First, DRL requires sufficient intertemporal overlap in state trajectories, meaning that the distribution of states observed early in the trajectory has support over the states that determine long-term outcomes. When this overlap is weak, DRL estimators become unstable and require prohibitively large samples—a difficulty that is especially severe in high-dimensional or unbounded state spaces (mehrabi2024off). Unlike standard policy overlap, intertemporal overlap is more restrictive because future states cannot be randomized: even perfect treatment randomization does not guarantee support in the state distribution over time. A closely related issue is well known in offline RL, where short-horizon behavior data do not cover the long-term state distribution induced by the target policy, leading to extrapolation error and instability under temporal distribution shift (buckman2020importance; szepesvari2022algorithms). Although value estimation under limited overlap has been widely studied, valid statistical inference in such settings remains largely unresolved (fujimoto2019off; levine2020offline; jin2021pessimism).
Second, DRL requires estimating high-dimensional nuisance components such as occupancy density ratios, typically via min–max optimization over adversarial function classes. State-of-the-art methods include DICE (nachum2019dualdice; zhang2020gendice) and minimax weight estimation approaches (liu2018breaking; uehara2020minimax; dikkala2020minimax). In practice, these adversarial estimators are computationally demanding, tuning-sensitive, and often unstable in finite samples (jiang2020minimax; kallus2020double). A common workaround is to restrict the function class used in the inner maximization (e.g., to linear or RKHS models), but this induces misspecification bias (kallus2020double; kim2022lobsdice; olivas2025source).
1.1 Contributions of this work
Motivated by practical challenges, we extend the DRL framework to incorporate semiparametric model restrictions and introduce a new class of superefficient estimators that address the twin challenges of limited intertemporal overlap and high-dimensional nuisance estimation. Our three main contributions are as follows.
Semiparametric and automatic inference for DRL. We develop a semiparametric framework for estimation and inference in DRL under model restrictions on the $Q$-function in time-homogeneous MDPs. The $Q$-function is the dynamic analogue of the outcome regression, linking long-term cumulative rewards to short-term data through the Bellman equation (bellman1952theory). By imposing structure on the $Q$-function, our approach relaxes the stringent intertemporal overlap conditions required by fully nonparametric methods and yields improved statistical efficiency. Related gains have been shown in the special case of linear MDPs, where both the transition kernel and reward function are assumed linear in a known feature map (xie2023semiparametrically). In contrast, we show that similar benefits arise under substantially weaker assumptions: it is sufficient to restrict only the $Q$-function, without imposing a parametric model on the underlying MDP. The intuition is that any restriction on the $Q$-function implicitly constrains both the reward function and the transition dynamics through the Bellman equation. Beyond policy values, we consider inference for general $L^{2}$-continuous linear functionals of the $Q$-function, and show that debiasing requires estimating only one additional nuisance component—the corresponding Riesz representer, which generalizes the discounted state–action occupancy ratio. We construct automatic DRL estimators whose form does not depend on the target functional and which are doubly robust to misspecification of both the $Q$-function and its Riesz representer.
Superefficient nonparametric estimators under limited overlap. When model restrictions are misspecified, DRL estimators may introduce bias, whereas fully nonparametric methods remain valid but can become unstable, highly variable, or nonregular under limited intertemporal overlap (imbensOverlapEstimand2006; benkeser2020nonparametric; mehrabi2024off). To address this, we develop a new class of superefficient nonparametric estimators for continuous linear functionals of the $Q$-function. These estimators are asymptotically normal with variance strictly below the generalized Cramér–Rao bound, enabling inference in irregular settings where root-$n$ regular estimation is infeasible, such as when occupancy ratios have infinite variance (mehrabi2024off). They achieve this by using the $Q$-function as a one-dimensional summary of the state–action space, thereby reducing high-dimensional overlap requirements to a single-dimensional condition.
Calibration removes the need to estimate density ratios. Existing DRL estimators rely on adversarial min–max optimization to estimate the discounted occupancy ratio, a step that is computationally intensive, tuning-sensitive, and unstable. Our proposed superefficient estimator bypasses this step entirely by using a calibrated plug-in estimator. The method is simple: estimate the $Q$-function using fitted $Q$-iteration, calibrate it, and plug the result into the target functional. We show that calibration implicitly carries out the necessary bias correction, obviating the need to estimate density ratios or, more broadly, the Riesz representer. Unlike existing DRL estimators, our approach requires only fitted $Q$-iteration (munos2008finite), making nuisance estimation computationally efficient and stable, and supported by well-developed theory (fan2020theoretical; hu2025fast; munos2008finite). To implement this idea, we introduce fitted $Q$-calibration, a dynamic extension of isotonic calibration (zadrozny2002transforming; niculescu2005predicting).
This paper is organized as follows. Section 2 introduces Markov decision processes, $Q$-functions, and the target estimand. Section 3 presents our semiparametric DRL estimators, the corresponding asymptotic theory, and extensions. In Section 4, we propose our superefficient estimator and fitted $Q$-calibration. Finally, Section 5 presents numerical experiments.
1.2 Related work
Off-policy evaluation. Our work builds on recent advances in inference for off-policy evaluation (murphy2003optimal; liu2018breaking; tang2019doubly; shi2022statistical; wang2023projected), particularly debiased and doubly robust estimation methods (tang2019doubly; kallus2020double; shi2021deeply; kallus2022efficiently; mehrabi2024off). Existing approaches typically model the $Q$-function using nonparametric or parametric methods and focus on a single linear functional—the policy value. Parametric inference in linear MDPs has been studied by xie2023semiparametrically and wu2024statistical. We extend this literature by developing debiased inference procedures for general linear functionals of the $Q$-function under semiparametric restrictions. We focus on time-homogeneous MDPs, which are especially relevant for long-term causal inference and reinforcement learning, and which avoid the curse of the infinite horizon that arises in non-homogeneous models (liu2018breaking; kallus2022efficiently).
Addressing intertemporal overlap. Early work on nonparametric inference in off-policy evaluation required strong intertemporal overlap, assuming the occupancy density ratio was uniformly bounded (van2018online; kallus2020double; kallus2022efficiently). Later work relaxed this to a finite-variance assumption via ratio truncation (mehrabi2024off), but fully nonparametric estimators can still require large samples when overlap is weak, and the ratio itself may have infinite variance or fail to exist, making root-$n$ regular inference impossible. In contrast, we address overlap through model restrictions, and our superefficient estimator enables valid nonparametric inference even when the density ratio is ill-behaved. In value estimation and policy optimization, limited overlap is often handled by regularizing value estimates in poorly supported regions (yu2020mopo; liu2020provably; jin2021pessimism), but these methods do not extend naturally to statistical inference. Our use of semiparametric structure achieves a similar form of extrapolation from well-supported to weakly supported regions, but is designed for valid inference rather than policy optimization.
Semiparametric restrictions and efficiency. Semiparametric restrictions have long been used to stabilize causal inference in cross-sectional settings. For example, while fully nonparametric estimators such as AIPW (robinsCausal) may suffer from high variance under limited treatment overlap, estimators based on partially linear models are far less sensitive to overlap violations (imbensOverlapEstimand2006; li2019overlapWeights; robinson1988root). Related ideas appear in data fusion, where model restrictions are used to address unobserved confounding (kallus2018removing), and in randomized trials, where such restrictions improve efficiency by leveraging external data (van2024adaptive).
Superefficient estimation. Superefficient ATE estimators have been developed to permit nonparametric inference under limited overlap (CTMLElasso; d2021deconfounding; moosavi2023costs; van2023adaptive), typically through data-driven confounder selection or dimension reduction (shortreed2017outcome; HALoutcomeAdapt; benkeser2020nonparametric). Our work is closest to benkeser2020nonparametric, who use the outcome regression function as a one-dimensional summary for adjustment. We extend this idea to dynamic settings, showing that the $Q$-function plays an analogous role in MDPs. We achieve debiasing via isotonic calibration rather than targeted minimum loss estimation, which simplifies the procedure and avoids additional nuisance estimation.
Calibration and isotonic regression. Calibration has a long history in forecasting and machine learning as a tool for improving the reliability of classification and prediction models (lichtenstein1977calibration; platt1999probabilistic; zadrozny2001obtaining). Isotonic calibration (zadrozny2002transforming; niculescu2005predicting) is a widely used instance based on isotonic regression (barlow1972isotonic). In causal inference, calibration has primarily been applied to improve nuisance estimators, such as CATEs and propensity scores (gutman2022propensity; van2023causal; ballinari2024improving; van2024stabilized; rabenseifner2025calibrationstrategiesrobustcausal). This literature is distinct from causal isotonic regression (westling2020causal), a shape-constrained method for inference on monotone dose–response curves.
In contrast, we use calibration not as a nuisance-refinement step but as a core ingredient for inference. Our approach is motivated by the calibrated DML framework of vanderlaan2024automatic, which employs calibration to achieve doubly robust inference for linear functionals of a regression function, yielding estimators that remain asymptotically normal even when one of the nuisance components is inconsistently estimated. While we extend this idea to the MDP setting, our objective differs. Rather than targeting double robustness, we use calibration to construct a super-efficient plug-in estimator that avoids estimation of the Riesz representer. The dynamic setting also introduces new challenges: the $Q$-function is defined as the solution to a linear inverse problem, requiring new calibration techniques and corresponding theory. To our knowledge, calibration of $Q$-functions in MDPs has not previously been studied, and our dynamic generalization of isotonic calibration is novel and of independent methodological interest.
2 Preliminaries
2.1 Data Structure and Markov Decision Model
We consider a randomized experiment or observational study in which individuals sequentially receive treatments (actions) according to a policy. At each time $t\in\{0,1,2,\dots\}$, an individual occupies a state $S_{t}\in\mathcal{S}$, which informs the choice of action $A_{t}\in\mathcal{A}$. The resulting state–action pair generates an intermediate outcome $Y_{t}\in\mathbb{R}$, interpreted as an immediate reward or cost, and determines the next state $S_{t+1}$. The observed data consist of $n$ i.i.d. samples of a single transition, $\mathcal{D}_{n}:=\{(S_{0,i},A_{0,i},Y_{0,i},S_{1,i}):i\in[n]\},$ drawn from a distribution $P_{0}$ in a nonparametric model $\mathcal{P}$. Although only one transition is observed per individual, datasets with multiple transitions can be decomposed into such units, and any remaining dependence can be handled using Markov chain limit theory (bibaut2021sequential). The goal is to perform long-term causal inference for a counterfactual policy $\pi(a\mid s)$, assumed stationary and admitting a density with respect to a measure $\nu$ on $\mathcal{A}$. For notational convenience, we write $f_{0}$ for any population quantity $f_{P_{0}}$ defined under $P_{0}$.
The infinite-horizon trajectory $(S_{0},A_{0},Y_{0},S_{1},A_{1},Y_{1},S_{2},A_{2},\dots)$ is generated by a time-homogeneous Markov decision process with joint law $\mathbb{P}_{0}$ (puterman1990markov). Formally, we represent the data-generating mechanism using the nonparametric structural equation model (NPSEM) (pearl2012causal):
$A_{t}$ $=f_{A}(S_{t},U_{A_{t}}),\qquad Y_{t}$ $=f_{Y}(A_{t},S_{t},U_{Y_{t}}),\qquad S_{t+1}$ $=f_{S}(Y_{t},A_{t},S_{t},U_{S_{t+1}}),$
where $f_{A}$, $f_{Y}$, and $f_{S}$ are unknown deterministic functions, and the latent variables $\{U_{A_{t}},U_{Y_{t}},U_{S_{t+1}}\}$ are mutually independent, unobserved, and stationary. This encoding implies the Markov property: $A_{t}$ depends on the past only through $S_{t}$, $Y_{t}$ depends only on $(A_{t},S_{t})$, and $S_{t+1}$ depends only on $(Y_{t},A_{t},S_{t})$. We also assume time-homogeneity, so the conditional laws of $Y_{t}$ and $S_{t+1}$ do not depend on $t$. Under these assumptions, the full trajectory law $\mathbb{P}_{0}$ is determined by the distribution $P_{0}$ of the first transition. These assumptions become more plausible when the state space is sufficiently rich—for example, by augmenting the state with a fixed number of past values, $\widetilde{S}_{t}:=(S_{t},S_{t-1},\dots,S_{t-k})$, or with low-dimensional summaries of historical information (van2018robust). The NPSEM allows us to define counterfactual MDPs, as illustrated in the next example.
Example 1 (Policy Value in an MDP).
Given a policy $\pi$, let $\{S_{t}(\pi),A_{t}(\pi),Y_{t}(\pi):t\in\mathbb{T}\}\sim\mathbb{P}_{0}^{\pi}$ denote the counterfactual trajectory under a Markov decision process, where $A_{t}$ is drawn from $\pi(\cdot\mid S_{t})$. This trajectory is defined by intervening on the structural equation for $A_{t}$ in the NPSEM so that $A_{t}^{\prime}$ follows $\pi(\cdot\mid S_{t})$. The value of the policy at time $t$ is $\mathbb{E}_{0}^{\pi}[Y_{t}(\pi)]$, and the expected discounted cumulative outcome, for discount factor $\gamma\in[0,1]$, is $\mathbb{E}_{0}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}Y_{t}(\pi)\right]$, where $\mathbb{E}_{0}^{\pi}$ denotes the expectation under $\mathbb{P}_{0}^{\pi}$. ∎
2.2 Inferential objective
Our goal is to perform inference on $L^{2}$-continuous linear functionals of the $Q$-function, which links long-term causal effects to short-term data. For a given discount factor $\gamma\in[0,1]$, the $Q$-function associated with a policy $\pi$ is defined as
$q_{0}^{\pi}(a,s):=\mathbb{E}_{0}^{\pi}\!\left[\sum_{t=0}^{\infty}\gamma^{t}Y_{t}(\pi)\mid A_{0}=a,S_{0}=s\right],$
where $\mathbb{E}_{0}^{\pi}$ denotes expectation under the counterfactual MDP induced by policy $\pi$ (kaelbling1996reinforcement). Throughout, when the policy is $\pi$, we omit the superscript and write $q_{0}$. The $Q$-function represents the expected cumulative reward for an individual starting in state $s$, taking action $a$, and subsequently following policy $\pi$. The discount factor $\gamma$ controls the weight placed on future rewards. Under the convention $0^{0}=1$, setting $\gamma=0$ reduces $q_{0}(a,s)$ to the reward function $r_{0}(a,s):=E_{0}[Y_{0}\mid A_{0}=a,S_{0}=s]$.
Our estimand is $\psi_{0}:=E_{0}[m(S_{0},A_{0},q_{0})]$, where $q\mapsto m(S_{0},A_{0},q)$ is a linear map. A key example is the long-term policy value $\mathbb{E}_{0}^{\pi}\!\left[\sum_{t=0}^{\infty}\gamma^{t}Y_{t}(\pi)\right],$ which corresponds to $m(s,a,q)=\int q(a^{\prime},s)\,\pi(a^{\prime}\mid s)\,\nu(da^{\prime})$. The $Q$-function $q_{0}$ is identified from the observed data distribution $P_{0}$ as the fixed point of the Bellman equation (bellman1966dynamic; sutton1998reinforcement):
$q_{0}(A_{0},S_{0})=E_{0}\!\left[Y_{0}+\gamma V^{\pi}(q_{0})(S_{1})\mid A_{0},S_{0}\right],$ (1)
where $V^{\pi}(q)(s^{\prime}):=\int q(a^{\prime},s^{\prime})\,\pi(a^{\prime}\mid s^{\prime})\,\nu(da^{\prime})$ denotes the value function. The Bellman equation expresses $q_{0}(A_{0},S_{0})$ as the sum of the immediate reward and the discounted future value $V^{\pi}(q_{0})(S_{1})$ under policy $\pi$. Setting $\gamma=0$ recovers the well-studied problem of inference on linear functionals of the outcome regression (chernozhukov2018double; chernozhukov2022automatic). A notable example of such a functional is the long-term causal effect in a randomized experiment (tran2023inferring).
Example 2 (Long-term causal effect in an A/B test).
Consider an A/B test in which individuals are randomly assigned to treatment or control, so the behavior policy is static: $A_{t}:=Z$ for all $t\geq 0$, where $Z\in\{0,1\}$ is the treatment indicator. Following tran2023inferring, the long-term causal effect is defined as $\psi_{0}:=\mathbb{E}_{0}[\sum_{t=0}^{\infty}\gamma^{t}\{Y_{t}(1)-Y_{t}(0)\}]$, where $Y_{t}(1)$ and $Y_{t}(0)$ denote the potential outcomes under treatment and control. Equivalently, $\psi_{0}=\psi_{0,1}-\psi_{0,0}$, with $\psi_{0,z}:=\mathbb{E}_{0}[q_{0}^{\pi_{z}}(z,S_{0})]$ for the deterministic policy $\pi_{z}(a\mid s)=\mathbbm{1}\{a=z\}$. Each component $\psi_{0,z}$ corresponds to the linear functional $m(s,a,q)=q(z,s)$. ∎
Challenges in nonparametric inference. Statistically efficient estimation of the policy value $E_{0}[V^{\pi}(q_{0})(S_{0})]$—a special case of a linear functional—via DRL has been studied under nonparametric models by van2018online, kallus2020double; kallus2022efficiently, and tran2023inferring. However, nonparametric inference is challenging because root-$n$ regular estimation requires the existence and finite variance of the unnormalized state–action discounted occupancy ratio
$d_{0}(a,s):=\frac{\pi(a\mid s)}{b_{0}(a\mid s)}\sum_{t=0}^{\infty}\gamma^{t}\frac{d\mathbb{P}_{0}^{\pi}(S_{t}=s)}{d\mathbb{P}_{0}^{b_{0}}(S_{0}=s)},$ (2)
where $b_{0}(a\mid s):=P_{0}(A_{0}=a\mid S_{0}=s)$ is the behavior policy (mehrabi2024off). This condition requires both policy overlap and intertemporal overlap, which can be restrictive in high-dimensional settings or when interventions induce rare or novel states (jin2021pessimism; d2021overlap; mehrabi2024off). Near-violations inflate estimator variance because the efficiency bound is determined by the variability of $d_{0}$ (kallus2022efficiently).
2.3 The Bellman equation as an inverse problem
We consider an equivalent representation of $\psi_{0}$ as a linear functional of the solution to a linear inverse problem (ai2003efficient; ai2012semiparametric; bennett2022inference; bennett2023minimax; bennett2023source). Let $\lambda$ be a measure on $\mathcal{S}\times\mathcal{A}$ that dominates the distributions of $(A_{0},S_{0})$ and $(A_{1},S_{1})$ for all $P\in\mathcal{P}$, and let $L^{\infty}(\lambda)$ denote the Banach space induced by the $\lambda$-essential supremum norm. By rearranging terms in (1), the $Q$-function $q_{P}$ for each $P\in\mathcal{P}$ is identified as the solution to the integral equation
$\mathcal{T}_{P}(q_{P})=r_{P},\quad\text{$P$-almost everywhere},$ (3)
where the reward function (or outcome regression) is defined by $r_{P}(a,s):=E_{P}[Y_{0}\mid A_{0}=a,S_{0}=s]$, and the Bellman integral operator $\mathcal{T}_{P}:L^{\infty}(\lambda)\rightarrow L^{\infty}(\lambda)$ is defined as as $\mathcal{T}_{P}(h)(a,s):=h(a,s)-\gamma E_{P}\left[V^{\pi}(h)(S_{1})\mid A_{0}=a,S_{0}=s\right].$ It holds that $\psi_{0}=\Psi(P_{0})$ where we define target parameter $\Psi:P\mapsto E_{P}\left[m(S_{0},A_{0},q_{P})\right]$, over the nonparametric model $\mathcal{P}$.
3 Semiparametric double reinforcement learning
3.1 Proposed estimator
In this section, we propose automatic DRL estimators for the estimand $\psi_{0}$ that impose semiparametric restrictions on the $Q$-function. In later sections, we describe how to modify these estimators to obtain model-robust inference and derive their asymptotic theory.
Suppose the true $Q$-function $q_{0}$ lies in a (possibly infinite-dimensional) subspace $H\subset L^{\infty}(\lambda)$, and define the corresponding semiparametric model $\mathcal{P}_{H}:=\{P\in\mathcal{P}:q_{P}\in H\}$. For example, $H$ may encode a partially linear structure of the form $q_{0}(A_{0},S_{0})=q_{0}(0,S_{0})+A_{0}\beta_{0}^{\top}S_{0}$, where the baseline function $q_{0}(0,S_{0})$ is estimated nonparametrically while the treatment contrast $q_{0}(1,S_{0})-q_{0}(0,S_{0})$ is modeled linearly in $S_{0}$ (robinson1988root). A natural estimator of $\psi_{0}=\Psi(P_{0})$ is the plug-in estimator $\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},q_{n,H})$, where $q_{n,H}\in H$ is an estimator of $q_{0}$. One way to obtain $q_{n,H}$ is through fitted $Q$-iteration, discussed in Section 3.5. However, when $q_{n,H}$ is estimated flexibly, the plug-in estimator generally fails to achieve $\sqrt{n}$-consistency or asymptotic normality because the first-order bias induced by $q_{n,H}-q_{0}$ dominates (van2011targeted; chernozhukov2018double). Debiasing is therefore required to eliminate this first-order bias and recover valid asymptotic inference.
To construct a bias correction, we require that the linear functional $\psi_{0}=E_{0}[m(S_{0},A_{0},q_{0})]$ be continuous in $q_{0}$, and that $q_{0}$ depend continuously on $r_{0}$ through the inverse relation $\mathcal{T}_{0}(q_{0})=r_{0}$. To formalize this, for $P\in\mathcal{P}$, let $\langle\cdot,\cdot\rangle_{P}$ denote the $L^{2}(P_{A_{0},S_{0}})$ inner product—where $P_{A_{0},S_{0}}$ is the marginal distribution of $(A_{0},S_{0})$ under PP—and let $\|\cdot\|_{P}$ be the associated norm. The closure of $H$ with respect to $\|\cdot\|_{P}$ is denoted $\overline{H}_{P}$, and we define $\mathcal{T}_{P}(\overline{H}_{P}):=\{\mathcal{T}_{P}(h):h\in\overline{H}_{P}\}.$ We impose the following conditions.
  1. (C1)
    For all $P\in\mathcal{P}$ in a Hellinger neighborhood of $P_{0}$, the following hold:
    1. (i)
      (Functional continuity) There exists $C<\infty$ such that $|E_{P}[m(S_{0},A_{0},q)]|\leq C\|q\|_{P}$ for all $q\in H$.
    2. (ii)
      (Bounded inverse) The operator $\mathcal{T}_{P}$ is continuous as a map from $(H,\|\cdot\|_{P})$ to $L^{2}(P_{A_{0},S_{0}})$, and its unique extension to $\overline{H}_{P}$ has a continuous inverse on its range.
Condition C1 ensures that $\psi_{0}=E_{0}[m(S_{0},A_{0},\mathcal{T}_{0}^{-1}(r_{0}))]$ is an $L^{2}$-continuous functional of the reward function $r_{0}$ on the range $\mathcal{T}_{0}(\overline{H}_{0})$, a standard requirement for regular estimation (chernozhukov2018learning; ichimura2022influence). Condition C1(ii) is standard in reinforcement learning and follows from the well-posedness of the Bellman equation (chen2022well). Appendix A further shows that $\mathcal{T}_{0}$ is a Fredholm operator of index zero and therefore admits a bounded inverse for all but finitely many discount factors $\gamma\in[0,1)$ (conway1994course).
Condition C1 also implies that the functional $q\mapsto E_{0}[m(S_{0},A_{0},q)]$ is continuous with respect to the weak norm $\|\mathcal{T}_{0}(\cdot)\|_{P_{0}}$ induced by the Bellman operator, in the sense that
$\sup_{q\in H}\frac{|E_{0}[m(S_{0},A_{0},q)]|}{\|\mathcal{T}_{0}(q)\|_{P_{0}}}<\infty.$ (4)
As a consequence, the target parameter $\psi_{0}$ admits a dual representation as a (signed) weighted expectation of the outcome $Y_{0}$. Specifically, by the Riesz representation theorem, there exists a representer $\alpha_{0,H}\in\overline{H}_{P_{0}}$ such that
$\psi_{0}=\langle\mathcal{T}_{0}(\alpha_{0,H}),\mathcal{T}_{0}(q_{0})\rangle_{P_{0}}=\langle\mathcal{T}_{0}(\alpha_{0,H}),r_{0}\rangle_{P_{0}},$
and hence $\psi_{0}=E_{0}[\mathcal{T}_{0}(\alpha_{0,H})(A_{0},S_{0})Y_{0}]$. The weighting function $\mathcal{T}_{0}(\alpha_{0,H})$ generalizes the state–action occupancy ratio $d_{0}$ in (2) and plays a central role in debiasing.
We construct an automatic DRL estimator by augmenting the plug-in estimator with an influence-function-based bias correction (bickel1993efficient). Given estimators $\widehat{\mathcal{T}}_{n}(\alpha_{n,H})$ and $q_{n,H}$ of $\mathcal{T}_{0}(\alpha_{0,H})$ and $q_{0}$, respectively, we define the doubly robust estimator
$\psi_{n,H}:=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},q_{n,H})+\frac{1}{n}\sum_{i=1}^{n}\widehat{\mathcal{T}}_{n}(\alpha_{n,H})(A_{0,i},S_{0,i})\left\{Y_{0,i}+\gamma V^{\pi}(q_{n,H})(S_{1,i})-q_{n,H}(A_{0,i},S_{0,i})\right\}.$ (5)
For the special case of the policy value and $H:=L^{\infty}(\lambda)$, this reduces to the nonparametric DRL estimator of kallus2020double. As in standard debiased machine learning, we recommend cross-fitting for nuisance estimation to mitigate overfitting and avoid Donsker conditions (van2011cross; chernozhukov2018double). For readability, we omit cross-fitting notation throughout.
The estimator (5) is automatic: its form does not depend on the choice of linear functional, but only on the Riesz representer. The representer optimizes the objective:
$\alpha_{0,H}=\arg\min_{\alpha\in\overline{H}_{P_{0}}}E_{0}\!\left[\{\mathcal{T}_{0}(\alpha)(A_{0},S_{0})\}^{2}-2\,m(S_{0},A_{0},\alpha)\right],$ (6)
which can be rewritten as a min–max saddle-point problem that avoids computing the conditional expectation inside $\mathcal{T}_{0}$ (see Section 3.5). Once $\alpha_{0,H}$ is estimated, the quantity $\mathcal{T}_{0}(\alpha_{0,H})$ is obtained via a second-stage regression. As in autoDML (chernozhukov2022automatic; bennett2023source; van2025automatic), the Riesz representer—and therefore the DRL estimator—can be learned using flexible machine learning methods for any choice of $m$.
3.2 Model-robust extension and asymptotic theory
Under semiparametric restrictions, the true $Q$-function $q_{0}$ may not lie exactly in the model $H\subset L^{\infty}(\lambda)$, but it may be well approximated by elements of $H$. Allowing small misspecification in exchange for efficiency gains can be advantageous, especially under limited overlap (imbensOverlapEstimand2006; van2023adaptive). For instance, linear MDP models yield substantial efficiency improvements (wang2020statistical; xie2023semiparametrically; wu2024statistical). More generally, $H$ may be adopted purely as a working model, with interest focused on the best approximation to $q_{0}$, such as the best linear predictor when $H$ is taken to be a linear class (whitney2020comment; vansteelandt2022assumption; chambaz2012estimation; chernozhukov2018generic).
To accommodate such settings, we extend our framework to model-robust inference on the projection estimand $\psi_{0,H}:=E_{0}\!\left[m(S_{0},A_{0},q_{0,H})\right],$ where $q_{0,H}$ is the Bellman projection of $q_{0}$ onto $H$. For any $P\in\mathcal{P}$, we define the Bellman projection
$q_{P,H}:=\arg\min_{q\in\overline{H}_{P}}E_{P}\!\left[\{Y_{0}-\mathcal{T}_{P}(q)(A_{0},S_{0})\}^{2}\right],$ (7)
that is, the best approximation to $q_{P}$ within $H$ under the norm $\|\mathcal{T}_{P}(\cdot)\|_{P}$. This induces a working parameter $\Psi_{H}(P):=E_{P}\bigl[m(S_{0},A_{0},q_{P,H})\bigr],$ which coincides with the original parameter $\Psi$ on $\mathcal{P}_{H}$ but remains well-defined over the full nonparametric model $\mathcal{P}$. The projected function $q_{P,H}$ satisfies the projected Bellman equation $\mathcal{T}_{P}(q_{P,H})=r_{P,H}$, where $r_{P,H}:=\arg\min_{r\in\mathcal{T}_{P}(\overline{H}_{P})}\|r_{P}-r\|_{P}$ is the $L^{2}(P)$-projection of the reward function $r_{P}$ onto the range $\mathcal{T}_{P}(\overline{H}_{P})$.
A model-robust and efficient estimator of the projection parameter $\Psi_{H}(P_{0})$ is obtained by adding a correction term to $\psi_{n,H}$:
$\psi_{n,H}^{*}$ $=\psi_{n,H}+\frac{1}{n}\sum_{i=1}^{n}\big(r_{n}-\widehat{\mathcal{T}}_{n}(q_{n,H})\big)(A_{0,i},S_{0,i})\left\{\alpha_{n,H}(A_{0,i},S_{0,i})-\gamma V^{\pi}(\alpha_{n,H})(S_{1,i})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H})(A_{0,i},S_{0,i})\right\}.$
Here $q_{n,H}\in H$, $\alpha_{n,H}\in H$, $\widehat{\mathcal{T}}_{n}(q_{n,H})$, $\widehat{\mathcal{T}}_{n}(\alpha_{n,H})$, and $r_{n}$ are estimators of $q_{0,H}$, $\alpha_{0,H}$, $\mathcal{T}_{0}(q_{0,H})$, $\mathcal{T}_{0}(\alpha_{0,H})$, and $r_{0}$, respectively. The estimator $\psi_{n,H}^{*}$ generalizes the DRL estimator $\psi_{n,H}$ from the previous section, which is recovered under correct specification by setting $r_{n}:=\widehat{\mathcal{T}}_{n}(q_{n,H})$, in which case the additional correction term vanishes.
The asymptotic properties of $\psi_{n,H}^{*}$ (and of $\psi_{n,H}$ as a special case) follow from the next theorem, which establishes pathwise differentiability of $\Psi_{H}$, identifies its efficient influence function (EIF), and provides a von Mises expansion (mises1947asymptotic; bickel1993efficient). For each $P\in\mathcal{P}$, we define the function
$\varphi^{*}_{P,H}(s,a,y,s^{\prime})$ $:=\mathcal{T}_{P}(\alpha_{P,H})(a,s)\,\{y+\gamma V^{\pi}(q_{P,H})(s^{\prime})-q_{P,H}(a,s)\}$
$\qquad+\bigl\{\alpha_{P,H}(a,s)-\gamma V^{\pi}(\alpha_{P,H})(s^{\prime})-\mathcal{T}_{P}(\alpha_{P,H})(a,s)\bigr\}\bigl\{r_{P}(a,s)-\mathcal{T}_{P}(q_{P,H})(a,s)\bigr\}$
$\qquad+m(s,a,q_{P,H})-\Psi_{H}(P).$
Theorem 1 (Pathwise differentiability).
Suppose C1 holds. Then the parameter $\Psi_{H}:\mathcal{P}\to\mathbb{R}$ is pathwise differentiable at $P_{0}$ with efficient influence function $\varphi^{*}_{0,H}$. Moreover, for any $\overline{P}\in\mathcal{P}$ for which $\varphi^{*}_{\overline{P},H}$ exists, the parameter admits the expansion
$\Psi_{H}(\overline{P})-\Psi_{H}(P_{0})$ $=-\,P_{0}\varphi^{*}_{\overline{P},H}+R_{H}^{*}(\overline{P},P_{0}),$
$R_{H}^{*}(\overline{P},P_{0})$ $:=\left\langle\mathcal{T}_{\overline{P}}(\alpha_{\overline{P},H})-\mathcal{T}_{0}(\alpha_{0,H}),\,\mathcal{T}_{0}(q_{0,H}-q_{\overline{P},H})\right\rangle_{P_{0}}$
$\qquad+\left\langle(\mathcal{T}_{\overline{P}}-\mathcal{T}_{0})(\alpha_{\overline{P},H}),\,r_{0}-r_{\overline{P}}+\mathcal{T}_{\overline{P}}(q_{\overline{P},H})-\mathcal{T}_{0}(q_{0,H})\right\rangle_{P_{0}}.$
The variance of the EIF $\varphi^{*}_{P_{0},H}$ determines the $P_{0}$-generalized Cramér–Rao efficiency bound for the projection parameter $\Psi_{H}$. When semiparametric restrictions are imposed through $H$, this efficiency bound can be strictly smaller than that of the unrestricted parameter $\Psi$; we return to this point in the next section. Specializing Theorem 1 to the case $H=L^{\infty}(\lambda)$ and to the policy value functional recovers the nonparametric setting of kallus2020double; kallus2022efficiently. Under correct specification ($P_{0},\overline{P}\in\mathcal{P}_{H}$), the second term in the EIF $\varphi^{*}_{\overline{P},H}$ and the second term in the remainder $R_{H}^{*}(\overline{P},P_{0})$ both vanish.
The following theorem establishes double robustness, asymptotic normality and efficiency of $\psi_{n,H}^{*}$ for $\Psi_{H}$. Let $\varphi_{n,H}^{*}$ denote the plug-in estimator of the efficient influence function $\varphi_{0,H}^{*}$ from Theorem 1, obtained by substituting the estimated nuisance functions.
  1. (C2)
    Negligible empirical process: $n^{-\frac{1}{2}}(P_{n}-P_{0})\{\varphi_{n,H}^{*}-\varphi_{0,H}^{*}\}=o_{p}(1)$.
  2. (C3)
    Nuisance estimation rate: Each of the following hold:
    1. (a)
      $\left\|\widehat{\mathcal{T}}_{n}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{0,H})\right\|_{P_{0}}\cdot\left\|\mathcal{T}_{0}(q_{n,H})-\mathcal{T}_{0}(q_{0,H})\right\|_{P_{0}}=o_{p}(n^{-\frac{1}{2}})$
    2. (b)
      $\left\|\mathcal{T}_{0}(\alpha_{n,H})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H})\right\|_{P_{0}}\cdot\left\|\widehat{\mathcal{T}}_{n}(q_{n,H})-\mathcal{T}_{0}(q_{0,H})+r_{0}-r_{n}\right\|_{P_{0}}=o_{p}(n^{-\frac{1}{2}})$
Theorem 2.
Assume C1, C2, and C3. Then, $\psi_{n,H}^{*}-\psi_{0,H}=(P_{n}-P_{0})\varphi_{0,H}^{*}+o_{p}(n^{-\frac{1}{2}})$. Moreover, $\psi_{n,H}^{*}$ is a $P_{0}$-regular and efficient estimator for $\Psi_{H}$ under the nonparametric model .
Condition C2 is an empirical process requirement that holds if $\|\varphi_{n,H}^{*}-\varphi_{0,H}^{*}\|_{P_{0}}=o_{p}(1)$ with the difference lying in a Donsker class, or if sample-splitting or cross-fitting is used to bypass Donsker conditions (van2011cross; chernozhukov2018double). Condition C3 is a doubly robust rate condition requiring sufficiently fast convergence of the nuisance estimators. The first rate condition, $\|\widehat{\mathcal{T}}_{n}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{0,H})\|_{P_{0}}\cdot\|\mathcal{T}_{0}(q_{n,H})-\mathcal{T}_{0}(q_{0,H})\|_{P_{0}}=o_{p}(n^{-1/2})$, generalizes the standard requirement for nonparametric inference on the policy value (kallus2022efficiently; bennett2023minimax; li2024regularized). The second condition is specific to the semiparametric setting and applies only to the model-robust estimator $\psi_{n,H}^{*}$. Under correct specification ($q_{0}\in H$), it is automatically satisfied by the uncorrected estimator $\psi_{n,H}$, which uses $r_{n}:=\widehat{\mathcal{T}}_{n}(q_{n,H})$ and $r_{0}:=\mathcal{T}_{0}(q_{0,H})$.
3.3 Efficiency considerations.
The limiting variance of the semiparametric DRL estimator $\psi_{n,H}^{*}$ is determined by the variability of the weighting function $\mathcal{T}_{0}(\alpha_{0,H})$. For the policy value, this weighting function equals the state–occupancy ratio $d_{0}$ in the fully nonparametric case, and otherwise its $L^{2}(P_{0})$-projection onto the range $\mathcal{T}_{0}(\overline{H}_{P_{0}})$ (assuming $d_{0}$ has finite variance). Thus, the semiparametric efficiency bound is strictly smaller whenever $H$ excludes the nonparametric Riesz representer $\alpha_{0,L^{\infty}(\lambda)}$. For example, if $H$ consists of functions linear in a feature map $\varphi:\mathcal{A}\times\mathcal{S}\to\mathbb{R}^{k}$, then the weighting function becomes the best linear predictor of $d_{0}$, namely $\mathcal{T}_{0}(\alpha_{0,H})(a,s)=\beta_{0}^{\top}\varphi_{0,\gamma}(a,s)$, where $\beta_{0}:=\arg\min_{\beta}E_{0}[(d_{0}(A_{0},S_{0})-\beta^{\top}\varphi_{0,\gamma}(A_{0},S_{0}))^{2}]$ and $\varphi_{0,\gamma}(a,s):=\varphi(a,s)-\gamma\,E_{0}[\int\varphi(a^{\prime},S_{1})\,\pi(a^{\prime}\mid S_{1})\,\nu(da^{\prime})\mid A_{0}=a,S_{0}=s]$, with expectations of vectors taken component-wise. Efficiency gains are largest when the high-variance components of $d_{0}$ lie outside the linear span of $\varphi_{0,\gamma}$.
Similar efficiency gains have been established in linear MDPs, where both the transition kernel and the reward function are assumed to be linear in a known feature map (xie2023semiparametrically). In contrast, we obtain the same benefits under substantially weaker assumptions: it is enough to restrict only the $Q$-function, without imposing a parametric model on the underlying MDP. Intuitively, constraining the $Q$-function implicitly restricts both the reward and the transition kernel through the Bellman equation. The following example makes this efficiency phenomenon explicit in a simple illustrative setting.
Example 3 (Eliminating overlap dependence via time-invariant state structure).
[…]
Interestingly, restricting only the reward function or only the transition kernel does not relax the intertemporal overlap requirement. In both cases, the EIF—and therefore the efficiency bound—continues to depend critically on the state–occupancy ratio $d_{0}$ in (2).
Theorem 3 (EIF with known reward or kernel).
Let $\Psi(P):=E_{0}\!\left[V^{\pi}(q_{P})(S_{0})\right]$ and assume the law of $S_{0}$ is known. The EIF of $\Psi$ when either the reward or kernel is known is:
  1. (i)
    ($r_{0}$ is known) $\varphi_{0,1}(S_{1},A_{0},S_{0})\;=\;d_{0}(A_{0},S_{0})\,\Big\{r_{0}(A_{0},S_{0})+\gamma V^{\pi}(q_{0})(S_{1})-q_{0}(A_{0},S_{0})\Big\},$
  2. (ii)
    ($P_{0}(S_{1}\mid A_{0},S_{0})$ is known) $\varphi_{0,2}(Y_{0},A_{0},S_{0})\;=\;d_{0}(A_{0},S_{0})\,\Big\{Y_{0}-r_{0}(A_{0},S_{0})\Big\}.$
3.4 Sieve estimation and model approximation error
Sieve methods are widely used in econometrics and statistics because they combine the simplicity and finite-sample stability of parametric models with the flexibility of nonparametric function approximation (shen1997methods; newey1997convergence; spnpsieve). Rather than fixing a single working model $H$, one may adopt a sieve strategy, in which model selection for the $Q$-function is performed over a sequence of linear spaces $H_{1}\subset H_{2}\subset\cdots\subset H_{k}\subset\cdots\subset H_{\infty}:=H$ of increasing complexity. For example, $H_{k}$ could be a finite-dimensional series space spanned by trinometric polynomial (JacksonOnAB), spline (Gordon1974BSPLINECA), or wavelet bases (Antoniadis_2007), or it could be infinite-dimensional, such as a space of additively decomposable functions in $k$ variables (stone1985additive; hastie1987generalized; zhang2022regression). Given $H_{n}:=H_{k(n)}$ for a potentially data-dependent sequence $k(n)$, a sieve DRL estimator of $\psi_{0}$ is constructed in the same manner as in Section 3.1, yielding $\psi_{n,H_{n}}$. The next theorem shows that the associated misspecification bias $\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ is second order in the sieve approximation error, extending related bounds from static settings (van2023adaptive; van2025automatic). In Appendix F, we derive the asymptotic theory for sieve DRL estimators.
Theorem 4 (Model approximation error).
Suppose that $q_{0}\in H_{0}$ for some submodel $H_{0}\subseteq H$ (possibly depending on $P_{0}$), and that Condition C1 holds for both $H$ and $H_{0}$. Then
$\Psi_{H_{n}}(P_{0})-\Psi(P_{0})=-\left\langle\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n,0}}),\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\right\rangle_{P_{0}},$
where $H_{n,0}:=H_{n}\oplus H_{0}$ denotes the direct-sum linear model.
For sieve DRL, when $k(n)\uparrow\infty$, the theorem may be applied with $H_{0}:=H$, in which case $\alpha_{0,H_{n,0}}=\alpha_{0,H}$. If $k(n)$ instead converges to a fixed finite value $k_{0}$, one may take $H_{0}:=H_{k_{0}}$. Applying the result with $H_{n}=H$ further implies that standard DRL estimators incur only second-order bias under mild misspecification of $H$.
3.5 Estimation of nuisance functions
We now describe how the $Q$-function $q_{0,H}$ and the Riesz representer $\alpha_{0,H}$ can be estimated using flexible machine learning methods. A standard approach for estimating $q_{0,H}$ in reinforcement learning is fitted $Q$-iteration (FQI) (munos2008finite), an iterative regression algorithm for solving the Bellman equation (1). This equation implies that $q_{0,H}(A_{0},S_{0})$ can be recovered by regressing the pseudo-outcome $Y_{0}+\gamma V^{\pi}(q_{0,H})(S_{1})$ on $(A_{0},S_{0})$. FQI initializes $q_{n,H}^{(0)}:=0$ and repeatedly regresses the updated Bellman target $Y_{0}+\gamma V^{\pi}(q_{n,H}^{(k)})(S_{1})$ on $(A_{0},S_{0})$ over the model class $H$ to obtain $q_{n,H}^{(k+1)}$. Iteration stops when successive updates change little in $\ell^{2}$-norm or when the cross-validated risk stabilizes. The finite-iteration error decreases at rate $\gamma^{k}$, so only $O(\log n)$ iterations are typically required in practice (munos2008finite; agarwal2019reinforcement; fan2020theoretical; hu2025fast). When implemented with gradient descent or boosting, the Bellman target may be updated at each gradient step, often accelerating convergence (baird1999reinforcement). A full description appears in Algorithm 2 in Appendix B.
Estimation of the Riesz representer $\alpha_{0,H}$ can be formulated as a convex–concave min–max problem analogous to those used for occupancy ratio estimation (liu2018breaking; nachum2019dualdice; uehara2020minimax; dikkala2020minimax; kallus2020double). The objective in (6) can be rewritten as $\alpha_{0,H}=\operatorname*{argmin}_{\alpha\in H}\max_{f\in L^{\infty}(\lambda)}L_{0}(\alpha,f)$, where the objective function is
$\begin{aligned} L_{0}(\alpha,f)=E_{0}\Big[\{\alpha(A_{0},S_{0})\}^{2}&-2\gamma\,\alpha(A_{0},S_{0})\,V^{\pi}(\alpha)(S_{1})-2\,m(S_{0},A_{0},\alpha)\\ &-\frac{\gamma^{2}}{2}\big\{\{f(A_{0},S_{0})\}^{2}-2\,V^{\pi}(\alpha)(S_{1})\,f(A_{0},S_{0})\big\}\Big].\end{aligned}$ (8)
Given an estimator $\alpha_{n,H}$ of $\alpha_{0,H}$, the remaining nuisance component in $\mathcal{T}_{0}(\alpha_{0,H})$ may be obtained by regressing $V^{\pi}(\alpha_{n,H})(S_{1})$ on $(A_{0},S_{0})$. Min–max empirical risk minimization is studied in dikkala2020minimax and bennett2023minimax. If the inner maximization is taken over a linear model or a reproducing kernel Hilbert space (RKHS), the maximizer can be computed efficiently in closed form using methods such as kernel ridge regression (kallus2020double; olivas2025source).
4 Superefficient inference via Bellman calibration
4.1 Proposed estimator
The previous section showed that imposing semiparametric structure on the $Q$-function can relax intertemporal overlap requirements and improve efficiency. However, if this structure is misspecified, such estimators may suffer from bias. In contrast, fully nonparametric DRL estimators remain valid but can exhibit high variance—or even lose asymptotic normality—when overlap is limited. To address this, we propose nonparametric superefficient estimators that use the $Q$-function as a one-dimensional adjustment variable in the Bellman equation. This reduction replaces the usual high-dimensional overlap condition with a single-dimensional one. As a result, the estimators are superefficient: their limiting variances lie strictly below the nonparametric efficiency bound (le1953some; van2000asymptotic; benkeser2020nonparametric).
Our key insight is that the $Q$-function $q_{0}$ satisfies a dimension-reduced Bellman equation:
$q_{0}(A_{0},S_{0})=E_{0}\!\left[Y_{0}+\gamma V^{\pi}(q_{0})(S_{1})\mid q_{0}(A_{0},S_{0})\right],$ (9)
which conditions only on the scalar value $q_{0}(A_{0},S_{0})$ rather than on the full state–action pair $(A_{0},S_{0})$. This identity follows by taking the conditional expectation given $q_{0}(A_{0},S_{0})$ of both sides of the Bellman equation in (1), and then applying the law of total expectation. Thus, if $q_{0}$ were known, we could apply DRL using $\{q_{0}(a,s):(a,s)\in\mathcal{A}\times\mathcal{S}\}$ as a sufficient one-dimensional summary of the state–action space. The resulting estimator would be superefficient in the fully nonparametric model—because it is efficient in the corresponding dimension-reduced model—with a limiting variance that depends only on intertemporal overlap conditional on the value of the $Q$-function, rather than on overlap across the full state–action space. Since $q_{0}$ is unknown, we instead apply this approach with an estimator $q_{n}^{*}$ of $q_{0}$. At first glance, this procedure may appear invalid, as it seems highly sensitive to errors in $q_{n}^{*}$. However, we show that this is not the case: the error incurred by substituting $q_{n}^{*}$ for $q_{0}$ enters only at second order.
Our proposed superefficient estimator of $\Psi(P_{0})$ is the calibrated plug-in estimator
$\psi_{n}^{*}:=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n}^{*}),$
where $q_{n}^{*}$ is a Bellman-calibrated estimator of the $Q$-function $q_{0}$, constructed to satisfy the empirical Bellman equation $q_{n}^{*}(a,s)=E_{P_{n}}\!\left[\,Y_{0}+\gamma V^{\pi}(q_{n}^{*})(S_{1})\,\middle|\,q_{n}^{*}(A_{0},S_{0})=q_{n}^{*}(a,s)\right],$ which can be written explicitly as
$q_{n}^{*}(a,s)=\frac{\sum_{i=1}^{n}1\{q_{n}^{*}(A_{0,i},S_{0,i})=q_{n}^{*}(a,s)\}\,\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})\}}{\sum_{i=1}^{n}1\{q_{n}^{*}(A_{0,i},S_{0,i})=q_{n}^{*}(a,s)\}}.$ (10)
As we show, calibration performs implicit debiasing by ensuring that $q_{n}^{*}$ satisfies (9) empirically. Given an initial estimator $q_{n}$, a Bellman-calibrated estimator $q_{n}^{*}$ can be constructed post hoc using fitted $Q$-calibration and isotonic regression, as described in Algorithm 1.
Before presenting the calibration algorithm, we briefly explain why enforcing (10) debiases the plug-in estimator. The key observation is that Bellman calibration makes $\psi_{n}^{*}$ equivalent to an oracle DRL estimator in the dimension-reduced model induced by $q_{n}^{*}$. In particular, (10) is equivalent to requiring orthogonality of the temporal-difference errors: for every transformation $f:\mathbb{R}\to\mathbb{R}$,
$\frac{1}{n}\sum_{i=1}^{n}f\!\left(q_{n}^{*}(A_{0,i},S_{0,i})\right)\Bigl\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}^{*}(A_{0,i},S_{0,i})\Bigr\}=0.$ (11)
The Riesz representer under the dimension-reduced model is of the form $f\circ q_{n}^{*}$ for some choice of $f$. Consequently, (11) guarantees that the bias-correction term in the corresponding DRL estimator vanishes when evaluated with $q_{n}^{*}$ and the true Riesz representer. In principle, one could instead estimate the Riesz representer directly and construct a superefficient DRL estimator by applying an explicit bias correction. However, although this yields an estimator that is asymptotically equivalent to ours, it requires consistent estimation of the representer—typically via a minimax optimization procedure—whereas calibration removes the need to estimate it at all.
4.2 Fitted $Q$-calibration
Our proposed algorithm, fitted $Q$-calibration with isotonic regression, is given in Algorithm 1, where $\mathcal{F}_{\text{iso}}$ denotes the class of monotone non-decreasing (isotonic) functions. Python Code is provided in Appendix E. The method combines isotonic calibration (zadrozny2002transforming), originally developed for classification, with fitted $Q$-iteration (munos2008finite). Operationally, it is equivalent to running fitted $Q$-iteration on the dimension-reduced data $\{q_{n}(A_{0,i},S_{0,i}),\,Y_{0,i}\}$, using monotone functions as the regression class. Its convergence properties therefore follow directly from standard results on fitted $Q$-iteration and isotonic regression. In particular, if $q_{n}$ is trained on data independent of the calibration sample, Section 5 of hu2025fast shows (under mild conditions) that
$\|q_{n}^{*(K)}-q_{0}\|_{P_{0}}\;\lesssim\;\|q_{n}-q_{0}\|_{P_{0}}+O_{p}(n^{-1/3})+O(\gamma^{K}),$
where $\|q_{n}-q_{0}\|_{P_{0}}$ upper bounds the misspecification error of the monotone class, $O_{p}(n^{-1/3})$ is the statistical error from estimating a monotone function, and $O(\gamma^{K})$ is the finite-iteration error. Consequently, calibration does not degrade the performance of $q_{n}$ beyond a typically negligible $O_{p}(n^{-1/3})$ term, and it often improves performance by learning an optimal monotone transformation. The number of iterations $K$ may be selected using a stopping rule such as $\|q_{n}^{*(k+1)}-q_{n}^{*(k)}\|_{P_{n}}<\varepsilon$. Because each isotonic regression update is computationally inexpensive, a tolerance on the order of $\varepsilon=1/n$ works well in practice.
Algorithm 1 Fitted $Q$-Calibration with isotonic regression
0:  Initial estimator $q_{n}$, max iteration $K$
1:  Initialize $q_{n}^{*(0)}:=q_{n}$
2:  for $k=0,1,2,\dots,K-1$ do
3:   $f_{n}^{(k+1)}:=\displaystyle\operatorname*{argmin}_{f\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\left\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*(k)})(S_{1,i})-f(q_{n}(A_{0,i},S_{0,i}))\right\}^{2}$
4:   $q_{n}^{*(k+1)}:=f_{n}^{(k+1)}\circ q_{n}$
5:  end for
6:  return $q_{n}^{*(K)}$
Algorithm 1 enforces (10) by exploiting the fact that isotonic regression is equivalent to histogram regression on data-dependent bins of the initial $Q$-values (barlow1972isotonic; van2023causal). A single isotonic regression fits a step function over a partition of the one-dimensional set $\{q_{n}(a,s):(a,s)\in\mathcal{A}\times\mathcal{S}\}$, with the partition chosen to satisfy monotonicity. For $K=\infty$, the calibrated estimate $q_{n}^{*}(a,s)$ equals the empirical average of the Bellman target $Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})$ over all observations whose initial values $q_{n}(A_{0,i},S_{0,i})$ fall in the same bin as $q_{n}(a,s)$. However, these are precisely the observations whose calibrated values $q_{n}^{*}(A_{0,i},S_{0,i})$ equal $q_{n}^{*}(a,s)$, ensuring (10). An alternative is fixed-bin histogram calibration (zadrozny2001obtaining; gupta2021distribution), but it requires manual bin selection, whereas isotonic regression performs binning automatically (van2025generalized).
To mitigate overfitting, the initial $Q$-function estimator $q_{n}$ in Algorithm 1 should be trained on data that is independent of the data used for calibration. This can be achieved via sample splitting: one part of the data is used to estimate $q_{n}$, and the remaining part is used to calibrate $q_{n}$ and compute the plug-in estimator of the linear functional. To improve data efficiency, we recommend cross-fitting $q_{n}$ and then calibrating the pooled out-of-fold predictions using all available data, as in van2024stabilized; vanderlaan2024automatic. A cross-fitted version of Algorithm 1 is given in Appendix C.
4.3 Asymptotic theory
We show that the calibrated plug-in estimator is efficient for an oracle parameter $\Psi_{q_{0}}$, from which superefficiency for $\Psi$ follows. Following the framework in Section 3.2, for any $q\in L^{\infty}(\lambda)$, we define the projection parameter
$\Psi_{q}(P):=E_{P}\!\left[m(S_{0},A_{0},q_{P,q})\right],\qquad q_{P,q}:=\operatorname*{argmin}_{f\circ q:\,f\in L^{\infty}(\mathbb{R})}E_{P}\!\left[\{Y_{0}-\mathcal{T}_{P,q}(f\circ q)(A_{0},S_{0})\}^{2}\right],$
where $\mathcal{T}_{P,q}(f\circ q)(a,s):=f(q(a,s))-\gamma E_{P}[V^{\pi}(f\circ q)(S_{1})\mid q(A_{0},S_{0})=q(a,s)],$ and $L^{\infty}(\mathbb{R})$ denotes the class of bounded real-valued transformations. Although $\Psi_{q_{0}}$ identifies the same estimand $\psi_{0}$ as $\Psi$, since $\Psi_{q_{0}}(P_{0})=\Psi(P_{0})$, its efficiency bound is typically smaller and less sensitive to violations of intertemporal overlap. The following example illustrates this phenomenon in the case of the policy value.
Example 4 (Efficiency bound under dimension reduction).
For the map $m(a,s,q):=V^{b_{0}}(q)(s)$ with $\pi:=b_{0}$, the $P_{0}$-EIF of $\Psi$ is $(s,a,y,s^{\prime})\mapsto d_{0}(a,s)\{y+\gamma V^{b_{0}}(q_{0})(s^{\prime})-q_{0}(a,s)\}+V^{b_{0}}(q_{0})(s)-\psi_{0},$ where $d_{0}(a,s):=\sum_{t=0}^{\infty}\gamma^{t}\,\frac{d\mathbb{P}_{0}((A_{t},S_{t})=(a,s))}{d\mathbb{P}_{0}((A_{0},S_{0})=(a,s))}$. The $P_{0}$-EIF of $\Psi_{q_{0}}$ is obtained by replacing $d_{0}$ with the dimension-reduced occupancy ratio
$d_{0,q_{0}}(a,s):=\sum_{t=0}^{\infty}\gamma^{t}\,\frac{d\mathbb{P}_{0}(q_{0}(A_{t},S_{t})=q_{0}(a,s))}{d\mathbb{P}_{0}(q_{0}(A_{0},S_{0})=q_{0}(a,s))},$
since $d_{0,q_{0}}(A_{0},S_{0})=E_{0}[d_{0}(A_{0},S_{0})\mid q_{0}(A_{0},S_{0})]$. Thus, the efficiency bound of $\Psi_{q_{0}}$ depends only on intertemporal overlap of the one-dimensional process $\{q_{0}(A_{t},S_{t})\}_{t\in\mathbb{N}}$, rather than the full state–action process.
Although $\Psi_{q_{0}}$ depends on the unknown $q_{0}$, it can be approximated by replacing $q_{0}$ with the estimator $q_{n}^{*}$. The theorem below shows that $\Psi_{q_{n}^{*}}$ agrees with $\Psi_{q_{0}}$ up to second-order error under suitable conditions. The result does not rely on calibration: the theorem remains valid if $q_{n}^{*}$ is replaced by any estimator $q_{n}$. To state the result, we introduce some notation. For a feature map $\phi:\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}^{m}$, define the induced model class $H_{\phi}:=\{f\circ\phi:f:\mathbb{R}^{m}\rightarrow\mathbb{R}\}\cap L^{\infty}(\lambda)$. Let $\alpha_{0,\phi}\in\overline{H}_{P_{0},\phi}$ denote the associated Riesz representer, defined by $E_{0}[m(A_{0},S_{0},q)]=\langle\mathcal{T}_{0,\phi}(\alpha_{0,\phi}),\,\mathcal{T}_{0,\phi}(q)\rangle_{P_{0}}$ for all $q\in\overline{H}_{P_{0},\phi}$. For shorthand, write $d_{0,\phi}:=\mathcal{T}_{0,\phi}(\alpha_{0,\phi})$.
  1. (D1)
    (Invertibility) For $\phi=(q_{0}),$ $(q_{n}^{*})$, and $(q_{0},q_{n}^{*})$, $\mathcal{T}_{0,\phi}:\overline{H}_{P_{0},\phi}\rightarrow\overline{H}_{P_{0},\phi}$ is continuous and invertible.
  2. (D2)
    (Lipschitz continuity) There exists $L\in(0,\infty)$ such that, for all sufficiently large $n$, the bivariate function $(t_{1},t_{2})\mapsto E_{0}\left[d_{0,(q_{n}^{*},q_{0})}(A_{0},S_{0})\,\middle|\,q_{n}^{*}(A_{0},S_{0})=t_{1},\ q_{0}(A_{0},S_{0})=t_{2},\ \mathcal{D}_{n}\right]$ is almost surely Lipschitz continuous with constant $L$.
Theorem 5 (Parameter approximation error is second-order).
Suppose that C1(i) holds with $H=H_{(q_{n}^{*},q_{0})}$ (continuity of linear functional) and D1 holds. Then,
$\Psi_{q_{n}^{*}}(P_{0})-\Psi(P_{0})$ $=\langle d_{0,(q_{n}^{*},q_{0})}-d_{0,q_{n}^{*}},\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}}-q_{0})\rangle_{P_{0}}.$
If D2 also holds, then $\Psi_{q_{n}^{*}}(P_{0})-\Psi(P_{0})=O_{p}\left(\|q_{n}^{*}-q_{0}\|_{P_{0}}\|\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0})\|_{P_{0}}\right).$
Theorem 5 shows that the bias term $\Psi_{q_{n}^{*}}(P_{0})-\Psi(P_{0})$ vanishes provided that the optimal transformation $q_{0,q_{n}^{*}}$ of $q_{n}^{*}$ converges to $q_{0}$, and the difference $d_{0,(q_{n}^{*},q_{0})}-d_{0,q_{n}^{*}}$ converges to zero. Heuristically, since $d_{0,q_{n}^{*}}(A_{0},S_{0})=E_{0}[d_{0,(q_{n}^{*},q_{0})}(A_{0},S_{0})\mid q_{n}^{*}(A_{0},S_{0})]$, this requires that conditioning on the pair $(q_{n}^{*}(A_{0},S_{0}),q_{0}(A_{0},S_{0}))$ asymptotically conveys no more information than conditioning on either component alone. Under D2, this convergence holds whenever $\|q_{n}^{*}-q_{0}\|_{P_{0}}=o_{p}(1)$. Similar conditions for DML with estimated features are assumed in benkeser2017doubly; benkeser2020nonparametric; wang2023super; bonvini2024doubly; vanderlaan2024automatic. Condition D1 is a dimension-reduced version of C1.
We now show that the estimator $\psi_{n}^{*}$ is asymptotically linear and efficient for the oracle parameter $\Psi_{q_{0}}$, and therefore superefficient for $\Psi$. In what follows, for $q\in\{q_{n}^{*},q_{0}\}$, let $\varphi_{0,q}(s,a,y,s^{\prime}):=\mathcal{T}_{0,q}(\alpha_{0,q})(a,s)\{y+\gamma V^{\pi}(q)(s^{\prime})-q(a,s)\}+m(s,a,q)-\Psi_{q}(P_{0})$. The following conditions are analogous to those of Theorem 2 and are discussed in Appendix C.4.
  1. (D3)
    Nuisance estimation rate: $\|q_{n}^{*}-q_{0}\|_{P_{0}}\|\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0})\|_{P_{0}}=o_{p}(n^{-\frac{1}{2}})$.
  2. (D4)
    Empirical process condition: $n^{-\frac{1}{2}}(P_{n}-P_{0})\{\varphi_{n,q_{n}^{*}}-\varphi_{0,q_{0}}\}=o_{p}(1)$.
Theorem 6 (Asymptotic linearity and superefficiency).
Suppose $q_{n}^{*}$ satisfies the empirical Bellman calibration condition (10), and D1D4 hold. Then:
  1. (i)
    $\psi_{n}^{*}$ is asymptotically linear for $\psi_{0}$ with influence function $\varphi_{0,q_{0}}$.
Moreover, $\varphi_{0,q_{0}}$ is the EIF for $\Psi_{q_{0}}$, and:
  1. (ii)
    $\psi_{n}^{*}$ is a $P_{0}$-regular and efficient estimator for $\Psi_{q_{0}}$ under the nonparametric model.
By Theorem 6, the calibrated estimator $\psi_{n}^{*}$ admits the expansion $\psi_{n}^{*}-\psi_{0}=P_{n}\varphi_{0,q_{0}}+o_{p}(n^{-1/2}),$ where $\varphi_{0,q_{0}}$ is the EIF for the oracle parameter $\Psi_{q_{0}}$. Hence $\sqrt{n}(\psi_{n}^{*}-\Psi(P_{0}))$ is asymptotically normal with variance $\sigma_{0}^{2}:=P_{0}\varphi_{0,q_{0}}^{2}$, and inference may be conducted using standard Wald intervals once a consistent estimator of $\sigma_{0}^{2}$ is available. Under regularity conditions, $\sigma_{0}^{2}$ can be consistently estimated by the empirical variance of an estimated influence function $\varphi_{0,q_{n}^{*}}$, which requires estimating $\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})$. Since $q_{n}^{*}$ is discretely valued, this quantity can be computed efficiently using matrix formulas for finite-state Markov chains (Appendix C.2). As an alternative, $\sigma_{0}^{2}$ may be estimated via a bootstrap that resamples only the calibration step while keeping the initial $Q$-function estimator $q_{n}$ fixed; theoretical justification for this approach is given in vanderlaan2024automatic.
Since $\psi_{n}^{*}$ is efficient for the dimension-reduced parameter $\Psi_{q_{0}}$, which leverages knowledge of $q_{0}$, it is superefficient for the target parameter $\Psi$. The superefficiency of $\psi_{n}^{*}$ comes at the cost of irregularity for $\Psi$: the estimator may exhibit non-vanishing asymptotic bias under sampling from a local alternative $P_{0,hn^{-1/2}}$ to $P_{0}$, with $\sqrt{n}\{\psi_{n}^{*}-\Psi(P_{0,hn^{-1/2}})\}\underset{P_{0,hn^{-1/2}}}{\rightsquigarrow}N(0,\sigma_{0}^{2})+O(h),$ and inference is therefore not uniformly valid over the nonparametric model (van1997superefficiency; van2023adaptive). Nonetheless, accepting irregularity in exchange for superefficiency can be advantageous in practice (benkeser2020nonparametric; van2023adaptive; moosavi2023costs). Theorem 6 guarantees that $\psi_{n}^{*}$ remains regular and efficient for the oracle parameter $\Psi_{q_{0}}$ at $P_{0}$, meaning $\sqrt{n}\{\psi_{n}^{*}-\Psi_{q_{0}}(P_{0,hn^{-1/2}})\}\underset{\text{\scriptsize$P_{0,hn^{-1/2}}$}}{\rightsquigarrow}N(0,\sigma_{0}^{2})$. Thus, even under sampling from $P_{0,hn^{-1/2}}$, the estimator $\psi_{n}^{*}$ continues to yield valid asymptotic inference for the projected target $\Psi_{q_{0}}(P_{0,hn^{-1/2}})$.
5 Numerical experiments
We consider Example 2 from Section 2.2, where the parameter $\Psi$ represents the long-term causal effect of an A/B test. Following tran2023inferring, the estimand is the long-term ATE $E_{0}\!\left[V^{\pi}(q_{0})\bigl(1,\widetilde{S}_{0}\bigr)-V^{\pi}(q_{0})\bigl(0,\widetilde{S}_{0}\bigr)\right].$ To mimic an online A/B test, we simulate a discrete-state Markov process with four state variables—engagement, churn risk, tenure, and an overlap variable parameterized by $\beta$—each taking values in $\{0,1,2\}$. The parameter $\beta$ governs the degree of intertemporal overlap: larger values correspond to weaker overlap between treatment arms and downstream states. This setup enables us to assess estimator performance under varying overlap conditions, a key challenge in long-term causal inference. Full simulation details are provided in Appendix D.2.
We estimate the $V$-function using fitted $Q$-iteration with gradient-boosted regression trees implemented in lightgbm. We compare three estimators: (i) the Bellman-calibrated plug-in estimator from Section 4, (ii) an adaptive DRL estimator that learns a data-driven model class for the $V$-function, and (iii) the nonparametric DRL estimator of kallus2020double and tran2023inferring. The adaptive DRL method builds $\mathcal{H}$ by one-hot encoding the leaf nodes of a boosted tree model for the $V$-function (Section 3.1 of he2014practical). The Riesz representer is then estimated via min–max optimization, with the inner maximization approximated in closed form by ridge regression on the induced tree features. Confidence intervals for the Bellman-calibrated estimator are obtained via the bootstrap procedure of vanderlaan2024automatic.
(a) Limited overlap ($\beta=0.6$).
(b) Varying overlap ($\gamma=0.8$).
Figure 2: Bias, standard error (SE), and coverage across discount factors $\gamma$ for setting with limited intertemporal overlap. Subfigure (d) compares Bellman calibration and nonparametric methods in low-overlap settings ($\beta=0.7,0.8,0.9$); adaptive DRL (tree) results closely resemble the nonparametric method and are omitted for clarity.
Figure 2 compares the three estimators across varying levels of intertemporal overlap ($\beta$) and discounting ($\gamma$). Results for the good ($\beta=0$) and moderate ($\beta=0.3$) overlap settings are given in Appendix D.2. Performance depends strongly on overlap: when overlap is good, all methods achieve low bias, small standard errors, and near-nominal coverage. As overlap worsens ($\beta=0.3,0.6$), the nonparametric estimator deteriorates—exhibiting high variance and undercoverage—reflecting its sensitivity to inverse weighting. Subfigure (d) highlights this at $\gamma=0.8$, where the nonparametric method shows large bias, variance, and coverage error, while Bellman calibration maintains coverage near 95% as $n$ increases. Across all settings, Bellman calibration is the most stable, consistently achieving low bias and variance. The adaptive DRL estimator improves upon the nonparametric baseline but remains more variable than Bellman calibration under poor overlap, likely due to the higher complexity of $\mathcal{H}$. Finally, as $\gamma$ increases, the nonparametric estimator becomes increasingly unstable, whereas Bellman calibration remains accurate and low-variance, demonstrating robustness in long-horizon settings.
6 Conclusion and future work
We introduced a calibrated plug-in approach for inference in dynamic settings by leveraging $Q$-function calibration. A natural direction for future research is to extend the calibrated DML framework of vanderlaan2024automatic from static to dynamic models. Such an extension could show that jointly calibrating the $Q$-function and its Riesz representer yields doubly robust inference (benkeser2017doubly), enabling valid inference even when one of the two components is inconsistently estimated or converges slowly. In particular, calibrating the $Q$-function in DRL can improve robustness to misspecification of the density ratio. For this reason, we recommend $Q$-function calibration in nonparametric DRL, even when the density ratio is used for bias correction.
Additional directions include extending the methodology to time-inhomogeneous MDPs and to correlated data settings in which multiple transitions are observed per individual. Finally, the framework extends naturally to well-posed linear inverse problems more broadly, by replacing the Bellman operator with an invertible linear operator.
Appendix A Sufficient conditions for C1
In this section, we provide sufficient conditions for the invertibility of $\mathcal{T}_{P}$ in Condition C1. The key idea is that $\mathcal{T}_{P}$ differs from the identity operator by a compact operator under mild conditions. As a result, $\mathcal{T}_{P}$ is a Fredholm operator of index zero on $\overline{H}_{P}$, and the inverse problem defining $q_{P,H}$ is a Fredholm equation of the second kind (conway1994course).
In what follows, we define the operator $\mathcal{K}_{P}:L^{\infty}(\lambda)\rightarrow L^{\infty}(\lambda)$ pointwise as $\mathcal{K}_{P}(h)(a,s):=E_{P}\left[V^{\pi}(h)(S_{1})\mid A_{0}=a,S_{0}=s\right]$. A key property of the Bellman operator $\mathcal{T}_{P}:=I-\gamma\mathcal{K}_{P}$ is that it is a $\gamma$-perturbation of the identity operator $I:L^{\infty}(\lambda)\rightarrow L^{\infty}(\lambda)$, which maps each function to itself, by the conditional expectation operator $\mathcal{K}_{P}$. The following condition ensures existence of $q_{P,H}$.
  1. (A1)
    Compactness and Fredholm property: The operator $\mathcal{K}_{P}$ is continuous on $H$ with respect to $\|\cdot\|_{P}$, and its unique continuous extension is compact on the closure $(\overline{H}_{P},\|\cdot\|_{P})$.
Condition A1 is a mild requirement that holds under appropriate assumptions on the state transition probabilities. Suppose that the operator $\mathcal{K}_{P}$ admits the following integral representation for $h\in L^{\infty}(\lambda)$:
$\mathcal{K}_{P}(h)(a,s)=\int h(a^{\prime},s^{\prime})\,K_{P}(a^{\prime},s^{\prime}\mid a,s)\,dP_{A_{0},S_{0}}(a^{\prime},s^{\prime}),$
where the kernel is given by $K_{P}(a^{\prime},s^{\prime}\mid a,s):=\frac{\pi(a^{\prime}\mid s^{\prime})}{b_{P}(a^{\prime}\mid s^{\prime})}\,\frac{dP(S_{1}=s^{\prime}\mid A_{0}=a,S_{0}=s)}{dP(S_{0}=s^{\prime})}.$ Condition A1 holds with $H=L^{\infty}(\lambda)$ when the state and action spaces are compact subsets of $\mathbb{R}^{d}$, and the kernel function $K_{P}$ is continuous and bounded. In this case, the kernel is square-integrable with respect to the product measure $P_{A_{0},S_{0}}\otimes P_{A_{0},S_{0}}$, so $\mathcal{K}_{P}$ defines a Hilbert–Schmidt operator on $L^{2}(P_{A_{0},S_{0}})$, and is therefore compact (conway1994course). Related conditions for the nonparametric well-posedness of the integral equation for the $Q$-function were proposed in Section 3.2 of chen2022well. Notably, Condition A1 becomes even less stringent under stronger semiparametric restrictions imposed through $H$, and holds trivially when $H$ is finite-dimensional.
The following theorem shows that Condition A1 ensures not only the existence but also the uniqueness of the Bellman projection for almost all discount factors $\gamma$. We begin by introducing the following condition.
  1. (A2)
    Invertibility: $\gamma^{-1}$ is not an eigenvalue of $\mathcal{K}_{P}$ when restricted to $(\overline{H}_{P},\|\cdot\|_{P})$.
Theorem 7 (Existence and uniqueness of solution).
Assume A1 holds at $P\in\mathcal{P}$. Then, the range $\mathcal{T}_{P}(\overline{H}_{P})$ is a closed subspace of $L^{2}(P_{A_{0},S_{0}})$ and there exists an element $q_{P,H}\in H$ satisfying (7). Moreover, if A2 also holds, then $\mathcal{T}_{P}\colon(\overline{H}_{P},\|\cdot\|_{P})\rightarrow L^{2}(P_{A_{0},S_{0}})$ has a bounded inverse on its range, and $q_{P,H}=\mathcal{T}_{P}^{-1}(r_{P,H})$ is the unique solution.
Condition A1 ensures that the Bellman integral operator $\mathcal{T}_{P}$ can be continuously extended to a map from $\overline{H}_{P}$ to $\overline{H}_{P}$, and this map differs from the identity by a compact operator. Consequently, $\mathcal{T}_{P}=I-\gamma\mathcal{K}_{P}$ is a Fredholm operator of index zero on $\overline{H}_{P}$, and the inverse problem defining $q_{P,H}$ is a Fredholm equation of the second kind (conway1994course). The closedness of the range $\mathcal{T}_{0}(\overline{H}_{0})$, guaranteed by Theorem 7, ensures the existence of the Bellman projection $q_{P,H}$ in (7). Condition A2 further guarantees uniqueness, and holds whenever $\gamma<\|\mathcal{K}_{P}\|_{P}^{-1}$. By the compactness of $\mathcal{K}_{P}$, the spectrum of $\mathcal{T}_{P}$ is countable with $\gamma^{-1}$ as the only possible accumulation point. Hence, the Bellman projection $q_{P,H}$ is unique for almost all discount factors $\gamma$.
Historically, the existence and uniqueness of the $Q$-function $q_{0}$ for all $\gamma\in(0,1)$ in $L^{\infty}(\lambda)$ is established using Banach’s fixed point theorem, leveraging the fact that $\mathcal{K}_{P}$ is a contraction on $L^{\infty}(\lambda)$. However, this argument does not apply to the Bellman projection $q_{0,H}$, since the $L^{2}$-projection operator is not a contraction in the supremum norm. Moreover, even for $q_{0}$, this approach does not guarantee uniqueness of the solution in $L^{2}(P_{0,A_{0},S_{0}})$, as it only ensures injectivity of $\mathcal{T}_{P}$ on the dense subspace $L^{\infty}(\lambda)$.
Appendix B Additional details on nuisance estimation
Algorithm 2 Fitted Q-Iteration
0:  Data, Function class $H$, number of iterations $K$;
1:  Initialize $q_{n,H}^{\pi,(0)}:=0$;
2:  for $k=0,1,\ldots,K-1$ do
3:   Set value function $V_{n,H}^{\pi,(k)}:s\mapsto\int q_{n,H}^{\pi,(k)}(a^{\prime},s)\pi(a^{\prime}\mid s)da^{\prime};$
4:   Update $q_{n,H}^{\pi,(k+1)}\in H$ by estimating:
$\arg\min_{q\in H}E_{0}\left[\{Y_{0}+\gamma V_{n,H}^{\pi,(k)}(S_{1})-q(A_{0},S_{0})\}^{2}\right];$
5:  end for
6:  Set $q_{n,H}:=q_{n,H}^{\pi,(K)}$;
7:  return $q_{n,H}$;
Appendix C Additional details for Section 4
C.1 Empirical Bellman calibration of Alg. 1
Lemma 8.
Suppose that $f_{n}^{*}$ is the fixed point isotonic regression solution to the calibrated fitted Q-iteration algorithm in Alg. 1, such that:
$f_{n}^{*}=\operatorname*{argmin}_{f\in\mathcal{F}_{iso}}\sum_{i=1}^{n}\{Y_{0,i}+\gamma V^{\pi}(f_{n}^{*}\circ q_{n})(S_{1,i}))-f(q_{n}(A_{0,i},S_{0,i}))\}^{2}.$
Then, $q_{n}^{*}:=f_{n}^{*}\circ q_{n}$ satisfies, for each transformation $f:\mathbb{R}\rightarrow\mathbb{R}$, the empirical orthogonality condition:
$\sum_{i=1}^{n}f(q_{n}^{*}(A_{0,i},S_{0,i}))\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}^{*}(A_{0,i},S_{0,i})\}=0.$
Proof.
The proof follows from Lemma 4 in van2023causal with minor notational changes. Recall that $f_{n}^{*}$ is the unique càdlàg piecewise constant solution of the isotonic regression problem with jumps occurring only at observed values of $q_{n}$. For any transformation $f:\mathbb{R}\rightarrow\mathbb{R}$, we claim that $f_{n}^{*}+\varepsilon(h\circ f_{n}^{*})$ is monotone nondecreasing for $\varepsilon$ sufficiently close to zero. To see this, note that $f_{n}^{*}$ is a step function with only finitely many jumps. As a consequence, $h\circ f_{n}^{*}$ is also a step function with the same jump points as $f_{n}^{*}$. By taking $\varepsilon$ close enough to zero, we can guarantee that the maximum jump size of $\varepsilon(h\circ f_{n}^{*})$ is smaller than the minimum jump size of $f_{n}^{*}$. For all $\varepsilon$ sufficiently close to zero, it must then be the case that $f_{n}^{*}+\varepsilon(h\circ f_{n}^{*})$ is also monotone nondecreasing and, thus, an element of $\mathcal{F}_{\text{iso}}$. Since $f_{n}^{*}$ is the empirical risk minimizer over $\mathcal{F}_{\text{iso}}$, we must have that
$\frac{d}{d\varepsilon}\sum_{i=1}^{n}\left\{Y_{0,i}+\gamma V^{\pi}(f_{n}^{*}\circ q_{n})(S_{1,i}))-(f_{n}^{*}+\varepsilon(h\circ f_{n}^{*}))(q_{n}(A_{0,i},S_{0,i}))\right\}^{2}\bigg|_{\varepsilon=0}=0,$
which implies that
$\sum_{i=1}^{n}f(q_{n}^{*}(A_{0,i},S_{0,i}))\left\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}^{*}(A_{0,i},S_{0,i})\right\}=0.$
Since the transformation $f$ was arbitrary, the result then follows. ∎
C.2 Estimation of Riesz representer for calibrated fitted Q-iteration
An empirical plug-in estimator $d_{n}^{*}$ of $\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})$ is given by $T_{n,q_{n}^{*}}(\alpha_{n,q_{n}^{*}})$, where
$T_{n,q_{n}^{*}}(\alpha)=(a,s)\mapsto\alpha(a,s)-\gamma E_{P_{n}}[\alpha(A,S_{1})\mid q_{n}^{*}(A_{0},S_{0})=q_{n}^{*}(a,s)]$
is the empirical Bellman operator induced by the empirical distribution $P_{n}$ of $\{(S_{0,i},A_{0,i},S_{1,i})\}_{i=1}^{n}$, and $\alpha_{n,q_{n}^{*}}$ is obtained by solving
$\operatorname*{argmin}_{f\circ q_{n}^{*};f:\mathbb{R}\rightarrow\mathbb{R}}\frac{1}{n}\sum_{i=1}^{n}\left[\{T_{n,q_{n}^{*}}(f\circ q_{n}^{*})(A_{0,i},S_{0,i})\}^{2}-2m(S_{0,i},A_{0,i},f\circ q_{n}^{*})\right],$
which is a parametric M-estimation problem that can be computed using numerical solvers.
C.3 Bellman Calibration with Cross-Fitting
Algorithm 3 presents a cross-fitted variant of fitted $Q$-calibration (Algorithm 1). We note that the isotonic calibration step in Algorithm 1 should not itself be cross-fitted. Instead, isotonic regression should be applied using the cross-fitted estimates obtained by pooling the out-of-fold predictions. Importantly, this additional fitting step does not compromise the theoretical guarantees of DML, as $\mathcal{F}_{\text{iso}}$ has controlled complexity, being a Donsker class (vanderlaan2024automatic).
Algorithm 3 Isotonic Fitted $Q$-calibration with Cross-Fitting
1:  Input: Data $\{(S_{0,i},A_{0,i},Y_{0,i},S_{1,i})\}_{i=1}^{n}$
2:      Cross-fitted estimators $\{q_{n}^{(-i)}\}_{i=1}^{n}$, with each $q_{n}^{(-i)}$ independent of $(S_{0,i},A_{0,i},Y_{0,i},S_{1,i})$
3:      Stopping threshold $\varepsilon\approx 0$
4:  Initialize $q_{n}^{*(-i,0)}(a,s):=q_{n}^{(-i)}(a,s)$ for each $i$;
5:  for $k=0,1,2,\ldots$ do
6:   Compute $f_{n}^{(k+1)}$ by solving:
$\operatorname*{argmin}_{f\in\mathcal{F}_{\text{iso}}}\sum_{i=1}^{n}\left\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*(-i,k)})(S_{1,i})-f(q_{n}^{(-i)}(A_{0,i},S_{0,i}))\right\}^{2};$
7:   Update $q_{n}^{*(-i,k+1)}(a,s):=f_{n}^{(k+1)}(q_{n}^{(-i)}(a,s))$ for each $i$;
8:   if $\|q_{n}^{*(-i,k+1)}-q_{n}^{*(-i,k)}\|_{P_{n}}<\varepsilon$ then
9:    Set $q_{n}^{*(-i)}:=q_{n}^{*(-i,k+1)}$ for each $i$;
10:    break;
11:   end if
12:  end for
13:  return $\{q_{n}^{*(-i)}\}_{i=1}^{n}$;
The procedure begins with an initial collection of Q-function estimators $\{q_{n}^{(-i)}\}_{i=1}^{n}$, each trained on a subsample that excludes the $i$th observation to preserve independence. These out-of-fold predictions serve as the input to an iterative calibration procedure that updates each fold-specific Q-function through composition with an isotonic regression fit. At each iteration, a global isotonic calibrator $f_{n}^{(k+1)}\in\mathcal{F}_{\text{iso}}$ is trained to regress the pseudo-outcomes $Y_{0,i}+\gamma V^{\pi}(q_{n}^{*(-i,k)})(S_{1,i})$ on the out-of-fold predictions $q_{n}^{(-i)}(A_{0,i},S_{0,i})$. The updated estimate $q_{n}^{*(-i,k+1)}$ is then defined as the composition $f_{n}^{(k+1)}\circ q_{n}^{(-i)}$. The process continues until convergence in $L_{2}(P_{n})$, at which point the final cross-fitted calibrated estimators $\{q_{n}^{*(-i)}\}_{i=1}^{n}$ are returned. This calibration procedure leverages the entire dataset while preserving fold-level independence in each update, thereby recovering full-sample efficiency without violating the theoretical guarantees of cross-fitted debiased machine learning. We refer the reader to van2024stabilized and vanderlaan2024automatic for additional examples of calibration on cross-fitted estimates. Importantly, this additional fitting from calibration does not compromise the theoretical guarantees of DML, as $\mathcal{F}_{\text{iso}}$ has controlled complexity, being a Donsker class (vanderlaan2024automatic; rabenseifner2025calibration).
C.4 Discussions of conditions of Theorem 6
Condition D1 ensures pathwise differentiability of $\Psi_{q_{0}}$, $\Psi_{q_{n}^{*}}$, and $\Psi_{(q_{n}^{*},q_{0})}$, and requires overlap only in the lower-dimensional feature space induced by the $Q$-function. It is therefore significantly weaker than the condition required for pathwise differentiability of $\Psi$ in the full nonparametric model. Condition D3 holds if the calibrated estimator $q_{n}^{*}$ is $o_{p}(n^{-1/4})$-consistent for $q_{0}$ in $L^{2}(P_{0})$, and if the best approximation $q_{0,q_{n}^{*}}^{\pi}$ to $q_{0}$ given $q_{n}^{*}$ satisfies $\|\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}}^{\pi})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0})\|_{P_{0}}=o_{p}(n^{-1/4})$. Together, these imply both $\|q_{n}^{*}-q_{0}\|_{P_{0}}=o_{p}(n^{-1/4})$ and the required approximation rate in the transformed space. General results on isotonic calibration suggest that $q_{n}^{*}$ converges at least as fast—and possibly faster—than the initial estimator $q_{n}$, up to an asymptotically negligible error of order $n^{-1/3}$ (van2025generalized). Condition D4 is satisfied if (i) $\|\varphi_{n,q_{n}^{*}}-\varphi_{0,q_{0}}\|_{P_{0}}=o_{p}(1)$, and (ii) the difference lies in a Donsker class or if the initial estimator $q_{n}$ is estimated via sample splitting or cross-fitting (van2024stabilized; rabenseifner2025calibration). Under boundedness and D2, the first condition holds whenever $\|q_{n}^{*}-q_{0}\|_{P_{0}}=o_{p}(1)$. This empirical process condition holds under mild conditions when Alg. 3 is used.
Appendix D Additional details on experiments
D.1 Simulation design
[…]
D.2 Additional experimental results
(a) Good overlap ($\beta=0$).
(b) Moderate overlap ($\beta=0.3$).
Figure 3: Bias, standard error (SE), and coverage across discount factors $\gamma$ for various values of $\beta$.
Appendix E Code
The Python implementation of fitted $Q$-calibration is provided below.
Our implementation provides a scalable and efficient version of fitted $Q$-calibration using regression tree estimation via xgboost. We exploit the fact that isotonic regression is equivalent to fitting a monotone-constrained regression tree of unconstrained depth. The max_depth and min_child_weight parameters in isotonic regression can be specified or tuned without affecting calibration. We recommend setting max_depth to a large value, as the monotonicity constraint automatically limits effective depth. Setting min_child_weight to a small nonzero value (e.g., 30) helps mitigate the poor boundary behavior of standard isotonic regression (van2024stabilized).
import numpy as np
import xgboost
def isoreg_with_xgboost(x, y, max_depth=15, min_child_weight=20, weights=None):
    Fits isotonic regression using XGBoost with monotonic constraints to ensure
    non-decreasing predictions as the predictor variable increases.
    Args:
        x (np.array): A vector or matrix of predictor variables.
        y (np.array): A vector of response variables.
        max_depth (int, optional): Maximum depth of the trees in XGBoost.
                                   Default is 15.
        min_child_weight (float, optional): Minimum sum of instance weights
                                            needed in a child node. Default is 20.
        weights (np.array, optional): A vector of weights for each instance.
                                      If None, all instances are equally weighted.
    Returns:
        function: A prediction function that takes a new predictor variable x
                  and returns the model’s predicted values.
    # Create an XGBoost DMatrix object from the data with optional weights
    data = xgb.DMatrix(data=np.asarray(x).reshape(len(y), -1),
    label=np.asarray(y), weight=weights)
    # Set parameters for the monotonic XGBoost model
    params = {
        ’max_depth’: max_depth,
        ’min_child_weight’: min_child_weight,
        ’monotone_constraints’: "(1)",  # Enforce monotonic increase
        ’eta’: 1,
        ’gamma’: 0,
        ’lambda’: 0
    # Train the model with one boosting round
    iso_fit = xgb.train(params=params, dtrain=data, num_boost_round=1)
    # Prediction function for new data
    def predict_fn(x):
        Predicts output for new input data using the trained isotonic regression model.
        Args:
            x (np.array): New predictor variables as a vector or matrix.
        Returns:
            np.array: Predicted values.
        x = np.atleast_2d(x).T if x.ndim == 1 else x
        data_pred = xgb.DMatrix(data=x)
        pred = iso_fit.predict(data_pred)
        return pred
    return predict_fn
def fitted_q_calibration_iso(q_n, Y0, A0, S0, S1, gamma, V_pi, K=5, sample_weight=None):
    Fitted Q-Calibration with isotonic regression and early stopping.
    Parameters
    q_n : callable
        Base Q-estimator. Must support vectorized evaluation:
        ‘q_n(A0, S0)‘ -> array of shape (n,), and ‘q_n(a, s)‘ for new inputs.
    Y0 : array-like of shape (n,)
[…]
    A0 : sequence length n
        Actions at time 0 used to compute base scores x_i = q_n(A0, S0).
    S0 : sequence length n
        States at time 0 corresponding to A0.
    S1 : sequence length n
[…]
    gamma : float in [0, 1)
        Discount factor.
    V_pi : callable
        Value operator. Called as ‘V_pi(q_fn, S1)‘ and must return the vector
[…]
        Here ‘q_fn(a, s)‘ is a callable Q-function supplied by this routine.
    K : int, default=5
        Maximum number of calibration iterations.
    sample_weight : array-like of shape (n,), optional
        Nonnegative weights for the isotonic fit. If None, uses uniform weights.
    Returns
    q_star : callable
[…]
        where K’ <= K due to early stopping.
    Y0 = np.asarray(Y0, float)
    n = len(Y0)
    eps = 1.0 / n
    sample_weight = np.ones(n) if sample_weight is None
    else np.asarray(sample_weight, float)
    # Fixed base scores x_i = q_n(A0, S0)
    x = q_n(A0, S0)
[…]
    q_star = q_n
    q_prev_vals = q_star(A0, S0)
    for _ in range(K):
        # Bellman targets with current q_star
        y = Y0 + gamma * V_pi(q_star, S1)
        # One-step isotonic fit via your monotone XGBoost helper
        iso = isoreg_with_xgboost(x, y,
        max_depth=15, min_child_weight=20, weights=sample_weight)
        # Compose: q_star(a, s) = iso( q_n(a, s) )
        q_star = (lambda iso=iso, q_n=q_n: (lambda a, s: iso(q_n(a, s))))()
        # Early stopping using empirical L2(P_n)
        q_new_vals = q_star(A0, S0)
        diff = np.sqrt(np.mean((q_new_vals - q_prev_vals) ** 2))
        if diff < eps:
            break
        q_prev_vals = q_new_vals
    return q_star
Appendix F DRL with sieve estimation and model selection
F.1 General Approach
Selecting an appropriate working model is challenging and can compromise inference due to model misspecification bias. Adaptive debiased machine learning (ADML) (van2023adaptive) provides a unified framework that combines debiased estimation with data-driven model selection to construct superefficient estimators of smooth functionals, adapting to the structure of the nuisance components. By learning model assumptions or feature representations directly from data, ADML facilitates valid inference while mitigating misspecification bias. The calibrated plug-in estimator introduced in the previous section is a special case of ADML. In this section, we extend the ADML framework to the MDP setting, showing how semiparametric DRL can be combined with model selection to construct estimators that adapt to the functional form of the $Q$-function $q_{0}$, going beyond calibration alone. We refer to this extension as Adaptive DRL (ADRL).
Let $H_{n}\subseteq H$ be a data-dependent working model for the $Q$-function $q_{P}$, selected via model selection. ADRL posits the existence of a fixed but unknown oracle submodel $H_{0}\subseteq H$, determined by the true $Q$-function $q_{0}$, such that the approximation error between $H_{n}$ and $H_{0}$ vanishes asymptotically. Suppose we have estimators $q_{n,H_{n}}\in H_{n}$ and $\alpha_{n,H_{n}}\in H_{n}$ for $q_{0,H_{n}}$ and $\alpha_{0,H_{n}}$, respectively. The ADRL estimator of $\Psi(P_{0})$ is the DRL estimator:
$\psi_{n,H_{n}}=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H_{n}})+\frac{1}{n}\sum_{i=1}^{n}\widehat{\mathcal{T}}_{n}(\alpha_{n,H})\left\{Y_{0,i}+\gamma V^{\pi}(q_{n,H_{n}})(S_{1,i})-q_{n,H_{n}}(A_{0,i},S_{0,i})\right\},$
which targets the data-adaptive parameter $\Psi_{H_{n}}(P_{0})$. Unlike the model-robust estimator $\psi_{n,H}^{*}$ in Section 3.2, we omit the bias correction term, as it is unnecessary when the model approximation error vanishes asymptotically. Using a novel expansion of the approximation error $\Psi_{H_{n}}(P_{0})-\Psi_{H_{0}}(P_{0})$, we show that $\psi_{n,H_{n}}$ remains $\sqrt{n}$-consistent, asymptotically normal, and efficient for the oracle parameter $\Psi_{H_{0}}$. This oracle parameter coincides with the target $\Psi(P_{0})$ when $q_{0}\in H_{0}$, and often has a smaller efficiency bound, yielding less variable estimates and narrower confidence intervals while preserving unbiasedness. The adaptive plug-in estimator $\psi_{n}^{*}$ based on $Q$-function calibration is a special case corresponding to $H_{n}=\{f\circ q_{n}:f\}$ and $H_{0}=\{f\circ q_{0}:f\}$.
For example, the working model $H_{n}$ could be selected via cross-validated FQI over a sieve of models, i.e., a sequence of increasingly complex classes $H_{1}\subset H_{2}\subset H_{3}\subset\dots\subset H_{\infty}:=H$, where $H$ is a correctly specified model containing $q_{0}$. A plausible oracle submodel $H_{0}$ is the smallest correctly specified class in the sieve that contains $q_{0}$, which can feasibly be approximated via cross-validation. Alternatively, $H_{n}$ could result from a variable selection procedure or a learned feature transformation, with $H_{0}$ corresponding to the asymptotically selected variables or limiting transformation. Data-adaptive methods for learning state-action feature representations have been proposed in pritz2021jointly and pavse2024state. Such transformations can also be derived directly from the fitted FQI model $q_{n,H_{n}}^{\pi}$, for example, by one-hot encoding the leaf nodes of trees in a random forest or gradient-boosted tree model, as in Section 3.1 of he2014practical.
F.2 Asymptotic theory
The following theorem is key to establishing the validity of our ADRL estimator, showing that the parameter approximation bias $\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ is second-order in the model approximation error and thus asymptotically negligible under certain conditions.
Theorem 9 (Second-order model approximation bias).
Suppose that $q_{0}\in H_{0}$ for some oracle submodel $H_{0}\subseteq H$, depending on $P_{0}$. Assume C1 holds for both $H:=H_{n}$ and $H:=H_{0}$. Then, the oracle approximation error of the working model $H_{n}$ satisfies:
$\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ $=-\langle\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n,0}}),\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\rangle_{P_{0}}$
where $H_{n,0}:=H_{n}\oplus H_{0}$ is the direct sum linear model.
For the approximation error $\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ to vanish, both $\mathcal{T}_{0}(\alpha_{0,H_{n}})$ and $\mathcal{T}_{0}(\alpha_{0,H_{n,0}})$ must converge in $L^{2}(P_{0})$, and $\mathcal{T}_{0}(q_{0,H_{n}})$ must converge to $\mathcal{T}_{0}(q_{0})$. This requires that the learned model $H_{n}$ approximates both the true $Q$-function $q_{0}$ and the union model representer $\alpha_{0,H_{n,0}}$ with vanishing error in the norm $\|\mathcal{T}_{0}(\cdot)\|$. In sieve-based model selection, the event $H_{n}\subseteq H_{0}$ typically holds with high probability, in which case $\mathcal{T}_{0}(\alpha_{0,H_{n,0}})=\mathcal{T}_{0}(\alpha_{0,H_{0}})$, and the condition reduces to requiring that $H_{n}$ grows sufficiently fast. For general model selection procedures, convergence of $\mathcal{T}_{0}(\alpha_{0,H_{n}})$ to $\mathcal{T}_{0}(\alpha_{0,H_{n,0}})$ further requires that any directions (e.g., variables or basis functions) in $H_{n,0}\cap H_{n}^{\perp}$ contribute negligibly to the union model representer $\alpha_{0,H_{n,0}}$.
To further clarify these conditions, suppose the working model $H_{n}:=H_{\phi_{n}}$ and the oracle model $H_{0}:=H_{\phi_{0}}$ are induced by feature transformations. For a transformation $\phi:\mathcal{A}\times\mathcal{Z}\times\mathcal{S}\rightarrow\mathbb{R}^{m}$, define $H_{\phi}:=\{f\circ\phi:f:\mathbb{R}^{m}\rightarrow\mathbb{R}\}$. The combined model $H_{n,0}$ is given by $H_{(\phi_{n},\phi_{0})}$, where $(\phi_{n},\phi_{0})$ denotes the feature map formed by stacking $\phi_{n}$ and $\phi_{0}$. Theorem 4 implies that the approximation bias vanishes if the nuisance functions derived from $\phi_{n}$ and $(\phi_{n},\phi_{0})$ converge to those derived from $\phi_{0}$. The special case $\phi_{n}=q_{n}$ and $\phi_{0}=q_{0}$ recovers Theorem 5 as a corollary. In Lemma 14 (Appendix J.1), we show that with features $X$ and outcome $Y$, the $L^{2}(P_{0})$ error of estimating $E_{0}[Y\mid\phi_{0}(X)]$ using either $E_{0}[Y\mid\phi_{n}(X),\mathcal{D}_{n}]$ or $E_{0}[Y\mid\phi_{n}(X),\phi_{0}(X),\mathcal{D}_{n}]$ is bounded by the feature approximation error $\sqrt{\int\|\phi_{n}(x)-\phi_{0}(x)\|_{\mathbb{R}^{m}}^{2}P_{0,X}(dx)}$. A sufficient condition for this bound is that the map $(t_{1},t_{2})\mapsto E_{0}[Y\mid\phi_{n}(X)=t_{1},\phi_{0}(X)=t_{2},\mathcal{D}_{n}]$ is almost surely Lipschitz continuous, generalizing Condition D2.
We now present our main result on the asymptotic linearity and superefficiency of the ADRL estimator $\psi_{n,H_{n}}$ for $\Psi(P_{0})$. In the following conditions, we define for each $P\in\mathcal{P}$ and model $H$:
$\varphi_{P,H}(s,a,y,s^{\prime}):=\mathcal{T}_{P}(\alpha_{P,H})(a,s)\left\{y+\gamma V^{\pi}(q_{P,H}^{\pi})(s^{\prime})-q_{P,H}^{\pi}(a,s)\right\}+m(s,a,q_{P,H})-\Psi_{H}(P).$
Let $\varphi_{n,H_{n}}$ denote the estimator of the influence function $\varphi_{0,H_{n}}$, obtained by plugging in our nuisance estimators.
  1. (C4)
    Consistency: $n^{-\frac{1}{2}}(P_{n}-P_{0})\{\varphi_{n,H_{n}}-\varphi_{0,H_{n}}\}=o_{p}(1)$.
  2. (C5)
    Nuisance estimation rate: $\|\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\|_{P_{0}}\|\mathcal{T}_{0}(q_{n,H_{n}})-\mathcal{T}_{0}(q_{0,H_{n}})\|_{P_{0}}=o_{p}(n^{-1/2})$.
  3. (C6)
    Stabilization of selected model: $n^{-\frac{1}{2}}(P_{n}-P_{0})\{\varphi_{0,H_{n}}-\varphi_{0,H_{0}}\}=o_{p}(1)$.
  4. (C7)
    Model approximation error: $\|\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n,0}})\|_{P_{0}}\|\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\|_{P_{0}}=o_{p}(n^{-\frac{1}{2}}).$
Theorem 10.
Assume C1 holds with $H=H_{n}$ and $H=H_{0}$. Suppose that $H_{n}$ converges to an oracle submodel $H_{0}$ with $q_{0}\in H_{0}$ in the sense that conditions C4-C7 hold. Then, $\psi_{n,H_{n}}-\Psi(P_{0})=(P_{n}-P_{0})\varphi_{0,H_{0}}+o_{p}(n^{-\frac{1}{2}})$. If, in addition, the conditions of Theorem 1 hold with $H:=H_{0}$, then $\psi_{n,H_{n}}$ is a $P_{0}$-regular and efficient estimator for the oracle parameter $\Psi_{H_{0}}$ under the nonparametric statistical model.
Together, C4 and C5 imply that $\psi_{n,H_{n}}-\Psi_{H_{n}}(P_{0})=(P_{n}-P_{0})\varphi_{0,H_{n}}+o_{p}(n^{-\frac{1}{2}})$, so that $\psi_{n,H_{n}}$ is debiased for the working parameter $\Psi_{H_{n}}(P_{0})$. Conditions C6 and C7, which ensures that data-driven model selection preserves the validity of the debiased machine learning estimator, appear in prior works on ADML (van2023adaptive; van2024adaptive). Condition C6 is an asymptotic stability condition requiring the EIF for the learned model $H_{n}$ to converge to the EIF for the oracle submodel $H_{0}$, which necessitates that $\mathcal{T}_{0}(\alpha_{0,H_{n}})$ and $q_{0,H_{n}}$ are asymptotically consistent with their oracle counterparts. Condition C7 ensures the parameter approximation bias satisfies $\Psi_{n}(P_{0})-\Psi(P_{0})=o_{p}(n^{-\frac{1}{2}})$ in view of Theorem 4.
Appendix G Data-fusion application of ADML
G.1 Background on confounding-robust data-fusion
We consider a data fusion setting in which experimental data ($Z=1$) is augmented with historical control data ($Z=0$), adapting the frameworks of kallus2018removing and van2024adaptive to Markov decision processes. Define the state as $S_{t}=(Z,\widetilde{S}_{t})$, as in Example 2. Suppose the study indicator $Z$ denotes whether a study unit belongs to an observational study $(Z=0)$, such as historical data, or a randomized experiment $(Z=1)$, such as a randomized control trial. We address the data-fusion problem of augmenting experimental data with potentially biased observational data to increase power by effectively enlarging the sample size. While randomization ensures unbiasedness in the experimental study, incorporating observational data can introduce bias from unmeasured confounding unless strong, untestable assumptions are made. Our goal is to combine these data sources in a confounding-robust manner that retains the unbiasedness of the “gold-standard” experiment while enhancing statistical efficiency.
A flexible approach for integrating randomized and observational data involves generating a biased estimate from the pooled data, learning a bias function between the data sources, and adjusting the biased estimate to obtain an unbiased causal effect estimate (kallus2018removing; wu2022integrative; shyr2023multi; van2024adaptive). However, without model assumptions on the data-generating distribution, nonparametric efficient estimators like those constructed via debiased machine learning asymptotically gain no efficiency from including biased observational data (dang2022cross). The efficiency gain from a larger sample size is completely offset by increased variance from learning the confounding bias function. When restricted to regular estimators, efficiency gains require imposing parametric or semiparametric restrictions on the bias function, which may induce estimation and confounding bias if these assumptions are violated. To address this limitation in cross-sectional studies, van2024adaptive proposed the ADML framework, which uses the highly adaptive lasso to learn model assumptions data-adaptively, yielding nonparametric superefficient estimators.
G.2 ADML Methodology
In this section, we extend the ADML framework of van2024adaptive to nonparametric data fusion in MDPs. Following the experimental grounding approach of kallus2018removing, we define the confounding bias function as the difference in $Q$-functions, $b_{0}^{\pi}(a,s)=q_{0}(a,1,s)-q_{0}(a,0,s)$, between the experimental and observational studies. Our parameter of interest is defined as $\Psi(P)=E_{P}[m(A_{0},Z,\widetilde{S}_{0},q_{P})]$, where the mapping $m(s,a,q)\mapsto\int q_{P}(a^{\prime},s)\pi(a^{\prime}\mid s)dz$ represents the long-term effect of policy $\pi$. This effect is averaged over the study-pooled covariate distribution; alternatively, the covariate distribution could be defined over the experimental study—see van2024adaptive for details. We approximate the parameter $\Psi$ by $\Psi_{H}=E_{P}[m(A_{0},Z,\widetilde{S}_{0},q_{P,H})]$, where $H$ is the partially linear working model that imposes the semiparametric restriftion that the bias function $b_{0}^{\pi}$ lies in a Hilbert space $\mathcal{B}\subseteq L^{2}(P_{A_{0},Z,\widetilde{S}_{0}})$. For example, one may posit that $b_{0}^{\pi}$ is well approximated by linear combinations of some finite set of features derived from $(A_{0},Z,\widetilde{S}_{0})$.
As outlined in Section F, an ADML estimator can be constructed by learning a model $H_{n}$ and performing inference for the data-adaptive parameter $\Psi_{H_{n}}$. According to the theory in Section F, the ADML estimator is, under certain conditions, asymptotically linear and efficient for the oracle parameter $\Psi_{H_{0}}$ corresponding to a limiting oracle model $H_{0}$ to which $H_{n}$ converges. For example, $H_{n}$ can be learned using fitted Q-iteration for $q_{0}(a,1,s)$ with the Highly Adaptive Lasso (benkeser2016highly), as in van2024adaptive, where a preliminary estimate of the observational $Q$-function $q_{0}(a,0,s)$ is used as an offset, allowing the bias function to be directly modeled. Consequently, the ADML estimator is asymptotically linear and superefficient for $\Psi(P_{0})$ under the nonparametric model, maintaining robustness to model misspecification while gaining efficiency by pooling the two studies when there is learnable structure in the bias function. This approach directly generalizes the approach of van2024adaptive for short-term causal effects in cross-sectional studies, which corresponds to the case where $\gamma=0$.
We now propose a specific ADML estimator that leverages fitted $Q$-calibration, as described in Section 4, to construct superefficient estimators of $\Psi(P_{0})$ while avoiding the computational challenges and instability of estimating the Riesz representer via min-max optimization. Suppose the observational data is far larger than the experimental study, as often occurs in industrial applications where historical data vastly outnumbers randomized data from A/B tests. In this setting, the observational $Q$-function $q_{0}(a,0,s)$ can be estimated very accurately and is effectively known. Assuming $q_{0}(a,0,s)$ is known, we define the oracle model:
$H_{0}:=\left\{q_{0}(a,0,s)+f(b_{0}^{\pi}(a,s)):f\text{ is a real-valued transformation}\right\}.$
This model consists of all $Q$-functions that agree with $q_{0}$ on the observational study and have a bias function differing from $b_{0}^{\pi}=q_{0}(a,1,s)-q_{0}(a,0,s)$ by a transformation of an arbitrary one-dimensional function. Notably, this model is necessarily correctly specified, as taking $f$ to be the identity function recovers $q_{0}$. Given an initial estimator $b_{n}^{(\pi)}$ of $b_{0}^{(\pi)}$, we approximate the oracle model by the working model: $H_{n}:=\left\{q_{0}(a,0,s)+f(b_{n}^{\pi}(a,s)):f\text{ is a real-valued transformation}\right\},$ and propose to obtain superefficient inference for $\Psi(P_{0})$ by constructing ADML estimators based on $\Psi_{H_{n}}$.
To construct an ADML estimator, we use a modified version of calibrated FQI that incorporates $q_{0}$ as an offset. Define the modified outcome $\widetilde{Y}_{0}=Y_{0}+\gamma V^{\pi}(q_{0})(A_{1},0,S_{1})-q_{0}(A_{0},0,S_{0})$. By the Bellman equation for $q_{0}$, we have $E_{0}[\widetilde{Y}_{0}\mid A_{0},Z,\widetilde{S}_{0},Z=0]=0$ and $E_{0}[\widetilde{Y}_{0}\mid A_{0},S_{0},Z=1]=E_{0}[b^{\pi}_{0}(A_{0},Z,\widetilde{S}_{0})-\gamma V_{b^{\pi}_{0}}^{\pi}(1,S_{1})\mid A_{0},S_{0},Z=1]$, so $b^{\pi}_{0}$ satisfies the Bellman equation for $\widetilde{Y}_{0}$ given $Z=1$. In our modification of Algorithm 1, $q_{n}$ is replaced by $b_{n}^{\pi}$, each $Y_{0,i}$ is replaced by $\widetilde{Y}_{0,i}:=Y_{0,i}+\gamma V^{\pi}(q_{0})(0,S_{1,i})-q_{0}(A_{0,i},0,S_{0,i})$, and calibration is applied only using observations with $Z_{i}=1$. This corresponds to calibrated fitted Q-iteration where calibration uses the class $\{(a,s)\mapsto q_{0}(a,0,s)+zf(b_{n}(a,s)):f\text{ is an isotonic function}\}$. Applying this procedure, we obtain an isotonic-calibrated bias function $b_{n}^{*}$ that satisfies the empirical orthogonality condition for each transformation $f:\mathbb{R}\rightarrow\mathbb{R}$:
$\sum_{i=1}^{n}Z_{i}f(b_{n}^{*}(A_{0,i},S_{0,i}))\{\widetilde{Y}_{0,i}+\gamma V_{b^{*}_{n}}^{\pi}(1,S_{1,i})-b_{n}^{*}(A_{0,i},S_{0,i})\}=0.$
A debiased plug-in estimator is then given by $\frac{1}{\sum_{i=1}^{n}Z_{0,i}}\sum_{i=1}^{n}Z_{0,i}m(A_{0,i},S_{0,i},q^{*}_{n,0})$, where $q^{*}_{n,0}(a,s)=q_{0}(a,0,s)+zb_{n}^{*}(a,s)$ is the calibrated estimator of $q_{0}^{(\pi)}$. An application of the results from Section 4 with the offset outcome $\widetilde{Y}_{0}:=Y_{0}+\gamma V^{\pi}(q_{0})(0,S_{1})-q_{0}(A_{0},0,S_{1,i})$ conditional on $Z_{0}=1$ establishes the asymptotic linearity and superefficiency of this estimator under the stated conditions.
Appendix H Proofs for Section 3
H.1 Proof of Theorem 7 on uniqueness of Bellman projection
Proof of Theorem 7.
(Existence) By Condition A1, the operator $\mathcal{K}_{P}$ is compact on $\overline{H}_{P}$, so $\mathcal{T}_{P}=I_{P}-\gamma\mathcal{K}_{P}$ is a Fredholm operator of index zero on the Banach space $(\overline{H}_{P},\|\cdot\|_{P})$ (Riesz–Schauder theory, conway1994course). In particular, its range $\mathcal{R}:=\mathcal{T}_{P}(\overline{H}_{P})$ is a closed subspace of $L^{2}(P_{A_{0},S_{0}})$.
Now write $Y_{0}=r_{P}(S_{0},A_{0})+\epsilon$, where $E_{P}[\epsilon\mid A_{0},S_{0}]=0$. Then for any $q\in\overline{H}_{P}$,
$E_{P}\left[(Y_{0}-\mathcal{T}_{P}(q))^{2}\right]=E_{P}[\epsilon^{2}]+\|r_{P}-\mathcal{T}_{P}(q)\|_{P}^{2}.$
Hence the minimizer in (7) corresponds to the projection $r_{P,H}:=\operatorname*{argmin}_{r\in\mathcal{R}}\|r_{P}-r\|_{P}^{2}$, which satisfies $r_{P,H}\in\mathcal{R}$. Therefore, there exists $q_{P,H}\in\overline{H}_{P}$ such that $\mathcal{T}_{P}(q_{P,H})=r_{P,H}$, i.e., a solution to (7) exists. Here, we used the standard fact that the projection onto a closed subspace of a Hilbert space always exists and is unique.
(Uniqueness) Under Condition A2, $\gamma^{-1}$ is not an eigenvalue of $\mathcal{K}_{P}$, so $\mathcal{T}_{P}=I_{P}-\gamma\mathcal{K}_{P}$ has a trivial null space on $\overline{H}_{P}$. Since $\mathcal{T}_{P}$ is Fredholm of index zero (conway1994course), injectivity implies surjectivity. Thus, $\mathcal{T}_{P}$ is bijective with a bounded inverse $\mathcal{T}_{P}^{-1}:\mathcal{R}\to\overline{H}_{P}$, and $q_{P,H}=\mathcal{T}_{P}^{-1}(r_{P,H})$ is the unique solution to (7). ∎
H.2 Derivation of EIF in Theorem 1
Proof of EIF in Theorem 1.
By Theorem 7, we have $q_{P,H}=\mathcal{T}_{P}^{-1}(r_{P,H})$, where $\mathcal{T}_{P}^{-1}$ is a bounded linear operator and $r_{P,H}:=\operatorname*{argmin}_{r\in\mathcal{T}_{P}(\overline{H}_{P})}\|r_{P}-r\|_{P}$ is the projection of $r_{P}$ onto the range of $\mathcal{T}_{P}$. Since $\mathcal{T}_{P}$ is invertible on $L^{2}(P)$, its range is closed, and $\overline{H}_{P}$ denotes the closure of $H$ in $L^{2}(P)$. Consequently, $q_{P,H}$ is uniquely identified as an element of the $L^{2}(P)$ closure of $H$.
Let $P\in\mathcal{P}$ be arbitrary, and let $(P_{\varepsilon,\phi}:\varepsilon\in\mathbb{R})$ denote a regular submodel satisfying: (i) $\frac{dP_{\varepsilon,\phi}}{dP}$ exists; (ii) $P_{\varepsilon,\phi}=P$ at $\varepsilon=0$; and (iii) the score at $\varepsilon=0$ is $\phi\in T_{\mathcal{P}}(P)$. We now show that the parameter $\Psi_{H}$ is pathwise differentiable along any such path and satisfies the inner product representation:
$\frac{d}{d\varepsilon}\Psi_{H}(P_{\varepsilon,\phi})\big|_{\varepsilon=0}=\langle\varphi_{P},\phi\rangle_{P},$
where $\varphi_{P}$ denotes the efficient influence function (EIF) of $\Psi_{H}$. To compute the pathwise derivative of $\Psi_{H}$, we will use the representation $\Psi_{H}(P)=\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P}(q_{P,H})\rangle_{P}=\langle\mathcal{T}_{P}(\alpha_{P,H}),\Pi_{P}(r_{P})\rangle_{P}$, which is guaranteed by C1.
We adopt the following notation. Let $W=(S_{0},A_{0},Y_{0},S_{1})$, and let $w=(s,a,y,s^{\prime})$ denote a generic realization of $W$. For each function $h$, define the next-state value function $\overline{V}^{\pi}(h)$ as $w\mapsto V^{\pi}(h)(s^{\prime})$. Throughout, we will view $\overline{V}^{\pi}(h)$ as an element of $L^{2}(P_{S_{1}})$ and $V^{\pi}(h)$ as an element of $L^{2}(P_{S_{0}})$. Define $\Pi_{P}:L^{2}(P)\rightarrow H$ as the $L^{2}(P_{A_{0},S_{0}})$ projection operator onto $H$, given pointwise by $\Pi_{P}f:=\operatorname*{argmin}_{h\in H}\|f-h\|_{P}$.
For each $h\in\mathcal{H}$, we denote the pathwise derivative $d\mathcal{T}_{P}(h):T_{\mathcal{P}}(P)\rightarrow L^{2}(P)$ of $\mathcal{T}_{P}(h)$ by the map $\phi\mapsto\frac{d}{d\varepsilon}\mathcal{T}_{P_{\varepsilon,\phi}}(h)\big|_{\varepsilon=0}$. We can compute this pathwise derivative as follows:
$\frac{d}{d\varepsilon}\mathcal{T}_{P_{\varepsilon,\phi}}(h)\big|_{\varepsilon=0}$ $=-\gamma\frac{d}{d\varepsilon}\int V^{\pi}(h)(s^{\prime},Z)P_{\varepsilon,\phi}(S_{1}=ds^{\prime}\mid A_{0},S_{0})\big|_{\varepsilon=0}$
$=-\gamma E_{P}[V^{\pi}(h)(S_{1})\left\{\phi(W)-E_{P}[\phi(W)\mid A_{0},S_{0}]\right\}\mid A_{0},S_{0}]$
$=-\gamma E_{P}[\left\{V^{\pi}(h)(S_{1})-E_{P}[V^{\pi}(h)(S_{1})\mid A_{0},S_{0}]\right\}\phi(W)\mid A_{0},S_{0}]$
$=-E_{P}[\left\{\gamma V^{\pi}(h)(S_{1})+\mathcal{T}_{P}(h)(A_{0},S_{0})-h(A_{0},S_{0})\right\}\phi(W)\mid A_{0},S_{0}]$
$=E_{P}[\left\{h(A_{0},S_{0})-\gamma\overline{V}^{\pi}(h)(S_{1})-\mathcal{T}_{P}(h)(A_{0},S_{0})\right\}\phi(W)\mid A_{0},S_{0}].$
In the final equality, we used the fact that $V^{\pi}(h)(S_{1})=\overline{V}^{\pi}(h)(S_{1})$ by definition. We will make use of the following expression:
$\langle f,d\mathcal{T}_{P}(h)(\phi)\rangle_{P}=\langle f,\phi\left\{h-\gamma\overline{V}^{\pi}(h)-\mathcal{T}_{P}(h)\right\}\rangle_{P}\quad\text{for all }f\in L^{2}(P),$
where $\overline{V}^{\pi}(h)$ is viewed as a function of $S_{1}$.
The first-order conditions of the optimization problem defining $q_{P,H}$ imply that $q_{P,H}$ satisfies the restricted moment equation:
$\langle\mathcal{T}_{P}(h),r_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}=0\text{ for all }h\in\mathcal{H}.$
By the product rule of differentiation, we have
$\frac{d}{d\varepsilon}\Psi_{H}(P_{\varepsilon,\phi})\Big|_{\varepsilon=0}$ $=\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P_{\varepsilon,\phi}}(q_{P_{\varepsilon,\phi}})\rangle_{P}\Big|_{\varepsilon=0}+\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P}(q_{P,H})\rangle_{P_{\varepsilon,\phi}}\Big|_{\varepsilon=0}.$
First Term.
We know $\mathcal{T}_{P}(q_{P,H})$ is determined by:
$\langle\mathcal{T}_{P}(h),\mathcal{T}_{P}(q_{P,H})\rangle_{P}=\langle\mathcal{T}_{P}(h),Y_{0}\rangle_{P}\text{ for all }h\in\mathcal{H}.$
Hence, taking the pathwise derivative of both sides, we find, for all $h\in\mathcal{H}$, that
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(h),\mathcal{T}_{P_{\varepsilon,\phi}}(q_{P_{\varepsilon,\phi}})\rangle_{P_{\varepsilon,\phi}}\Big|_{\varepsilon=0}=\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(h),Y_{0}\rangle_{P_{\varepsilon,\phi}}\Big|_{\varepsilon=0}.$
Thus, by the chain rule, we have
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(h),\mathcal{T}_{P_{\varepsilon,\phi}}(q_{P_{\varepsilon,\phi}})\rangle_{P}\Big|_{\varepsilon=0}+\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(h),\mathcal{T}_{P}(q_{P,H})\rangle_{P_{\varepsilon,\phi}}\Big|_{\varepsilon=0}=\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(h),Y_{0}\rangle_{P_{\varepsilon,\phi}}\Big|_{\varepsilon=0}.$
Therefore,
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(h),\mathcal{T}_{P_{\varepsilon,\phi}}(q_{P_{\varepsilon,\phi}})\rangle_{P}\Big|_{\varepsilon=0}$ $=\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(h),Y_{0}\rangle_{P}\Big|_{\varepsilon=0}+\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(h),Y_{0}\rangle_{P}\Big|_{\varepsilon=0}$
$\qquad-\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(h),\mathcal{T}_{P}(q_{P,H})\rangle_{P}\Big|_{\varepsilon=0}-\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(h),\mathcal{T}_{P}(q_{P,H})\rangle_{P}\Big|_{\varepsilon=0}$
$=\langle d\mathcal{T}_{P}(h)(\phi),Y_{0}\rangle_{P}+\langle\phi,\mathcal{T}_{P}(h)Y_{0}-E_{P}[\mathcal{T}_{P}(h)Y_{0}]\rangle_{P}$
$\qquad-\langle\phi,\mathcal{T}_{P}(h)\mathcal{T}_{P}(q_{P,H})-E_{P}[\mathcal{T}_{P}(h)\mathcal{T}_{P}(q_{P,H})]\rangle_{P}-\langle d\mathcal{T}_{P}(h)(\phi),\mathcal{T}_{P}(q_{P,H})\rangle_{P}$
$=\langle d\mathcal{T}_{P}(h)(\phi),r_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}+\langle\phi,\mathcal{T}_{P}(h)\{Y_{0}-\mathcal{T}_{P}(q_{P,H})\}\rangle_{P}$
$\qquad+\langle\phi,E_{P}[\mathcal{T}_{P}(h)\mathcal{T}_{P}(q_{P,H})]-E_{P}[\mathcal{T}_{P}(h)Y_{0}]\rangle_{P}.$
The above holds for all $h\in H$, and therefore also for $h=\alpha_{P,H}$ by continuity of the inner product and of $\mathcal{T}_{P}$, since $\alpha_{P,H}$ lies in the $L^{2}(P)$-closure of $H$. Hence, taking $h=\alpha_{P,H}$ and using that $E_{P}[\mathcal{T}_{P}(\alpha_{P,H})\mathcal{T}_{P}(q_{P,H})]=E_{P}[\mathcal{T}_{P}(\alpha_{P,H})Y_{0}]$, it follows that
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P_{\varepsilon,\phi}}(q_{P_{\varepsilon,\phi}})\rangle_{P}\Big|_{\varepsilon=0}$ $=\langle d\mathcal{T}_{P}(\alpha_{P,H})(\phi),r_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}+\langle\phi,\mathcal{T}_{P}(\alpha_{P,H})\{Y_{0}-\mathcal{T}_{P}(q_{P,H})\}\rangle_{P}$
$\qquad+\langle\phi,E_{P}[\mathcal{T}_{P}(\alpha_{P,H})\mathcal{T}_{P}(q_{P,H})]-E_{P}[\mathcal{T}_{P}(\alpha_{P,H})Y_{0}]\rangle_{P}$
$=\langle d\mathcal{T}_{P}(\alpha_{P,H})(\phi),r_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}+\langle\phi,\mathcal{T}_{P}(\alpha_{P,H})\{Y_{0}-\mathcal{T}_{P}(q_{P,H})\}\rangle_{P}$
$=\langle\phi\{\alpha_{P,H}-\gamma\overline{V}_{\alpha_{P,H}}^{\pi}-\mathcal{T}_{P}(\alpha_{P,H})\},r_{P}-\mathcal{T}_{P}(q_{P,H})\rangle_{P}$
$\qquad+\langle\phi,\mathcal{T}_{P}(\alpha_{P,H})\{Y_{0}-r_{P}\}\rangle_{P}.$
Consequently, this derivative component can be expressed as the inner product $\langle\varphi_{P,1},\phi\rangle$ for the gradient component:
$\varphi_{1,P}:w\mapsto\{\alpha_{P,H}(a,s)-\gamma V^{\pi}(\alpha_{P,H})(s^{\prime})\}\{r_{P}(a,s)-\mathcal{T}_{P}(q_{P,H})(a,s)\}+\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y-r_{P}(a,s)\}.$
Second Term.
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P}(q_{P,H})\rangle_{P_{\varepsilon,\phi}}\Big|_{\varepsilon=0}$ $=\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P}(q_{P,H})\rangle_{P}\Big|_{\varepsilon=0}$
$\qquad+\frac{d}{d\varepsilon}\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P}(q_{P,H})\rangle_{P_{\varepsilon,\phi}}\Big|_{\varepsilon=0}.$
To compute this term, we use the Riesz representation property of $\alpha_{P,H}$, which implies:
$\langle\mathcal{T}_{P}(\alpha_{P,H}),\mathcal{T}_{P}(h)\rangle_{P}=E_{P}[m(S_{0},A_{0},h)]\text{ for all }h\in\mathcal{H}.$
Taking the pathwise derivative of both sides and applying the chain rule, we find:
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P_{\varepsilon,\phi}}(h)\rangle_{P_{\varepsilon,\phi}}\big|_{\varepsilon=0}$ $=\frac{d}{d\varepsilon}E_{P_{\varepsilon,\phi}}[m(S_{0},A_{0},h)]\big|_{\varepsilon=0};$
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P}(h)\rangle_{P_{\varepsilon,\phi}}\big|_{\varepsilon=0}+\langle\mathcal{T}_{P}(\alpha_{P,H}),d\mathcal{T}_{P}(h)(\phi)\rangle_{P}\big|_{\varepsilon=0}$ $=\frac{d}{d\varepsilon}E_{P_{\varepsilon,\phi}}[m(S_{0},A_{0},h)]\big|_{\varepsilon=0},$
and, hence,
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P}(h)\rangle_{P_{\varepsilon,\phi}}\big|_{\varepsilon=0}=\frac{d}{d\varepsilon}E_{P_{\varepsilon,\phi}}[m(S_{0},A_{0},h)]\big|_{\varepsilon=0}-\langle\mathcal{T}_{P}(\alpha_{P,H}),d\mathcal{T}_{P}(h)(\phi)\rangle_{P}\big|_{\varepsilon=0},$
where we compute
$\frac{d}{d\varepsilon}E_{P_{\varepsilon,\phi}}[m(S_{0},A_{0},h)]\big|_{\varepsilon=0}=E_{P}[\phi_{S_{0}}(S_{0})m(S_{0},A_{0},h)]=\langle\phi,m(S_{0},A_{0},h)-E_{P}[m(S_{0},A_{0},h)]\rangle_{P}.$
The above holds for all $h\in H$ and, therefore, also for $q_{P,H}$ by continuity of the inner product and of $\mathcal{T}_{P}$, since $q_{P,H}$ lies in the $L^{2}(P)$-closure of $H$ by Theorem 7. Thus, setting $h=q_{P,H}$, we find that:
$\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P}(q_{P,H})\rangle_{P_{\varepsilon,\phi}}\big|_{\varepsilon=0}=\langle\phi,m(S_{0},A_{0},q_{P,H})-E_{P}[m(S_{0},A_{0},q_{P,H})]\rangle_{P}-\langle\mathcal{T}_{P}(\alpha_{P,H}),d\mathcal{T}_{P}(q_{P,H})(\phi)\rangle_{P}.$
By the definition of $d\mathcal{T}_{P}(q_{P,H})(\phi)$, we have:
$\langle\mathcal{T}_{P}(\alpha_{P,H}),d\mathcal{T}_{P}(q_{P,H})(\phi)\rangle_{P}$ $=-E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\gamma V^{\pi}(q_{P,H})(S_{1})+\mathcal{T}_{P}(q_{P,H})-q_{P,H}(A_{0},S_{0})\right\}\phi(Z)\right].$
Thus, $\frac{d}{d\varepsilon}\langle\mathcal{T}_{P_{\varepsilon,\phi}}(\alpha_{P_{\varepsilon,\phi}}),\mathcal{T}_{P}(q_{P,H})\rangle_{P_{\varepsilon,\phi}}\big|_{\varepsilon=0}:=\langle\varphi_{P,2},\phi\rangle_{P}$ for the gradient component:
$\varphi_{P,2}:w\mapsto\mathcal{T}_{P}(\alpha_{P,H})(a,s)\left\{\gamma V^{\pi}(q_{P,H})(a,s^{\prime})+\mathcal{T}_{P}(q_{P,H})(a,s)-q_{P,H}(a,s)\right\}\,+\,m(s,a,q_{P,H})-\Psi_{H}(P).$
EIF.
Putting it all together, the EIF $\varphi_{P}:=\varphi_{P,1}+\varphi_{P,2}$ is:
$w\mapsto$ $\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y-r_{P}(a,s)\}+\{\alpha_{P,H}(a,s)-\gamma V^{\pi}(\alpha_{P,H})(a,s^{\prime})\}\{r_{P}(a,s)-\mathcal{T}_{P}(q_{P,H})(a,s)\}$
$\qquad+\mathcal{T}_{P}(\alpha_{P,H})(a,s)\left\{\gamma V^{\pi}(q_{P,H})(a,s^{\prime})+\mathcal{T}_{P}(q_{P,H})(a,s)-q_{P,H}(a,s)\right\}$
$\qquad+m(s,a,q_{P,H})-\Psi_{H}(P).$
Assuming a correct model ($\mathcal{T}_{P}(q_{P,H})=r_{P}$), it simplifies to:
$\varphi_{P}:w\mapsto$ $\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y+\gamma V^{\pi}(q_{P,H})(a,s^{\prime})-q_{P,H}(a,s)\}$
$\qquad+m(s,a,q_{P,H})-\Psi_{H}(P).$
H.3 Derivation of von Mises expansion in Theorem 1
We establish the following generalization of the von Mises expansion in Theorem 1, which does not require the nuisance components in the influence function to be compatible with any single distribution $P\in\mathcal{P}$. Let $\widehat{q}_{H},\widehat{\alpha}_{H}\in H$, $\widehat{r}_{P}\in L^{\infty}(\lambda)$, and let $\widehat{\mathcal{T}}:L^{\infty}(\lambda)\rightarrow L^{\infty}(\lambda)$ be an arbitrary map. Let $\widehat{P}\in\mathcal{P}$ be a distribution such that $q_{\widehat{P}}^{\pi}=\widehat{q}_{H}$, so that $\Psi(\widehat{P}):=E_{\widehat{P}}[m(S_{0},A_{0},\widehat{q}_{H})]$. Define $\widehat{\varphi}^{*}_{H}$ as the function
$(s,a,y,s^{\prime})\mapsto\;$ $\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(a,s)\left\{y+\gamma V^{\pi}(\widehat{q}_{H})(s^{\prime})-\widehat{q}_{H}(a,s)\right\}$
$+\left\{\widehat{\alpha}_{H}(a,s)-\gamma V^{\pi}(\widehat{\alpha}_{H})(s^{\prime})-\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(a,s)\right\}\left\{r_{0}(a,s)-\widehat{\mathcal{T}}(\widehat{q}_{H})(a,s)\right\}$
[…]
Theorem 11 (Functional von Mises expansion).
Assume that C1 holds at $P_{0}$. Then, the parameter expansion satisfies: $\Psi_{H}(\widehat{P})-\Psi_{H}(P_{0})=-P_{0}\widehat{\varphi}^{*}_{H}+\widehat{R}_{H}^{*}(P_{0})$, where:
$\widehat{R}_{H}^{*}(P_{0})$ $:=P_{0}\left[\{\mathcal{T}_{0}(\widehat{\alpha}_{H})-\mathcal{T}_{0}(\alpha_{0,H})\}(\mathcal{T}_{0}(q_{0,H})-\mathcal{T}_{0}(\widehat{q}_{H}))\right]$
[…]
$\qquad+P_{0}\left[\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})-\mathcal{T}_{0}(\widehat{\alpha}_{H})\}(\widehat{\mathcal{T}}(\widehat{q}_{H})-\mathcal{T}_{0}(\widehat{q}_{H}))\right].$
Proof of Theorem 11.
[…]
$\widehat{R}_{H}^{*}(P)$ $:=\Psi_{H}(\widehat{P})-\Psi_{H}(P)+P\widehat{\phi}^{*}_{H}$
$=E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\{r_{P}(A_{0},S_{0})-\widehat{r}(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\{\widehat{\alpha}_{H}(A_{0},S_{0})-\gamma V^{\pi}(\widehat{\alpha}_{H})(S_{1})\}\{\widehat{r}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\left\{\gamma V^{\pi}(\widehat{q}_{H})(S_{1})+\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\widehat{q}_{H}(A_{0},S_{0})\right\}\right]$
$\qquad+E_{P}\left[m(S_{0},A_{0},\widehat{q}_{H})-m(S_{0},A_{0},q_{P,H})\right].$
By Riesz representation theorem, it holds that
$E_{P}\left[m(S_{0},A_{0},\widehat{q}_{H})-m(S_{0},A_{0},q_{P,H})\right]$ $=E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(q_{P,H})(A_{0},S_{0})\right\}\right]$
$=E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})-r_{P}(A_{0},S_{0})\right\}\right],$
where we used that $\mathcal{T}_{P}(q_{P,H})$ is the $L^{2}(P)$ projection of $r_{P}$ onto $\mathcal{T}_{P}(\overline{H}_{P})$. In addition, applying the law of iterated expectations applied to the second and third terms, we find:
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\{r_{P}(A_{0},S_{0})-\widehat{r}(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\{\widehat{r}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right]$
$\qquad+E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})-r_{P}(A_{0},S_{0})\right\}\right].$
Next, adding and subtracting, the first and third term can be rewritten as
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{r_{P}(A_{0},S_{0})-\widehat{r}(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\{r_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right]$
$\qquad+E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})-r_{P}(A_{0},S_{0})\right\}\right].$
Adding and subtracting again, the third and fourth terms can be rewritten as
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{r_{P}(A_{0},S_{0})-\widehat{r}(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\{r_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right]$
$\qquad+E_{P}\left[\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-r_{P}(A_{0},S_{0})\right\}\right].$
Combining the second and fourth term, we find
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{r_{P}(A_{0},S_{0})-\widehat{r}(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\left\{\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\{r_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right].$
Using that $\mathcal{T}_{P}(q_{P,H})$ is the $L^{2}(P)$ projection of $r_{P}$ onto $\mathcal{T}_{P}(\overline{H}_{P})$, we can show that
$E_{P}\left[\left\{\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\{r_{P}(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\hskip 28.45274pt=E_{P}\left[\left\{\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\{\mathcal{T}_{P}(q_{P})(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right].$
Substituting this expression, we conclude that
$\widehat{R}_{H}^{*}(P)$ $=E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})\right\}\{r_{P}(A_{0},S_{0})-\widehat{r}(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\left\{\mathcal{T}_{P}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\{\mathcal{T}_{P}(q_{P})(A_{0},S_{0})-\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})\}\right]$
$\qquad+E_{P}\left[\left\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\alpha_{P,H})(A_{0},S_{0})\right\}\left\{\widehat{\mathcal{T}}(\widehat{q}_{H})(A_{0},S_{0})-\mathcal{T}_{P}(\widehat{q}_{H})(A_{0},S_{0})\right\}\right].$
The first result then follows. In the case where $P,\widehat{P}\in\mathcal{P}_{H}$, we have that $r_{P}=\mathcal{T}_{P}(q_{P,H})$ and $\widehat{r}=\widehat{\mathcal{T}}(\widehat{q}_{H})$. In this case, the expression simplifies to:
$\widehat{R}_{H}^{*}(P)=E_{P}\left[\{\widehat{\mathcal{T}}(\widehat{\alpha}_{H})-\mathcal{T}_{P}(\alpha_{P,H})\}(\mathcal{T}_{P}(q_{P,H})-\mathcal{T}_{P}(\widehat{q}_{H}))]\right].$
Proof of Theorem 3.
We use the standard fact that the EIF in a restricted model is obtained by projecting the unrestricted (nonparametric) EIF onto the tangent space of that restricted model (bickel1993efficient). By kallus2022efficiently, the nonparametric EIF is
$\varphi_{0}(S_{1},Y_{0},A_{0},S_{0})=d_{0}(A_{0},S_{0})\,\Big\{Y_{0}+\gamma V^{\pi}(q_{0})(S_{1})-q_{0}(A_{0},S_{0})\Big\}+V^{\pi}(q_{0})(S_{0})-\psi_{0}.$
When the law of $S_{0}$ is known, the EIF is obtained by projecting $\varphi_{0}$ onto functions with mean zero given $S_{0}$, yielding
$\widetilde{\varphi}_{0}(S_{1},Y_{0},A_{0},S_{0}):=d_{0}(A_{0},S_{0})\,\Big\{Y_{0}+\gamma V^{\pi}(q_{0})(S_{1})-q_{0}(A_{0},S_{0})\Big\}.$
We may decompose $\widetilde{\varphi}_{0}$ as
$\widetilde{\varphi}_{0}(S_{1},Y_{0},A_{0},S_{0})$ $=\varphi_{0,1}(S_{1},A_{0},S_{0})+\varphi_{0,2}(Y_{0},A_{0},S_{0}),$
where $\varphi_{0,1}$ is mean zero in $(S_{1},A_{0},S_{0})$ given $(A_{0},S_{0})$, and $\varphi_{0,2}$ is mean zero in $(Y_{0},A_{0},S_{0})$ given $(A_{0},S_{0})$. Moreover, $\varphi_{0,1}$ and $\varphi_{0,2}$ are orthogonal in $L^{2}(P_{0})$.
The tangent space for the model where the law of $S_{0}$ and $r_{0}$ are known is the Hilbert space of square-integrable, mean-zero functions of $(S_{1},A_{0},S_{0})$ given $(A_{0},S_{0})$. The projection of $\widetilde{\varphi}_{0}$ onto this space is $\varphi_{0,1}$, and hence $\varphi_{0,1}$ is the EIF in this model.
Similarly, when the law of $S_{0}$ and the transition kernel are known, the tangent space is the Hilbert space of square-integrable, mean-zero functions of $(Y_{0},A_{0},S_{0})$ given $(A_{0},S_{0})$. The projection of $\widetilde{\varphi}_{0}$ onto this space is $\varphi_{0,2}$, so $\varphi_{0,2}$ is the EIF in this model. ∎
Appendix I Proofs for Section 3 on semiparametric DRL
I.1 Asymptotic linearity of DRL estimator under correct specification
The following theorem establishes the pathwise differentiability of the parameter $\Psi$ under the semiparametric model $\mathcal{P}_{H}$. For each $P\in\mathcal{P}_{H}$, we define the influence function
$\varphi_{P,H}(s,a,y,s^{\prime}):=\mathcal{T}_{P}(\alpha_{P,H})(a,s)\{y+\gamma V^{\pi}(q_{P})(s^{\prime})-q_{P}(a,s)\}+m(s,a,q_{P})-\Psi(P),$ (12)
where $\alpha_{P,H}\in\overline{H}_{P}$ is the Riesz representer of $h\mapsto E_{P}[m(S_{0},A_{0},h)]$ with respect to $\|\mathcal{T}_{P}(\cdot)\|_{P}$.
Theorem 12 (Pathwise differentiability).
Suppose Condition C1 holds and that $P_{0}\in\mathcal{P}_{H}$. Then, $\Psi:\mathcal{P}_{H}\rightarrow\mathbb{R}$ is pathwise differentiable at $P_{0}$, with influence function $\varphi_{0,H}$. Moreover, for any $\overline{P}\in\mathcal{P}_{H}$ for which $\varphi_{\overline{P},H}$ exists, the following von Mises expansion holds:
$\Psi(\overline{P})-\Psi(P_{0})+P_{0}\varphi_{\overline{P},H}=\left\langle\mathcal{T}_{\overline{P}}(\alpha_{\overline{P},H})-\mathcal{T}_{0}(\alpha_{0,H}),\,\mathcal{T}_{0}(q_{0}-q_{\overline{P}})\right\rangle_{P_{0}}.$
In the following conditions and theorem, let $\varphi_{n,H}$ denote the estimator of the IF $\varphi_{0,H}$ from Theorem 12, obtained by plugging in our nuisance estimators.
  1. (C8)
    Consistency: $n^{-\frac{1}{2}}(P_{n}-P_{0})\{\varphi_{n,H}-\varphi_{0,H}\}=o_{p}(1)$.
  2. (C9)
    Nuisance estimation rate: $\left\|\mathcal{T}_{0}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{0,H})\right\|_{P_{0}}\cdot\|\mathcal{T}_{0}(q_{n,H})-\mathcal{T}_{0}(q_{0,H})\|_{P_{0}}=o_{p}(n^{-\frac{1}{2}})$
Theorem 13 (Asymptotic linearity under correct specification).
Suppose that $P_{0}\in\mathcal{P}_{H}$, meaning $q_{0,H}\in\overline{H}_{P}$. Assume C1 holds, as well as C8 and C9. Then, $\psi_{n,H}-\Psi_{H}(P_{0})=(P_{n}-P_{0})\varphi_{0,H}+o_{p}(n^{-\frac{1}{2}})$. Moreover, $\psi_{n,H}$ is locally robust to misspecification as it is a $P_{0}$-regular and efficient estimator for the working parameter $\Psi_{H}$ under the nonparametric model .
Proof of Theorem 13 .
By C1, we can apply the von Mises expansion in Theorem 12 to conclude that
$\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H}^{\pi})-\Psi(P_{0})=-P_{0}\varphi_{n,H}+R_{n,H}(P_{0}),$
where:
$R_{n,H}(P_{0})$ $:=P_{0}\left[\left\{\widehat{\mathcal{T}}_{n}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{n,H})\right\}\mathcal{T}_{0}(q_{n,H}^{\pi}-q_{0})\right]$
By C9 and the Cauchy-Schwarz inequality, we have that $R_{n,H}(P_{0})=o_{p}(n^{-1/2})$. Thus,
$\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H}^{\pi})-\Psi(P_{0})=-P_{0}\varphi_{n,H}+o_{p}(n^{-1/2}).$
Using the definition of the one-step estimator $\psi_{n,H}=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H}^{\pi})+P_{n}\varphi_{n,H}$, we have
$\psi_{n,H}-\Psi(P_{0})=(P_{n}-P_{0})\varphi_{n,H}+o_{p}(n^{-1/2}).$
By C8, it follows that
$\psi_{n,H}-\Psi(P_{0})$ $=(P_{n}-P_{0})\varphi_{0,H}+(P_{n}-P_{0})\{\varphi_{n,H}-\varphi_{0,H}\}+o_{p}(n^{-1/2})$
$=(P_{n}-P_{0})\varphi_{0,H}+o_{p}(n^{-1/2}),$
as desired. Under A2 and correct specification of $H$, we have that the influence function $\varphi_{0,H}=\varphi_{0,H}^{*}$ is the $P_{0}$-EIF of $\Psi_{H}$ by Theorem 1. Thus, $\psi_{n,H_{n}}$ is a $P_{0}$-regular and efficient estimator for the working parameter $\Psi_{H}$ under the nonparametric model. ∎
I.2 Asymptotic linearity and efficiency of model-robust DRL estimator
Proof of Theorem 2.
By C1, we can apply the von Mises expansion in Theorem 1 to conclude that
$\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H}^{\pi})-\Psi_{H}(P_{0})=-P_{0}\varphi_{n,H}^{*}+R_{n,H}^{*}(P_{0}),$
where:
$R_{n,H}^{*}(P_{0})$ $:=P_{0}\left[\left\{\widehat{\mathcal{T}}_{n}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{n,H})\right\}(r_{0}-r_{n})\right]$
$\qquad+P_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H})-\mathcal{T}_{0}(\alpha_{0,H})\}(\mathcal{T}_{0}(q_{0,H})-\mathcal{T}_{0}(q_{n,H}^{\pi}))\right]$
$\qquad+P_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H})\}(\mathcal{T}_{0}(q_{n,H}^{\pi})-\widehat{\mathcal{T}}_{n}(q_{n,H}^{\pi}))\right].$
By C3 and the Cauchy-Schwarz inequality, we have that $R_{n,H}^{*}(P_{0})=o_{p}(n^{-1/2})$. Thus,
$\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H}^{\pi})-\Psi_{H}(P_{0})=-P_{0}\varphi_{n,H}^{*}+o_{p}(n^{-1/2}).$
Using the definition of the one-step estimator $\psi_{n,H}^{*}=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n,H}^{\pi})+P_{n}\varphi_{n,H}^{*}$, we have
$\psi_{n,H}^{*}-\Psi_{H}(P_{0})=(P_{n}-P_{0})\varphi_{n,H}^{*}+o_{p}(n^{-1/2}).$
By C2, it follows that
$\psi_{n,H}^{*}-\Psi_{H}(P_{0})$ $=(P_{n}-P_{0})\varphi_{0,H}^{*}+(P_{n}-P_{0})\{\varphi_{n,H}^{*}-\varphi_{0,H}^{*}\}+o_{p}(n^{-1/2})$
$=(P_{n}-P_{0})\varphi_{0,H}^{*}+o_{p}(n^{-1/2}),$
as desired. We have that the influence function $\varphi_{0,H}^{*}$ is the EIF of $\Psi_{H}$ by Theorem 1. Thus, $\psi_{n,H_{n}}^{*}$ is a $P_{0}$-regular and efficient estimator for the working parameter $\Psi_{H}$ under the nonparametric model.
Appendix J Proofs for Section 4 on calibrated FQI
J.1 Lemma bounding approximation error of estimated features
In the following lemma, let $X$ be a covariate and $Y\in\mathbb{R}$ be an outcome. For feature transformations $\varphi_{n},\varphi_{0}$, denote $f_{(\varphi_{n},\varphi_{0})}:x\mapsto E_{0}[Y_{0}\mid\varphi_{n}(X)=\varphi_{n}(x),\varphi_{0}(X)=\varphi_{0}(x)]$, $f_{\varphi_{n}}:x\mapsto E_{0}[Y_{0}\mid\varphi_{n}(X)=\varphi_{n}(x)]$, and $f_{\varphi_{0}}:x\mapsto E_{0}[Y_{0}\mid\varphi_{0}(X)=\varphi_{0}(x)]$. The following is a generalization of Lemma 1 in vanderlaan2024automatic.
Lemma 14.
Suppose that $(t_{1},t_{2})\mapsto E_{0}[f_{(\varphi_{n},\varphi_{0})}(X)\mid\varphi_{n}(X)=t_{1},\varphi_{0}(X)=t_{2},\mathcal{D}_{n}]$ is almost surely $L$-Lipschitz continuous. Then,
$\|f_{\varphi_{n}}-f_{(\varphi_{n},\varphi_{0})}\|_{P_{0}}\lesssim\|\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}\|_{P_{0}}\text{ and }\|f_{\varphi_{n}}-f_{\varphi_{0}}\|_{P_{0}}\lesssim\|\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}\|_{P_{0}}.$
Proof.
For any real-valued function $f:\mathcal{X}\rightarrow\mathbb{R}$ and a vector-valued function $v:\mathcal{X}\rightarrow\mathbb{R}^{k}$ with $k\in\mathbb{N}$, we define the conditional expectation projection operator $\Pi_{v}:\mathcal{H}\rightarrow\mathcal{H}$ pointwise as $\Pi_{v}f:=\operatorname*{argmin}_{\theta\in\Theta}\|f-\theta\circ v\|$, where $\Theta$ consists of all functions from $\mathbb{R}^{k}\rightarrow\mathbb{R}$. Whenever $v$ and $f$ are nonrandom functions, we have that $\Pi_{v}f:(a,w)\mapsto E_{0}[f(A,W)\,|\,v(A,W)=v(a,w)]$.
Let $g:\mathbb{R}^{k}\times\mathbb{R}^{k}\rightarrow\mathbb{R}$ be a Lipschitz continuous function with constant $L>0$. By Lipschitz continuity, we have that
$\left|g(\varphi_{n}(x),\varphi_{0}(x))-E[g(\varphi_{n}(X),\varphi_{0}(X))|\varphi_{n}(X)=\varphi_{n}(x)]\right|$
$=\left|E[g(\varphi_{n}(x),\varphi_{0}(x))-g(\varphi_{n}(x),\varphi_{0}(X))|\varphi_{n}(X)=\varphi_{n}(x)]\right|$
$\leq E[\left|g(\varphi_{n}(x),\varphi_{0}(x))-g(\varphi_{n}(x),\varphi_{0}(X))\right||\varphi_{n}(X)=\varphi_{n}(x)]$
$\leq LE[\left\|\varphi_{0}(x)-\varphi_{0}(X)\right\|_{\mathbb{R}^{d}}|\varphi_{n}(X)=\varphi_{n}(x)].$
On the event $\{\varphi_{n}(X)=\varphi_{n}(x)\}$, we know
$\|\varphi_{0}(x)-\varphi_{0}(X)\|_{\mathbb{R}^{d}}$ $\leq\|\varphi_{0}(x)-\varphi_{n}(x)\|_{\mathbb{R}^{d}}+\|\varphi_{n}(x)-\varphi_{n}(X)\|_{\mathbb{R}^{d}}+\|\varphi_{0}(X)-\varphi_{n}(X)\|_{\mathbb{R}^{d}}$
$\leq\|\varphi_{0}(x)-\varphi_{n}(x)\|_{\mathbb{R}^{d}}+\|\varphi_{0}(X)-\varphi_{n}(X)\|_{\mathbb{R}^{d}}.$
Therefore,
$\big|g(\varphi_{n}(x),\varphi_{0}(x))$- $E[g(\varphi_{n}(X),\varphi_{0}(X))|\varphi_{n}(X)=\varphi_{n}(x)]\big|$
$\lesssim E[\left\|\varphi_{0}(x)-\varphi_{0}(X)\right\|_{\mathbb{R}^{d}}\mid\varphi_{n}(X)=\varphi_{n}(x)]$
$\lesssim E[\|\varphi_{0}(x)-\varphi_{n}(x)\|_{\mathbb{R}^{d}}]+E[\|\varphi_{0}(X)-\varphi_{n}(X)\|_{\mathbb{R}^{d}}\mid\varphi_{n}(X)=\varphi_{n}(x)].$
Now, for some function $f$, suppose that $(\varphi_{n}(x),\varphi_{0}(x))\mapsto(\Pi_{\varphi_{n},\varphi_{0}}f)(x)$ is Lipschitz continuous with fixed constant $L$. Then, defining $g:(\widehat{m},m)\mapsto E_{0}[f(X)|\varphi_{n}(X)=\widehat{m},\varphi_{0}(X)=m,\mathcal{D}_{n}]$ and noting by the law of iterated expectation that $\Pi_{\varphi_{n}}\Pi_{\varphi_{n},\varphi_{0}}f=\Pi_{\varphi_{n}}f$, we obtain the following pointwise error bound:
$\left|\Pi_{\varphi_{n},\varphi_{0}}f-\Pi_{\varphi_{n}}f\right|\lesssim\left\|\varphi_{n}-\varphi_{0}\right\|_{\mathbb{R}^{d}}+\Pi_{\varphi_{n}}(\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}).$
Since $\|\Pi_{\varphi_{n}}(\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}})\|_{L^{2}(P_{0})}\leq\|\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}\|_{L^{2}(P_{0})}$ by the properties of projections, it follows that
$\|\Pi_{\varphi_{n},\varphi_{0}}f-\Pi_{\varphi_{n}}f\|_{L^{2}(P)}\lesssim\|\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}\|_{L^{2}(P)}.$
Taking $f:=f_{(\varphi_{n},\varphi_{0})}$ and noting that $\Pi_{\varphi_{n},\varphi_{0}}f_{(\varphi_{n},\varphi_{0})}:=f_{(\varphi_{n},\varphi_{0})}$ and that $\Pi_{\varphi_{n}}f_{(\varphi_{n},\varphi_{0})}:=f_{(\varphi_{n})}$, we conclude that
$\|f_{(\varphi_{n},\varphi_{0})}-f_{\varphi_{n}}\|_{L^{2}(P)}\lesssim\|\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}\|_{L^{2}(P)}.$
By an symmetric argument, swapping $\varphi_{n}$ with $\varphi_{0}$, we conclude that
$\|f_{(\varphi_{n},\varphi_{0})}-f_{\varphi_{0}}\|_{L^{2}(P)}\lesssim\|\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}\|_{L^{2}(P)}.$
Hence, by the triangle inequality, we have that
$\|f_{\varphi_{n}}-f_{\varphi_{0}}\|_{L^{2}(P)}\lesssim\|\|\varphi_{n}-\varphi_{0}\|_{\mathbb{R}^{d}}\|_{L^{2}(P)}.$
J.2 Proofs of main results
Proof of Theorem 5.
We prove the theorem for any given $q_{n}$ (e.g., $q_{n}^{*}$) as calibration is not necessary. We introduce the following notation. For any feature transformation $\phi:\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}^{m}$, let $d_{0,\phi}=\mathcal{T}_{0,\phi}(\alpha_{0,\phi})$, where $\alpha_{0,\phi}:=\operatorname*{argmin}_{\alpha\in\overline{H}_{\phi}}E_{0}\left[\{\mathcal{T}_{0,\phi}(f\circ\phi)(A_{0},S_{0})\}^{2}-2m(S_{0},A_{0},\alpha)\right]$ is the Riesz representer for the function class $H_{\phi}:=\{f\circ\phi;f:\mathbb{R}^{m}\rightarrow\mathbb{R}\}\cap L^{\infty}(\lambda)$ induced by $\phi$. Denote $r_{0,\phi}$ by the map $(a,s)\mapsto E_{0}[Y_{0}\mid\phi(A_{0},S_{0})=\phi(a,s)]$. Let $\alpha_{0,q_{n}}:=\operatorname*{argmin}_{\alpha\in\overline{H}_{q_{n}}}\left\|\mathcal{T}_{0,(q_{n},q_{0})}(\alpha_{0,(q_{n},q_{0})})-\mathcal{T}_{0,q_{n}}(\alpha)\right\|_{P_{0}}$. Then, we have $d_{0,q_{n}}=\mathcal{T}_{0,q_{n}}(\alpha_{0,q_{n}})$.
By Riesz representation theorem, we have that:
$\Psi_{n}(P_{0})-\Psi(P_{0})$ $=\langle d_{0,(q_{n},q_{0})},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}$
$=\langle d_{0,(q_{n},q_{0})}-d_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}$
$\qquad+\langle d_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}.$
Note $\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})=r_{0}$ by correct specification of $H_{(q_{n},q_{0})}$ for $q_{0}$. Thus, by the law of total expectation, we have
$\langle d_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}$ $=\langle d_{0,q_{n}},\mathcal{T}_{0,q_{n}}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}$
$=\langle d_{0,q_{n}},\mathcal{T}_{0,q_{n}}(q_{0,q_{n}})-r_{0}\rangle_{P_{0}}$
$=0,$
where the final equality follows from the orthogonality conditions of the Bellman projection $q_{0,q_{n}}$ of $q_{0}$ with respect to the norm $\|\mathcal{T}_{0,q_{n}}(\cdot)\|_{P_{0}}$ and the fact that $d_{0,q_{n}}$ is in the closed range of $\mathcal{T}_{0,q_{n}}$. Hence,
$\Psi_{n}(P_{0})-\Psi(P_{0})$ $=\langle d_{0,(q_{n},q_{0})}-d_{0,q_{n}},\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\rangle_{P_{0}}.$
Moreover, by the Cauchy–Schwarz inequality, it holds that
$|\Psi_{n}(P_{0})-\Psi(P_{0})|$ $\leq\|d_{0,(q_{n},q_{0})}-d_{0,q_{n}}\|_{P_{0}}\cdot\|\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\|_{P_{0}}.$
For the second part of the theorem, take $d_{0,q_{n}}$ equal to $\mathcal{T}_{0,q_{n}}(\alpha_{0,q_{n}})$. By the invertibility condition in D1, $\mathcal{T}_{0,q_{n}}(\overline{H}_{q_{n}})$ equals $\overline{H}_{q_{n}}$ from the first part of (D2). Hence, we have that $d_{0,q_{n}}$ equals the projection $\operatorname*{argmin}_{d\in\overline{H}_{q_{n}}}\|d_{0,(q_{n},q_{0})}-d\|_{P_{0}}.$ Next, using the second part of (D2), we apply Lemma 14 with $Y:=d_{0,(q_{n},q_{0})}$, $\varphi_{n}:=q_{n}$, and $\varphi_{0}:=q_{0}$ to conclude that
$\|d_{0,(q_{n},q_{0})}-d_{0,q_{n}}\|_{P_{0}}\lesssim\|q_{n}-q_{0}\|_{P_{0}}.$
Thus,
$|\Psi_{n}(P_{0})-\Psi(P_{0})|$ $\leq\|q_{n}-q_{0}\|_{P_{0}}\|\mathcal{T}_{0,(q_{n},q_{0})}(q_{0,q_{n}})-\mathcal{T}_{0,(q_{n},q_{0})}(q_{0})\|_{P_{0}},$
as desired. ∎
Lemma 15 (Calibration corrects plug-in bias).
Suppose $q_{n}^{*}$ is Bellman calibrated (10). Then,
$\psi_{n}^{*}=\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n}^{*})+\frac{1}{n}\sum_{i=1}^{n}\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})(A_{0,i},S_{0,i})\left\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}^{*}(A_{0,i},S_{0,i})\right\}.$
Proof of Lemma 15.
By empirical calibration, for any transformation $f:\mathbb{R}\rightarrow\mathbb{R}$, we have that
$\sum_{i=1}^{n}f(q_{n}^{*}(A_{0,i},S_{0,i}))\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}^{*}(A_{0,i},S_{0,i})\}=0.$
Taking $f$ such that $f\circ q_{n}^{*}=d_{0,q_{n}^{*}}$, we find that
$\sum_{i=1}^{n}d_{0,q_{n}^{*}}(A_{0,i},S_{0,i})\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}^{*}(A_{0,i},S_{0,i})\}=0.$
Therefore, the plug-in estimator $\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n}^{*})$ is equal to the DRL estimator:
$\frac{1}{n}\sum_{i=1}^{n}m(S_{0,i},A_{0,i},q_{n}^{*})+\sum_{i=1}^{n}d_{0,q_{n}^{*}}(A_{0,i},S_{0,i})\{Y_{0,i}+\gamma V^{\pi}(q_{n}^{*})(S_{1,i})-q_{n}^{*}(A_{0,i},S_{0,i})\}.$
Proof of Theorem 6.
By Lemma 8, $q_{n}^{*}$ is empirically calibrated for $q_{0}$. Thus, by Lemma 15, it holds that $P_{n}\varphi_{n,q_{n}^{*}}=0$ and, therefore,
$\psi_{n}^{*}-\Psi_{q_{n}^{*}}(P_{0})$ $=\psi_{n}^{*}+P_{n}\varphi_{n,q_{n}^{*}}-\Psi_{q_{n}^{*}}(P_{0})$
$=P_{n}\varphi_{0,q_{0}}+(P_{n}-P_{0})\{\varphi_{n,q_{n}^{*}}-\varphi_{0,q_{0}}\}$
$\qquad+\psi_{n}^{*}-\Psi_{q_{n}^{*}}(P_{0})+P_{0}\varphi_{n,q_{n}^{*}}.$
We first inspect the term $\psi_{n}^{*}-\Psi_{q_{n}^{*}}(P_{0})+P_{0}\varphi_{n,q_{n}^{*}}$. Note,
$\psi_{n}^{*}-\Psi_{q_{n}^{*}}(P_{0})+P_{0}\varphi_{n,q_{n}^{*}}$ $=P_{0}m(\cdot,q_{n}^{*})-P_{0}m(\cdot,q_{0,q_{n}^{*}})$
$\qquad+\int\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})(a,s)\{y+\gamma V^{\pi}(q_{n}^{*})(a,s^{\prime})-q_{n}^{*}(a,s)\}dP_{0}(s,a,y,s^{\prime})$
$=\langle\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}}),\mathcal{T}_{0,q_{n}^{*}}(q_{n}^{*})-\mathcal{T}_{0,q_{n}^{*}}(q_{0,q_{n}^{*}})\rangle_{P_{0}}$
$\qquad+\int\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})(a,s)\{y+\gamma V^{\pi}(q_{n}^{*})(a,s^{\prime})-q_{n}^{*}(a,s)\}dP_{0}(s,a,y,s^{\prime}),$
where the final equality uses the Riesz representation property of ${d}_{0,q_{n}^{*}}$ and that $q_{n}^{*}\in H_{q_{n}^{*}}$ and $q_{0,q_{n}^{*}}\in H_{q_{n}^{*}}$. Next, note, by the law of iterated expectation, that
$\int\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}})(a,s)\{y+\gamma V^{\pi}(q_{n}^{*})(a,s^{\prime})-q_{n}^{*}(a,s)\}dP_{0}(s,a,y,s^{\prime})$ $=\langle\mathcal{T}_{0,q_{n}^{*}}(\alpha_{0,q_{n}^{*}}),\mathcal{T}_{0,q_{n}^{*}}(q_{0,q_{n}^{*}})-\mathcal{T}_{0,q_{n}^{*}}(q_{n}^{*})\rangle_{P_{0}}.$
Putting it all together, we find that
$\psi_{n}^{*}-\Psi_{q_{n}^{*}}(P_{0})+P_{0}\varphi_{n,q_{n}^{*}}$ $=\langle{d}_{0,q_{n}^{*}},\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{n}^{*})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}})\rangle_{P_{0}}$
$\qquad+\langle{d}_{0,q_{n}^{*}},\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{n}^{*})\rangle_{P_{0}}$
$=0.$
Using that $\psi_{n}^{*}-\Psi_{q_{n}^{*}}(P_{0})+P_{0}\varphi_{n,q_{n}^{*}}=0$, we find that
$\psi_{n}^{*}-\Psi_{q_{n}^{*}}(P_{0})$ $=P_{n}\varphi_{0,q_{0}}+(P_{n}-P_{0})\{\varphi_{n,q_{n}^{*}}-\varphi_{0,q_{0}}\}$
$=P_{n}\varphi_{0,q_{0}}+o_{p}(n^{-\frac{1}{2}}),$
where we used that $(P_{n}-P_{0})\{\varphi_{n,q_{n}^{*}}-\varphi_{0,q_{0}}\}=o_{p}(n^{-\frac{1}{2}})$ by D4.
Finally, applying Theorem 5 and D3, we find that
$\Psi_{q_{n}}(P_{0})-\Psi(P_{0})=O_{p}\left(\|q_{n}^{*}-q_{0}\|_{P_{0}}\|\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0,q_{n}^{*}})-\mathcal{T}_{0,(q_{n}^{*},q_{0})}(q_{0})\|_{P_{0}}\right)=o_{p}(n^{-\frac{1}{2}}).$
Consequently,
$\psi_{n}^{*}-\Psi(P_{0})$ $=P_{n}\varphi_{0,q_{0}}+o_{p}(n^{-\frac{1}{2}}).$
Thus, $\psi_{n}^{*}$ is an asymptotically linear estimator of $\Psi(P_{0})=\Psi_{q_{0}}(P_{0})$ with influence function given by the $P_{0}$-efficient influence function of $\Psi_{q_{0}}$. It follows that $\psi_{n}^{*}$ is a regular and efficient estimator for $\Psi_{q_{0}}$ at $P_{0}$. The result then follows.
Appendix K Proofs for ADML in Section F
Proof of Theorem 4.
By C1 applied with $H:=H_{n,0}$ and Riesz representation theorem, we have that
$\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ $=\langle\mathcal{T}_{0}(\alpha_{0,H_{n,0}}),\mathcal{T}_{0}(q_{0,H_{n}})\rangle_{P_{0}}-\langle\mathcal{T}_{0}(\alpha_{0,H_{n,0}}),\mathcal{T}_{0}(q_{0})\rangle_{P_{0}}$
$=\langle\mathcal{T}_{0}(\alpha_{0,H_{n,0}}),\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\rangle_{P_{0}}.$
Note that $\mathcal{T}_{0}(q_{0,H_{n}})$ is the orthogonal projection in $L^{2}(P)$ of $\mathcal{T}_{0}(q_{0})$ onto $\mathcal{T}_{P}(H_{n})$. The orthogonality conditions of the projection imply that
$\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ $=\langle\mathcal{T}_{0}(\alpha_{0,H_{n,0}})-\mathcal{T}_{0}(\alpha_{0,H_{n}}),\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\rangle_{P_{0}}$
$=-\langle\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n,0}}),\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\rangle_{P_{0}}.$
In the event $H_{n}\subseteq H_{0}$, we have that $H_{n,0}=H_{0}$ and, hence,
$\Psi_{H_{n}}(P_{0})-\Psi(P_{0})$ $=-\langle\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{0}}),\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\rangle_{P_{0}},$
as desired. ∎
K.1 Proof of Theorem 10
Proof of Theorem 10.
Note that
$\psi_{n,H_{n}}-\Psi_{H_{0}}(P_{0})=\psi_{n,H_{n}}-\Psi_{H_{n}}(P_{0})+\Psi_{H_{n}}(P_{0})-\Psi_{H_{0}}(P_{0}).$
Observe that $\psi_{n,H_{n}}=\Psi_{H_{n}}(\widehat{P_{n}})+P_{n}\varphi_{n,H_{n}}$, where $\widehat{P}_{n}\in\mathcal{P}$ is any distribution such that $q_{\widehat{P}_{n},H_{n}}=q_{n,H_{n}}$, $r_{\widehat{P}_{n}}=\widehat{\mathcal{T}}_{n}(q_{n,H_{n}})$, and $T_{\widehat{P}_{n}}(\alpha_{\widehat{P}_{n},H_{n}})=T_{\widehat{P}_{n}}(\alpha_{n,H_{n}})$. Thus, it holds that:
$\psi_{n,H_{n}}-\Psi_{H_{n}}(P_{0})$ $=\Psi_{H_{n}}(\widehat{P_{n}})+P_{n}\varphi_{n,H_{n}}-\Psi_{H_{n}}(P_{0})$
$=P_{n}\varphi_{0,H_{n}}+(P_{n}-P)\{\varphi_{n,H_{n}}-\varphi_{0,H_{n}}\}+R_{n,H_{n}}(P_{0}),$
where $R_{n,H_{n}}(P_{0})=\Psi_{H_{n}}(\widehat{P}_{n})-\Psi_{H_{n}}(P_{0})+P_{0}\varphi_{n,H_{n}}$. By a direct application of C4, we have that $(P_{n}-P)\{\varphi_{n,H_{n}}-\varphi_{0,H_{n}}\}=o_{p}(n^{-\frac{1}{2}})$. Moreover, by application of C6,
$P_{n}\varphi_{0,H_{n}}$ $=(P_{n}-P_{0})\varphi_{0,H_{n}}$
$=(P_{n}-P_{0})\varphi_{0,H_{0}}+(P_{n}-P_{0})\{\varphi_{n,H_{n}}-\varphi_{0,H_{n}}\}$
$=P_{n}\varphi_{0,H_{0}}+o_{p}(n^{-\frac{1}{2}}),$
where we used that $P_{0}\varphi_{0,H_{n}}=0$ and $P_{0}\varphi_{0,H_{0}}=0$. Thus,
$\psi_{n,H_{n}}-\Psi_{H_{n}}(P_{0})$ $=P_{n}\varphi_{0,H_{0}}+R_{n,H_{n}}(P_{0})+o_{p}(n^{-\frac{1}{2}}).$
Next, applying Theorem 1, we find that
$R_{n,H_{n}}(P_{0})$ $=E_{0}\left[\left\{\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{n,H_{n}})\right\}(\mathcal{T}_{0}(q_{0})-\widehat{\mathcal{T}}_{n}(q_{n,H_{n}})\right]$
$\qquad+E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\}(\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{n,H_{n}}))\right]$
$\qquad+E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})\}(\mathcal{T}_{0}(q_{n,H_{n}})-\widehat{\mathcal{T}}_{n}(q_{n,H_{n}}))\right]$
$=E_{0}\left[\left\{\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{n,H_{n}})\right\}(\mathcal{T}_{0}(q_{0})-\mathcal{T}_{0}(q_{n,H_{n}})\right]$
$\qquad+E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\}(\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{n,H_{n}}))\right].$
Since $\mathcal{T}_{0}(q_{0,H_{n}})$ is the $L^{2}(P_{0})$ projection of $q_{0}$ onto $\mathcal{T}_{0}(H_{n})$, we have the orthogonality conditions:
$E_{0}\left[\left\{\mathcal{T}_{0}(\alpha)\right\}(\mathcal{T}_{0}(q_{0})-\mathcal{T}_{0}(q_{0,H_{n}})\right]=0\text{ for all }\alpha\in H_{n}.$
Hence, since $\alpha_{n,H_{n}},\alpha_{0,H_{n}}\in H_{n}$, we have that
$E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\}(\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{n,H_{n}}))\right]=E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\}(\mathcal{T}_{0}(q_{0})-\mathcal{T}_{0}(q_{n,H_{n}}))\right].$
Substituting the above expression, we find that
$R_{n,H_{n}}(P_{0})$ $=E_{0}\left[\left\{\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{n,H_{n}})\right\}(\mathcal{T}_{0}(q_{0})-\mathcal{T}_{0}(q_{n,H_{n}})\right]$
$\qquad+E_{0}\left[\{\mathcal{T}_{0}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\}(\mathcal{T}_{0}(q_{0})-\mathcal{T}_{0}(q_{n,H_{n}}))\right]$
$=E_{0}\left[\left\{\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\right\}(\mathcal{T}_{0}(q_{0})-\mathcal{T}_{0}(q_{n,H_{n}})\right]$
$=O_{p}\left(\|\widehat{\mathcal{T}}_{n}(\alpha_{n,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n}})\|_{P_{0}}\|\mathcal{T}_{0}(q_{n,H_{n}})-\mathcal{T}_{0}(q_{0,H_{n}})\|_{P_{0}}\right)$
$=o_{p}(n^{-\frac{1}{2}}),$
where the final two equalities follow from the Cauchy-Schwarz inequality and C5. Thus,
$\psi_{n,H_{n}}-\Psi_{H_{n}}(P_{0})$ $=P_{n}\varphi_{0,H_{0}}+o_{p}(n^{-\frac{1}{2}}).$
Next we turn to the term $\Psi_{H_{n}}(P_{0})-\Psi_{H_{0}}(P_{0})$. Note, by Theorem 4 and C7, it holds that:
$\Psi_{H_{n}}(P_{0})-\Psi_{H_{0}}(P_{0})$ $=-\langle\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n,0}}),\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\rangle_{P_{0}}$
$=O_{p}(\|\mathcal{T}_{0}(\alpha_{0,H_{n}})-\mathcal{T}_{0}(\alpha_{0,H_{n,0}})\|_{P_{0}}\|\mathcal{T}_{0}(q_{0,H_{n}})-\mathcal{T}_{0}(q_{0})\|_{P_{0}})$
$=o_{p}(n^{-\frac{1}{2}}),$
where the final two equalities follow from the Cauchy-Schwarz inequality and C7.
Putting it all together, we conclude that
$\psi_{n,H_{n}}-\Psi_{H_{0}}(P_{0})$ $=P_{n}\varphi_{0,H_{0}}+o_{p}(n^{-\frac{1}{2}}).$
Since $\varphi_{0,H_{0}}$ is the $P_{0}$-efficient influence function of $\Psi_{0}$, it follows that $\psi_{n,H_{n}}$ is an asymptotically linear, regular, and efficient estimator for $\Psi_{0}$ at $P_{0}$.